### v1

# Can Graph Learning Improve Task Planning?

Abstract
Task planning is emerging as an important research topic alongside the development of large language models (LLMs). It aims to break down complex user requests into solvable sub-tasks, thereby fulfilling the original requests. In this context, the sub-tasks can be naturally viewed as a graph, where the nodes represent the sub-tasks, and the edges denote the dependencies among them. Consequently, task planning is a decision-making problem that involves selecting a connected path or subgraph within the corresponding graph and invoking it. In this paper, we explore graph learning-based methods for task planning, a direction that is orthogonal to the prevalent focus on prompt design. Our interest in graph learning stems from a theoretical discovery: the biases of attention and auto-regressive loss impede LLMs’ ability to effectively navigate decision-making on graphs, which is adeptly addressed by graph neural networks (GNNs). This theoretical insight led us to integrate GNNs with LLMs to enhance overall performance. Extensive experiments demonstrate that GNN-based methods surpass existing solutions even without training, and minimal training can further enhance their performance. Additionally, our approach complements prompt engineering and fine-tuning techniques, with performance further enhanced by improved prompts or a fine-tuned model 111 The code and datasets are available at https://github.com/WxxShirley/GNN4TaskPlan.
1 Introduction
LLM-empowered autonomous agents have recently emerged as a rapidly growing field of research and are considered a significant step towards artificial general intelligence (AGI) [43, 4]. These agents have achieved remarkable successes across a variety of domains, as evidenced by their ability to address complex AI challenges (e.g., HuggingGPT [31] and Visual ChatGPT [46]), excel in gaming environments (e.g., Voyager [40]), and drive innovation in chemical research (e.g., [3]). Within this burgeoning field, task planning emerges as a critical area of study. It involves LLMs autonomously interpreting user instructions, breaking these instructions down into concrete and solvable sub-tasks, and then fulfilling the user’s request by executing each sub-task [29, 31, 30]. For instance, in the case of HuggingGPT [31], task planning involves invoking expert AI models from the HuggingFace website to solve complex AI tasks beyond the capabilities of GPT alone.
Given its practical significance, numerous algorithms have been proposed, with a major focus on prompt design [30, 31, 23, 20, 42, 34, 53, 2, 45]. This paper proposes to explore an orthogonal direction, i.e., graph-learning-based approaches. In task planning, solvable sub-tasks can be naturally represented as a task graph, wherein each node corresponds to a distinct sub-task, and each edge signifies the dependencies between these sub-tasks. The crux of task planning, therefore, involves selecting a connected path or subgraph to satisfy the user’s request, which is a decision-making problem on graphs. Adopting this framework, we analyze the task planning capabilities of LLMs, specifically within the context of HuggingGPT [31]. Our empirical investigation uncovers that a considerable portion of planning failures can be ascribed to the LLMs’ inefficacy in accurately discerning the structure of the task graph. This finding presents intriguing questions from both theoretical and empirical perspectives. Theoretically, it initiates a discussion on the inherent limitations of LLMs in processing graph-based tasks. Empirically, it highlights the urgent need for developing effective and efficient strategies to mitigate this deficiency and improve task planning performance.
For the theoretical question, we first investigate the expressiveness of Transformer architectures when applied to graph tasks with sequential graph input, such as edge list representations. Our initial hypothesis is that the format of sequential graph input might not align with the inductive bias inherent to graph structures, potentially reducing expressiveness. Contrary to this hypothesis, it is proved that by taking edge lists as the input, a constant-width Transformer can solve graph decision-making problems by simulating dynamic programming algorithms on edge lists. Nevertheless, we find that LLMs’ solutions lack invariance under graph isomorphism, an important property for graph decision-making problems. In addition, the expressiveness is weakened if the attention is sparse, which is typically observed in LLMs [48]. Beyond expressiveness, we also examine the influence of auto-regressive loss, demonstrating that it introduces spurious correlations that can be harmful to graph decision-making tasks. These insights expose the inherent limitations of LLMs in task planning and, more broadly, in graph-related problems (e.g., the challenges in [10, 41, 24]).
To tackle the limitations, we take the use of GNNs, which have been shown to adeptly handle graph decision-making problems, both in theory and in practice [16, 50]. Initially, we deploy LLMs to interpret an ambiguous user request, breaking it down into more detailed steps. Subsequently, we utilize a GNN to retrieve the relevant sub-tasks based on these detailed steps and the corresponding sub-task descriptions. Notably, this approach can be implemented without training if we adopt parameter-free GNN models such as SGC [47]. In the case of training-based methods, we apply the Bayesian Personalized Ranking (BPR) loss [27] to facilitate learning from the implicit sub-task rankings. Extensive experiments demonstrate that the proposed methods achieve better performance than baselines. Specifically, our main contributions are summarized as follows:
  1. 1.
    Task Planning Formulation: This study presents a formulation of task planning as a graph decision-making problem. In the realm of task planning, our work initiates the exploration of graph learning methodologies to enhance performance. Concurrently, it introduces task planning as a new application in the graph learning domain.
  2. 2.
    Theoretical Insights: We prove that Transformers have expressiveness to solve graph decision-making problems based on edge list input, but inductive biases of attention and the auto-regressive loss function may serve as obstacles to their full potential.
  3. 3.
    Novel Algorithms: Based on the theoretical analysis, we introduce an additional GNN for sub-task retrieval, available in both training-free and training-based variants. The experiments on diverse LLMs and planning datasets demonstrate that the proposed method outperforms existing solutions with much less computation time. Furthermore, the performance is further enhanced by improved prompts or a fine-tuned model.
2 Preliminaries and Formulations
In this section, we introduce the task planning and GNNs. More importantly, we connect them with the classic concepts, i.e., decision-making problems and dynamic programming. This connection will play a pivotal role in the theoretical analysis and algorithm development.
2.1 Task Planning and Decision Making
In task planning, the input comprises an ambiguous user request along with a pool of available sub-tasks. The output is a selection of specific sub-tasks to be invoked and the manner in which they are to be executed. An illustrative example within HuggingGPT is depicted in Figure 1.
The sub-tasks form a task graph $G=(V,E)$ [30], where the nodes $V=\{v_{1},\cdots,v_{n}\}$ represent a set of solvable tasks, and an edge from $v_{i}$ to $v_{j}$ is established if the sub-task $v_{j}$ is dependent on the sub-task $v_{i}$. The goal of task planning is to identify a path or a connected sub-graph within the task graph that fulfills the users’ requests. In the context of HuggingGPT [31], the nodes correspond to expert AI models from the HuggingFace website. An edge is drawn from $v_{i}$ to $v_{j}$ if the input format of $v_{j}$ matches the output format of $v_{i}$. Task planning involves invoking a sequence of AI models (i.e., a path). Viewed from this angle, task planning bears resemblance to traditional decision-making problems on graphs, such as planning for the shortest path. The key distinction from classic decision-making problems lies in the specification of the objective value: it is explicitly defined in the former, while in task planning, it is implicitly inferred from the users’ requests.
Figure 1: Illustration of Task Planning in HuggingGPT.
2.2 GNNs and Dynamic Programming
GNNs are a class of neural networks preserving equivariance under graph isomorphism. The most popular GNNs are based on message passing scheme, where the node representation is iteratively updated by aggregating the representations from its neighbors. Formally, the $k$-th layer of a GNN is given by
$\bm{m}_{u}^{(k)}=\text{Aggregate}^{(k)}(\{[\bm{h}_{v}^{(k-1)},\bm{e}_{u,v}]|v\in\mathcal{N}(u)\}),\quad\bm{h}_{u}^{(k)}=\text{Combine}^{(k)}(\bm{h}_{u}^{(k-1)},\bm{m}^{(k)}_{u}),$ (1)
where $\mathcal{N}(u)$ denotes the set of in-neighbors of node $u$, $\bm{e}_{u,v}$ is the edge feature, and $\bm{h}_{u}^{(k)}$ is the representation of node $u$ with $\bm{h}_{u}^{(0)}$ representing initial node features. The aggregation function $\text{Aggregate}^{(k)}$ and update function $\text{Combine}^{(k)}$ are parameterized by neural networks and the choice of these two functions leads to different types of GNNs.
DP [1] presents a general framework for solving decision-making problems. It breaks a complex problem into simpler sub-tasks, each of which can be solved based on the results of previous sub-tasks. The general form of DP is expressed as
$\text{Answer}[k][i]=f\left(\square_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]\oplus c[i][j])\right),$ (2)
where $\text{Answer}[k][i]$ is the solution to state $i$ in the $k$-th iteration, $f$ is a task-specific update function, $c[i][j]$ is a cost associated with state $i$ and $j$, $\mathcal{T}(i)$ is the set of state can be transited to $i$, $\square_{j\in\mathcal{T}(i)}$ is an aggregation function such as MAX or SUM, $\oplus$ is an operator such as + or XOR. We give the formulation of typical DP algorithms in Appendix C.1. DP updates in (2) can be expressed by GNNs’ updates in Eq. (1) if we set the states as the nodes, $\mathcal{N}(u)=\mathcal{T}(u)$, and $\bm{e}_{u,v}=c[u][v]$. More formal relationships are established in [50, 7], showing that GNNs are well-suited for solving DP problems.
Figure 2: Illustration of performance and hallucination ratio across LLMs.
3 Diagnosing Task Planning in LLMs
3.1 Empirical Observations
We adopt the experimental settings as outlined in the work of HuggingGPT [31], where the prompts are specifically optimized for task planning on HuggingFace. The evaluation metric calculates the F1 score to assess the accuracy of the tasks identified by LLMs against the ground-truth tasks. Additionally, we report two task-graph-related metrics: the node hallucination ratio and the edge hallucination ratio. These metrics measure the frequency of non-existent nodes (i.e., tasks) and edges (i.e., relations) outputted by LLMs, respectively, indicative of the models’ misinterpretation of the graph input.
Our empirical findings reveal that (1) LLMs exhibit a certain hallucination ratio, and (2) there is a strong correlation between the hallucination ratio and planning performance. This suggests that LLMs struggle to accurately interpret the task graph while the task graph is the key to the performance.
3.2 Theoretical Insights
Besides the example mentioned above, existing studies demonstrate that LLMs struggle with understanding graph-structured inputs [41, 10]. Despite these empirical evaluations, there remains a limited theoretical understanding of this phenomenon. In this subsection, we explore the performance of LLMs in graph decision-making problems, including task planning. In contrast to previous graph learning approaches for graph decision-making problems, LLMs process the graph input by flattening it into a sequence and are trained using an auto-regressive loss. We will then examine the impact of these two factors.
Sequential Graph Input: We consider general graph decision-making problems that can be resolved using DP as described in (2). The input comprises the edge list and initial states:
$\underbrace{u_{1}\ v_{1}\ c[u_{1}][v_{1}]\ u_{1}\ v_{2}\ c[u_{1}][v_{2}]\ \ldots}_{\text{edge list}}\ \underbrace{u_{1}\ \text{Answer}[0][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[0][u_{n}]}_{\text{initial states}}$ (3)
The intended output format is $u_{1}\ \text{Answer}[k][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[k][u_{n}]$. To our surprise, although the edge list input does not directly reflect the geometric structures of graphs, it enables Transformers to simulate DP efficiently, in terms of expressiveness, as demonstrated by the following theorem.
Theorem 1.
(Expressiveness) Assume the input format is given in (3) and $f,\oplus,\square$ in DP update (2) satisfy the assumptions 1 and 2. There exists a log-precision constant-depth and constant-width Transformer that simulates one step of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
The proof is presented in Appendix B.1. However, certain aspects of the proof’s constructions are challenging to be realized in Transformers that have been pretrained on natural language. First, the embedding process must be carefully filtered to ensure invariance under graph isomorphism. This invariance property does not align with the inductive biases inherent in natural language, making it difficult to achieve. Consequently, if an LLM can accurately produce the correct answer for a specific ordering of nodes, it might not maintain this accuracy after the nodes have been reordered (experiments given in Appendix C.2). Second, each token needs to synchronize its hidden states with all other tokens sharing the same token ID, which is of order $O(|V|)$. In practice, the attention trained from natural language is typically sparse [48], leading to intractability issues. The formal lower bound is provided in the following proposition and the proof is given in Appendix B.2.
Proposition 1.
Assume the input format is described (3) and that the attention mechanism is limited to attending to a constant number of tokens. There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
Auto-regressive Loss: Our investigation next focuses on the auto-regressive loss and considers the following scenario: given a fixed task graph, user data is collected to perform instruction tuning with next-token-prediction loss. A question arises concerning whether auto-regressive loss might introduce spurious correlations harmful to planning problems. We conceptualize this issue through the lens of a path-planning problem. The training dataset comprises input sequences of the form $s\ t\ s\ v_{1}\ v_{2}\ \cdots\ t$, where $s$ represents the source node, $t$ the target node, and the sequence $s\ v_{1}\ v_{2}\ \cdots\ t$ is a path that adheres to specified constraints. During testing, given the initial and target nodes $s$ and $t$, the model is expected to generate a path with the same constraint. Our findings indicate that auto-regressive loss can lead to the emergence of a frequency-based spurious correlation, as substantiated by the following theorem and the proof is given in Appendix B.3.
Theorem 2.
(Spurious correlations of auto-regressive loss) Assume (1) the loss employed is a next-token-prediction loss utilizing cross-entropy, applied to the sub-sequence $v_{1}\ v_{2}\ \cdots\ t$ during training; (2) the output logits are determined by target node $t$ and the current node $v_{i-1}$. Let $N_{t,v_{i-1},u}$ be the number of times in the training dataset such that $t$ is the target node, $v_{i-1}$ is the current node and $v_{i}=u$ is the next node. The optimal logits for predicting the next node $u$ from current node $v_{i-1}$ towards target node $t$ is given by $\hat{\bm{v}}_{i}[u]=\frac{N_{t,v_{i-1},u}}{\sum_{u}N_{t,v_{i-1},u}}$ if $\sum_{u}N_{t,v_{i-1},u}>0$. If $\sum_{u}N_{t,v_{i-1},u}=0$, $\hat{\bm{v}}_{i}[u]$ can be any non-negative number subject to $\sum_{u}\hat{\bm{v}}_{i}[u]=1$.
In our setup, $s$ $t$ is the instruction and the third token is a duplicate of the first token. It is reasonable to exclude these tokens in the loss calculation, which is the first assumption. The second assumption assumes that the output only depends on the current node and target node, which is a minimal requirement for path-related problems. For DP problems, the frequency-based prediction contradicts to the value-based ground-truth. We then give an example that auto-regressive loss even cannot find a valid path.
Example 1.
Consider a training dataset consisting of a sufficient number of valid paths. Suppose the dataset contains two paths $a\ b\ c$ and $b\ c\ d$ and there are no other paths such that $t=d$ and the current node $v_{i}=a$ for all $i$. Then we have $N_{d,a,u}\equiv 0$ for all $u$ and the logits for the next node can be arbitrary. This results in the model’s inability to predict the next node of $a$ when given $a$ as the source node and $d$ as the target node.
For a human, finding a path from $a$ to $d$ simply involves concatenating the paths $a\ b\ c$ and $b\ c\ d$. However, auto-regressive loss fails under such circumstances. In task planning datasets, we indeed observe that the performance of finetuned LLMs is inferior to that of GNNs trained on the same dataset, as shown in Figure 3(b).
4 A Simple Fix: Introducing GNNs
In contrast to LLMs, GNNs can strictly operate on the task graph, thereby avoiding hallucinations. Additionally, they leverage the graph structure as input, rather than flattening the graph into a sequence, thus overcoming the theoretical limitations discussed previously. Furthermore, GNNs have demonstrated proficiency in handling graph decision-making problems, both theoretically and empirically [50, 7, 16]. As a result, the simplest fix is to integrate GNNs into the task-planning algorithm. In this section, we first briefly introduce the current LLM-based solutions and then illustrate how GNNs can be integrated into them. We propose both training-free and training-based approaches to enhance performance. Training-free methods are necessary when the available tasks are continuously changing, or new tasks are emerging consistently. This scenario is common when the task planning module is deployed in a new system. Once the task planning module has been deployed for a period, it becomes possible to collect users’ requests and label a small proportion of the data, enabling lightweight training-based methods.
4.1 Purely LLM-based Task Planning
We outline a two-stage pipeline for task planning adopted in HuggingGPT [30, 31]. The first stage involves request decomposition, where a user’s ambiguous request is broken down into concrete steps via LLMs. For instance, the request illustrated in Figure 1 is decomposed into the following steps: (1) analyze the pose of the boy; (2) take that pose and generate a new image; (3) generate the caption for newly generated image; (4) convert the generated text into audio. The second stage is task retrieval. For each decomposed step, LLMs are employed to retrieve an appropriate task from the task pool and execute them in sequence.
4.2 A Training-free GNN-based Approach
We introduce GNNs into the task retrieval stage by formulating it as a node classification problem. The node embedding is the embedding of this node’s description. For example, the node description in HuggingFace can be “Translation. Translation is the task of converting text from one language to another”. Given that the node description is usually simple, node embedding is obtained by e5-355M [44] for efficiency. Then we use the parameter-free GNN (i.e., SGC [47]) to refine the node embeddings. The update is given by $\bm{H}^{(k)}=\alpha\bm{H}^{(k-1)}+(1-\alpha)\hat{\bm{A}}\bm{H}^{(k-1)}$, where $\bm{H}^{(k)}$ is the embedding after $k$-layer SGC, $\bm{H}^{(0)}$ is the initial node feature, $\hat{\bm{A}}$ is the normalized adjacency matrix of the task graph, and $\alpha$ is a hyperparameter.
When selecting the task node for a step $s_{i}$, e5-355M is used to embed the LLMs’ generated step as $\bm{x}_{i}^{\text{step}}$. Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the node embedding. Particularly, $v_{1}$ can be selected from the whole graph. The one-by-one node selection approach is inspired by the traditional research of GNNs for decision-making problems like combinatorial optimization [16].
4.3 A Training-required GNN-based Approach
The inference process in training-required methods mirrors that of the training-free approach, with the difference being the substitution of parameter-free GNNs with parametric counterparts, such as GAT [39] or GraphSAGE [11]. Here we specify the training process of GNNs.
Data Preparation: We assume that each entry in the task planning dataset comprises a user request, a sequence of decomposed steps, and the corresponding ground-truth tasks, denoted as $(\text{request},\{s_{1},\ldots,s_{n}\},\{v_{1},\ldots,v_{n}\})$. If the dataset does not adhere to this format, we reformat it accordingly using GPT-4, with details provided in Appendix E.2. It is important to note that there is a one-to-one correspondence between the steps and tasks in the dataset. The node classification dataset is represented as $\{s_{i},v_{i}\}_{i=1}^{n}$, where $s_{i}$ is a step described in natural language, and $v_{i}$ is its corresponding task.
Training Loss: Inspired by GNN for recommendation [12], we adopt the Bayesian Personalized Ranking (BPR) loss [27] designed for recommendation with binary rankings. The loss function is given by
$\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle),$
where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is ground-truth task, and $v^{\prime}$ is a negative task. We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task. The trainable parameters may merely include GNNs or both GNNs and e5-355M.
5 Experiments and Analysis
5.1 Experimental Setup
Datasets: We utilize four datasets across two task planning benchmarks: HuggingFace tasks, Multimedia tasks, and Daily Life API tasks from TaskBench [30], as well as TMDB API tasks from RestBench [35]. The HuggingFace dataset includes AI models on the HuggingFace. The Multimedia dataset provides a wide range of user-centric tasks, such as file downloading and video editing. The Daily Life APIs cater to everyday services like web search and shopping functionalities. TMDB focuses on movie-related search and retrieval tasks. Statistics for each dataset are presented in Table 5 with illustrative examples shown in Figure 8 in Appendix. Other benchmarks, such as ToolBench [25] and ToolAlpaca [38], are less suitable for our experiments due to (1) the absence of a well-defined task graph detailing tasks and their dependencies, and (2) a scarcity of samples involving multi-task planning, with a focus on single-task retrieval.
Evaluation: For the datasets from TaskBench, we split $3000$ samples for training and $500$ samples for testing, each containing an invocation path with at least two tasks. For the TMDB dataset, we first filter to include the samples with two or more invoked tasks, and then randomly select a sample served as the in-context learning example. The remaining $94$ samples are designated for testing. We adopt the evaluation metric in TaskBench [30] and HuggingGPT [31], i.e., Node F1-Score (n-F1) and Link F1-Score (l-F1), which measure the accuracy of invoked tasks and invoked dependencies, respectively. We also measure the token consumption (# tok) as the efficiency metric.
Choices of LLMs: We consider close-sourced LLMs, i.e., GPT-3.5-turbo and GPT-4-turbo, as well as open-sourced LLMs with different parameter scales, including CodeLlama-13B(or 7B)-Instruct-hf [28], Mistral-7B-Instruct-v0.2 [15], Vicuna-13B-v1.5 [54], and Baichuan2-13B-Chat [51].
Choices of GNNs: To comprehensively investigate the effectiveness of different graph learning methods in enhancing task planning, we consider a wide range of graph neural networks, including SGC [47], GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32].
5.2 Performance of the Training-free Approach
We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 4.1. (2) GraphSearch [23, 36, 22] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score to the selection. Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores. (3) SGC [47]: Employ a training-free SGC for task retrieval based on decomposed task steps. Illustrations of these methods are shown in Figures 5 and 6 in Appendix. Table 1 shows both the overall performance and token consumption costs.
[…]
TaskBench RestBench
HuggingFace Multimedia Daily Life TMDB
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ #Tok $\downarrow$
Baichuan2 13B Direct 45.85 19.00 2.43 47.57 4.08 2.59 33.45 9.52 3.72 30.87 9.92 1.96
GreedySearch 30.58 4.89 6.42 18.74 4.45 5.69 15.60 1.61 5.91 22.52 2.98 3.62
AdaptiveSearch 39.30 10.41 10.81 33.24 9.22 11.17 34.39 12.73 16.71 30.33 10.00 8.71
BeamSearch 41.06 9.59 24.69 32.24 9.09 21.60 36.18 13.18 23.83 30.97 7.61 9.08
SGC 56.53 29.94 2.28 56.75 31.62 2.43 62.31 36.69 3.53 32.97 9.11 1.84
Vicuna 13B Direct 50.46 21.27 2.50 53.57 23.19 2.64 73.70 45.80 3.82 44.66 14.01 2.02
GreedySearch 52.94 25.73 6.23 46.99 23.11 5.55 42.98 13.33 7.18 45.22 13.69 3.42
AdaptiveSearch 54.36 25.67 9.81 51.24 24.32 11.25 62.71 31.15 13.92 41.32 7.02 6.51
BeamSearch 56.64 26.93 24.11 54.09 26.19 25.42 54.55 23.60 24.86 46.91 15.41 7.79
SGC 59.62 31.98 2.31 61.78 37.60 2.43 83.33 63.77 3.82 48.79 15.99 1.89
CodeLlama 7B Direct 58.06 29.39 2.44 59.44 30.83 2.57 84.12 62.89 3.82 65.67 41.99 1.94
GreedySearch 58.71 31.56 5.84 62.83 38.12 5.35 82.51 63.83 7.08 65.51 42.60 3.12
AdaptiveSearch 60.42 33.18 6.84 62.32 36.81 5.50 83.42 64.15 7.83 65.37 40.64 5.00
BeamSearch 60.34 31.36 17.95 64.12 38.99 21.48 83.25 63.48 24.48 64.60 40.50 5.78
SGC 63.98 39.27 2.30 67.04 45.04 2.43 87.73 70.49 3.59 66.15 42.62 1.88
Mistral 7B Direct 60.60 30.23 2.49 69.83 39.85 2.64 84.26 53.63 3.77 62.23 22.02 1.96
GreedySearch 65.91 38.13 6.52 58.92 34.72 6.26 75.18 49.47 8.27 60.64 23.18 4.38
AdaptiveSearch 67.30 38.90 7.68 71.59 44.84 10.66 86.39 63.65 10.92 54.04 21.35 9.99
BeamSearch 67.13 36.73 25.66 73.55 47.12 31.10 85.87 61.53 39.16 63.41 26.79 11.26
SGC 67.43 42.08 2.32 74.07 49.90 2.43 87.13 66.49 3.54 64.72 25.67 1.89
CodeLlama 13B Direct 57.55 28.88 2.45 68.57 41.79 2.59 91.20 76.07 3.88 68.91 43.74 2.02
GreedySearch 61.67 34.02 5.95 67.98 42.04 4.95 91.50 76.56 5.54 66.67 42.16 3.81
AdaptiveSearch 60.85 31.66 11.10 68.14 41.71 6.77 91.34 76.09 7.18 63.74 37.17 8.16
BeamSearch 62.65 34.31 20.14 69.53 43.35 19.51 91.74 76.60 19.19 68.08 42.92 8.88
SGC 65.51 39.44 2.31 73.32 53.28 2.43 92.96 79.57 3.64 71.40 47.55 1.90
GPT- 3.5-turbo Direct 73.85 45.73 2.14 82.85 62.07 2.26 96.09 83.65 3.36 81.70 57.52 1.67
GreedySearch 67.75 43.88 5.29 81.11 63.02 4.92 93.77 81.26 7.36 76.19 50.11 3.06
AdaptiveSearch 72.18 47.55 7.47 81.86 62.71 5.71 93.79 81.41 8.53 77.57 53.65 5.89
BeamSearch 75.51 49.62 14.22 83.57 64.50 12.91 95.66 82.72 22.05 81.24 57.98 6.42
SGC 76.37 50.04 2.02 83.65 63.65 2.09 96.38 86.19 3.16 82.63 59.15 1.61
GPT- 4-turbo Direct 77.60 52.18 2.19 88.29 69.38 2.28 97.36 84.58 3.37 82.56 56.67 1.75
GreedySearch 74.75 50.44 5.78 86.81 69.80 5.52 97.36 85.78 7.37 75.34 49.95 3.73
AdaptiveSearch 76.17 51.30 8.94 88.02 69.99 7.14 97.30 85.80 9.04 81.78 55.15 6.35
BeamSearch 77.56 52.54 8.98 88.16 70.39 6.90 97.35 85.78 8.99 80.11 51.00 5.18
SGC 77.79 52.20 2.03 88.54 69.83 2.10 97.35 85.76 3.16 82.27 56.37 1.62
Compared with direct inference, integrating SGC consistently improves performance, underscoring the effectiveness of the proposed method. GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes. Notably, our proposed method achieves comparable or superior performance to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 6). The case studies are provided in Appendix F.4. However, we observed only marginal improvements with GPT-4-turbo. A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets. In addition, the language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output. A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11.
5.3 Performance of the Training-based Approaches
Settings: We further explore the efficacy of training-based GNNs in three TaskBench datasets. The TMDB dataset is excluded due to its limited sample size. Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM, whose role is to generate node embeddings derived from task names and descriptions. Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 8 and Table 9.
Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method. However, the co-training of GNNs with e5-355M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance. Further analysis across a broad spectrum of GNNs (as shown in Table 8 and Table 9) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE. This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize.
Efficiency: The training cost is remarkably low because we use e5-355M [44] as the text embedding model for GNNs. If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes. Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-355M model. This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs. The details of the training time are given in Table 7.
Table 2: Performance of training-based GNN. The performance of all other GNNs are given in Table 8 and Table 9 in the Appendix.
HuggingFace Multimedia Daily Life
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
GraphSAGE (co-train) 59.76 35.59 57.97 33.29 63.21 38.10
Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
GraphSAGE (co-train) 62.82 37.04 65.89 42.18 84.23 65.44
CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
GraphSAGE (co-train) 67.19 42.94 70.00 48.28 87.81 70.20
Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
GraphSAGE (co-train) 67.61 43.14 76.96 55.46 87.61 66.75
CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
GraphSAGE (co-train) 68.92 44.85 76.28 55.41 93.30 79.51
GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
GraphSAGE (co-train) 77.87 53.04 85.51 66.56 96.34 86.09
GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
GraphSAGE (co-train) 78.49 52.62 88.86 70.25 97.42 85.80
(a) Different Prompts
(b) Fine-tuned LLMs
Figure 3: Orthogonal Effectiveness of GNN on either different prompt templates (on HuggingFace dataset) or fine-tuned LLMs.
5.4 Improved Prompts and Fine-tuned LLMs
In this subsection, we show that the proposed method is orthogonal to two dominant methods, i.e., the prompt engineering and finetuning.
Orthogonal to Improved Prompts: We investigate GNN’s effectiveness when applied to improved prompt templates, i.e., strategically designed prompts that enhance the task planning abilities of LLMs. Specifically, we consider two types of prompts: (1) In-context Learning with Increased Examples [30] During main experiments, we maintain the consistent 1-shot in-context learning example for LLM’s direct inference. To realize further improvements, we increase the number of examples to $2$, and results under this setting are denoted as “2-shot Prompt”; (2) Plan like a Graph (PlaG) [20] We adopt the prompt in [20] to encourage LLM to think and plan in a graph-like manner, we convert the entire task graph into plain text and then integrate PlaG instructions to enhance LLM’s task planning. Results under this prompt template are denoted as “PlaG Prompt”.
From the results shown in Figure 3(a), where we apply three different prompts to CodeLlama-13B and Mistral-7B on HuggingFace, it is clear that applying a GNN to improved prompts, where task steps are more concisely decomposed and predictions are more accurate, can also boost performance.
Orthogonal to LLMs’ Fine-tuning: To explore whether our framework maintains effectiveness on fine-tuned LLMs, which have acquired dataset-specific task planning capabilities, we conduct further experiments. For each dataset, we use LoRA [13] to fine-tune two LLMs of different parameter scales, including CodeLlama-7B and Vicuna-13B, based on the same training data as GNNs. Details of fine-tuning process are provided in Appendix D.4. The finetuned model is named as “FT-CodeLlama” and “FT-Vicuna” in Figure 3(b).
The results depicted in Figure 3(b) demonstrate that fine-tuning markedly enhances the task-planning capabilities of LLMs. Furthermore, applying GraphSAGE to the decomposed tasks of LLMs further improves the accuracy of the task planning.
6 Related Works
Due to space limitations, this section focuses on related works of task planning, while discussions of other relevant areas are deferred to Appendix A. The existing studies of task planning approaches can be categorized into several directions, including task decomposition, multi-plan selection, the use of external planners, reflection, and memory-aided planning [14]. Task decomposition methods, such as the chain-of-thought approach [45], employ the divide-and-conquer strategy, utilizing LLMs for both task decomposition and sub-task planning. The application of this method to task planning is detailed in Section 4.1 and is referred to as “Direct” in the baseline comparison. Multi-plan selection strategies, exemplified by the tree-of-thought [53] and graph-of-thought [2], leverage search-based methods to generate plans. Subsequently, LLMs evaluate these plans to select the most effective one. The GraphSearch methods used in our baselines fall into this category. External planner approaches [21] formalize tasks using the Planning Domain Definition Language (PDDL) and employ classic solvers to address the planning problem, which is closely related to the proposed approaches in this paper. The task planning considered here is more flexible and cannot be directly translated into formal languages. However, we demonstrate that GNNs can serve as an effective external planner. Reflection-based methods [33] focus on reflecting upon experiences to refine the plan, while memory-aided planning approaches [55] utilize external experiences, such as those from search engines. These approaches are deployed in interactive environments and orthogonal to this paper.
7 Conclusions
This paper presents an initial exploration into graph-learning-based approaches for task planning. Through theoretical analysis, we demonstrate the inductive bias of the attention mechanism and the utility of auto-regressive loss impedes their effectiveness in task planning. We propose to integrate GNNs for task graph analysis, which yields performance improvements across a range of LLMs and datasets.
Limitations: Despite the encouraging performance, there are limitations that highlight significant opportunities for enhancement. Firstly, our proposed method, while effective, is straightforward; more sophisticated GNN-based decision-making algorithms could potentially offer further improvements. Secondly, in our current framework, GNNs function as an external module with limited interaction with LLMs. Exploring the synergies between GNNs and LLMs, particularly by incorporating GNN outputs as tokens within LLMs, presents an interesting avenue for research. Thirdly, the construction of the task graph currently requires manual effort. Investigating automated graph generation techniques for this application is another promising direction for future work.
Appendix A More Related Works
A.1 LLMs for Graphs
With the breakthroughs in LLMs, there has been a surge of interest in applying LLMs to graph-related problems. GPT4Graph [10] and NLGraph [41] are two prominent benchmarks designed to evaluate the performance of LLMs in the context of graph tasks. They encompass a wide spectrum of challenges, various input formats, and state-of-the-art prompting techniques, demonstrating that LLMs possess basic graph processing capabilities. Importantly, the choice of prompts and formats significantly influences performance. However, these benchmarks also expose the models’ susceptibility to spurious correlations within graphs. For instance, GPT-4 achieves only about $50\%$ accuracy on shortest-path tasks, even with the use of complex prompts. GraphInstruct [24] attempts to fine-tune LLMs on graph-theory-related tasks, resulting in improved performance, though it remains far from satisfactory. Despite these empirical efforts, there is a limited theoretical understanding of these evaluation results. The analysis in Section 3.2 aims to shed light on the empirical observations reported in these studies.
Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [37] and GraphLLM [6]. These approaches have demonstrated significant improvements in performance on GNN-related tasks. However, they have not yet been applied to task planning due to the lack of extensive training data. A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [37], and applying them to task planning.
A.2 Theoretical Analysis of Reasoning
Reasoning is closely related to task planning and decision-making. The theoretical exploration of the reasoning abilities of neural networks was initiated by [50]. This work unifies various reasoning tasks, such as intuitive physics, visual question answering, and shortest path calculations, into DP problems. It then analyzes the generalization capabilities of MLPs, DeepSets, and GNNs. It is demonstrated that GNNs exhibit the best generalization bounds, attributed to their architecture’s resemblance to the Bellman-Ford algorithm, which is adept at solving DP problems. In terms of reasoning abilities within LLMs, [8] examines how the Chain of Thought (CoT) approach aids in solving arithmetic and DP problems without graphs. By decomposing challenging problems into simpler subproblems, CoT extends the expressive capabilities of Transformers from $\text{TC}^{0}$ to P. This analysis is further applied to linear and sparse Transformers in [52].
Our proof of Theorem 1 builds upon the proof of Theorem 4.7 in [8]. However, while [8] addresses DP problems without graph structures, Theorem 1 specifically focuses on DP problems with graph edge list inputs. Moreover, unlike [8], which decomposes and solves the DP problem sequentially, Theorem 1 proposes a method to simulate DP on edge lists in parallel. In addition, we analyze the negative results rising from the inductive bias of attention mechanism and auto-regressive loss. These theoretical contributions are novel and promise to be valuable for general reasoning and planning tasks.
A.3 GNNs for Decision-making
GNNs are popular approaches for solving decision-making problems on graphs. The problems investigated are typically NP-hard, such as the minimum vertex cover, maximum cut, and the traveling salesman problem [5]. The basic approach involves selecting nodes one by one in a manner that satisfies the constraints [16]. In this paper, we adopt this method to sequentially select task nodes. Furthermore, reinforcement learning can be used to enhance the performance of GNNs beyond what is achievable with supervised labels alone [16]. In [16], the node with the highest score is selected exclusively. Conversely, [19] employs beam search to improve performance by selecting the top-$k$ nodes in a single iteration. Additionally, GNNs are utilized as the method for variable selection in exhaustive searches for exact solutions to combinatorial optimization problems [9]. This paper conceptualizes task planning as a graph-based decision-making problem. Both greedy and beam search algorithms have been adopted in task planning [23, 22]. Given this connection, a promising future direction involves repurposing GNNs for decision-making approaches in task-planning applications.
Appendix B Proofs
B.1 Proof of Theorem 1
Assumption 1.
Each function $f,\oplus$ in (2) can be approximated by constant size MLP.
Assumption 2.
The aggregation function $\square$ in (2) is one of min, max, sum, mean.
The first assumption is mild as MLPs are universal approximators. The second assumption is mild because these are the most commonly used aggregation functions.
Theorem 3.
(Expressiveness) Assume the input format is given in (3) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2. There exists a log-precision constant-depth and constant-width Transformer that simulates $k$ steps of DP update in (2). As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
Proof.
Token Embedding and Positional Embedding: The three-dimensional token embedding includes the token type $e^{\text{type1}}$ (00 for answer; $1$ for node; $2$ for edge cost), refined token type $e^{\text{type2}}$ (00 for answer, $1$ for the node tokens in initial states, $2$ for the target node tokens in the edge list, $3$ for the source node tokens in the edge list, $4$ for edge cost), and the token id $e^{\text{token}}$ (from 00 to $|V|-1$). The two-dimensional positional embedding includes the embedding for initial state tokens $e^{\text{pos1}}$ (00 for edge list tokens, $1$ for the first two elements of initial states, $2$ for the second two elements of initial states, etc.), embedding for edge list tokens $e^{\text{pos2}}$ (00 for initial state tokens, $1$ for the first three elements of the edge list, $2$ for the second three elements of the edge list, etc.). There are also constant-dimensional placeholders to put the states of DP.
Block 1 - Initial State Broadcast: The goal of the first block is to broadcast the initial states from the initial state token to node tokens. (1) Use MLPs to recover the digits of the answer tokens and put them in the first placeholder if $e^{type}_{k}==0$; (2) Copy the first placeholder from answer token to its previous node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos1}}_{k}-e^{\text{pos1}}_{j})^{2}<\delta\}$; (3) Broadcast the first placeholder with MEAN in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{type1}}_{k}-e^{\text{type1}}_{j})^{2}+(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}]]$.
Block 2 - Edge Feature Operations: The goal of the second block is to copy the edge features from the edge feature token to the corresponding node tokens. (1) Use MLPs to recover the digits of the edge feature tokens and put them in the second placeholder if $e^{\text{type1}}==2$; (2) Copy the second placeholder from the edge feature token to the node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos2}}_{k}-e^{\text{pos2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}]]$.
Block 3 - Message Preparation: (1) Use MLPs to compute $\oplus$ and place the results in the third placeholder. Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}],\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$; (2) Use MLPs to clean up the first and second placeholder. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos1}},\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$
Block 4 - Message Passing: The goal of the fourth block is to compute $\square$ and $f$. (1) Use one or two attention heads (one for max, min, mean aggregations, and two for sum aggregations) to perform the aggregation operation. This is achieved by using MEAN or MAX or SUM for the first placeholder and setting $\mathcal{S}_{k}=\{j|(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}+(e^{\text{type2}}_{k}-e^{\text{type2}}_{j})^{2}<\delta\}$. Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\square_{v_{j}\in\mathcal{T}(u_{i})}\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]]$; (2) Use MLPs to compute $f$. Now the state for every node token is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},f(\square_{v_{j}\in\mathcal{T}(i)}(\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]))]$.
After four blocks, the final state for every node token $u_{i}$ is given by $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[1][u_{i}]]$. $\text{Answer}[k][u_{i}]$ can be obtained by repeating the above four blocks $k$ times. ∎
Lemma 1.
[8] Let $n\in\mathbb{N}$ be an integer and $\bm{x}_{1},\cdots,\bm{x}_{n}$ be a sequence of vectors where $\bm{x}_{i}=(\tilde{\bm{x}}_{i},r_{i},1)\in[-M,M]^{d+2}$ where $M$ is a large constant. Let $\bm{K},\bm{Q},\bm{V}\in\mathbb{R}^{d^{\prime}\times(d+2)}$ be any matrices with $\|\bm{V}\|_{\infty}\leq 1$ and let $0<\rho,\delta<M$ be any real numbers. Denote $\bm{q}_{i}=\bm{Q}\bm{x}_{i},\bm{k}_{j}=\bm{K}\bm{x}_{i},\bm{v}_{j}=\bm{V}\bm{x}_{j}$. Define a matching set $\mathcal{S}=\{j||\bm{q}_{i}^{T}\bm{k}_{j}|\leq\rho\}$. Define two following operations
  • COPY: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$ with $\bm{u}_{i}=\bm{v}_{\text{pos}(i)}$, where $\text{pos}(i)=\arg\max_{j\in\mathcal{S}_{i}}r_{j}$.
  • MEAN, MAX, SUM: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$, where $\bm{u}_{i}=\square_{j\in\mathcal{S}_{i}}\bm{v}_{j}$ and $\square$ is min or max or sum or mean.
Specifically, for any sequence of vectors $\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}$, denote the corresponding output of the attention layer as $\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}$. Then, we have $\|\bm{u}_{i}-\bm{o}_{i}\|_{\infty}\leq\epsilon$ for all $i\in[n]$ and $\mathcal{S}\neq\emptyset$.
B.2 Proof of Proposition 2
Proposition 2.
Assume the input format is as described in Equation (3) and that the attention mechanism is limited to attending to a constant number of tokens. There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
Proof.
We present a proof by contradiction. Assume that a token in a Transformer with a constant depth, constant width, and log-precision can attend to only a constant number of nodes. Under this assumption, the total information accessible to the token in such a Transformer architecture amounts to $O(\log n)$ bits. However, for a graph with $|V|$ nodes, the number of possible outcomes from executing one-step DP is $O(e^{|V|})$, necessitating $\Theta(|V|)$ bits for representation. By the pigeonhole principle, this scenario inevitably leads to at least two distinct DP outcomes being represented by the same output sequence generated by the model, thereby constituting a contradiction. ∎
B.3 Proof of Theorem 2
Theorem 4.
(Spurious correlations of auto-regressive loss) Assume (1) the loss employed is a next-token-prediction loss utilizing cross-entropy, applied to the sub-sequence $v_{1}\ v_{2}\ \cdots\ t$ during training; (2) the output logits are determined by target node $t$ and the current node $v_{i-1}$. Let $N_{t,v_{i-1},u}$ be the number of times in the training dataset such that $t$ is the target node, $v_{i-1}$ is the current node and $v_{i}=u$ is the next node. The optimal logits for predicting the next node $u$ from current node $v_{i-1}$ towards target node $t$ is given by $\hat{\bm{v}}_{i}[u]=\frac{N_{t,v_{i-1},u}}{\sum_{u}N_{t,v_{i-1},u}}$ if $\sum_{u}N_{t,v_{i-1},u}>0$. If $\sum_{u}N_{t,v_{i-1},u}=0$, $\hat{\bm{v}}_{i}[u]$ can be any non-negative number subject to $\sum_{u}\hat{\bm{v}}_{i}[u]=1$.
Proof.
We denote $\mathcal{D}$ as the training dataset, $L_{i}$ as the sequence length of the $i$-th sequence in the dataset, $\bm{v}_{i,j}$ as the one hot embedding of the $j$-th token in the $i$-th training sequence, and $\hat{\bm{v}}_{i,j,u}$ as the $u$-th logit at the $j$-th token in the $i$-th sequence. The cross-entropy loss is given by
$-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\bm{v}_{i,j,u}\log\hat{\bm{v}}_{i,j,u}=-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\mathbb{I}_{u=v_{i,j}}\log\hat{\bm{v}}_{i,j,u}\overset{(a)}{=}-\sum_{t,v_{j-1},u}N_{t,v_{j-1},u}\log\hat{\bm{v}}_{i,j,u}$
=$\overset{(b)}{=}$ $-\sum_{t,v_{j-1},u}\left(\sum_{u}N_{t,v_{j-1},u}\right)\left[\left(\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}\right)\log\hat{\bm{v}}_{i,j,u}\right],$
where (a) uses the assumption that the output logits are determined by target node $t$ and the current node $v_{i-1}$. In (b), we assume that $\sum_{u}N_{t,v_{i-1},u}\neq 0$. If $\sum_{u}N_{t,v_{i-1},u}=0$, then the corresponding logits will not affect the loss function and $\hat{\bm{v}}_{i,j,u}$ can take any number. The cross-entropy is minimized when $\hat{\bm{v}}_{i,j,u}=\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}$. ∎
Appendix C More Discussions
C.1 Examples of Dynamic Programming
Longest Increasing Subsequence: The Longest Increasing Subsequence (LIS) problem is a classic dynamic programming problem that involves finding the length of the longest subsequence within a given array arr where the elements are in strictly increasing order. The state transition function for the LIS problem can be expressed as:
$\text{Answer}[k][i]=\max_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+(\mathbb{I}(j\neq i)\times 1)),$
where $\mathcal{T}(i)=\{i\}\cup\{j\;|j<i\;\text{and}\;\text{arr}[j]<\text{arr}[i]\}$ denotes the set of states that can transfer to state $i$, the aggregation function $\square$ is implemented as $\max(\cdot)$, and the cost $c[i][j]$ is $1$ for those candidate states that are not equal to state $i$ as adding the element leads to a longer subsequence while 00 for the state itself.
Bellman-Ford Algorithm: The Bellman-Ford algorithm is also a classic dynamic programming algorithm used to find the shortest path from a single source vertex to all other vertices in a weighted graph. The core idea behind the Bellman-Ford algorithm is that the distance from the source vertex to a target vertex can be computed as the minimum distance from the source to any of the target’s neighboring vertices, plus the weight of the edge connecting the neighbor to the target. Therefore, the state transition function for the Bellman-Ford algorithm can be expressed as:
$\text{Answer}[k][i]=\min_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+w[j][i]),$
where $\text{Answer}[k][i]$ represents the length of the shortest path from the source vertex to node $i$ at the $k$-th iteration, $\mathcal{T}(i)=\mathcal{N}^{-}(i)$ denotes the set of in-neighbors of node $i$, and $w[j][i]$ denotes the weight from node $j$ to $i$. The aggregation function $\square$ is implemented as $\min(\cdot)$ as we try to find the shortest path.
Travelling Salesman Problem: This problem is defined as, given a set of cities and the distances between every pair of cities, finding the shortest possible route that visits every city exactly once and returns to the starting point. If we regard the set of already visited city $\mathcal{S}$ ending at $i$-th city as the current state, then states that can transfer to current state are those that ending city can reach $k$. Therefore, the state transition function for the TSP problem can be expressed as:
$\text{Answer}[k][i][\mathcal{S}]=\min_{j\in\mathcal{S},j\neq i}(\text{Answer}[k-1][j][\mathcal{S}\setminus\{i\}]+w[j][i]),$
where $\text{Answer}[k][i][\mathcal{S}]$ represents the cost of the shortest tour that visits all the cities in the set $\mathcal{S}$ and ends at the $i$-th city, the aggregation function is still implemented as $\min(\cdot)$ since we aim to find the shortest path, and $w[j][i]$ denotes the distance from city $j$ to city $i$.
C.2 Permutation Invariance Test of LLMs
Figure 4: Illustrative Examples of LLMs Failure to Solve Graph Problems under Node-ID Permutation. Experiments were repeated for 30 times.
Appendix D Implementation Details
D.1 Details of Prompt Templates
To facilitate reproducibility, we provide prompt templates for all LLM-related queries, including direct inference and GraphSearch in Tables 3 and 4, respectively. GraphSearch consists of two types of LLM queries: Task Assessment, where LLM evaluates the suitability of each candidate task to address current step’s requirement during the iterative search, and Path Selection, where LLM selects the optimal path to fulfill user’s request after the search.
Table 3: Prompt template for LLM’s direct inference [30]
# TASK LIST #
{{ task list }}
# GOAL #
Based on the above tasks, I want you to generate task steps and a task invocation graph (including nodes and edges) to address the # USER REQUEST #. The format must be in strict JSON format, like:
[…]
# REQUIREMENTS #
1. Generated task steps and task nodes can resolve the user request # USER REQUEST # perfectly. Task name must be selected from # TASK LIST #.
2. The task steps should strictly align with the task nodes, and the number of task steps should be same with the task nodes.
3. The task links should reflect the temporal and resource dependencies among task nodes, i.e., the order in which the tasks are invoked.
# EXAMPLE #
{{ in-context learning examples }}
# USER REQUEST #
{{ user request }}
Now, please generate your response in a strict JSON format: # RESULT #
Table 4: Prompt templates of GraphSearch [23]
Scenario Prompt
Task
Assessment
# CANDIDATE TASK LIST #
{{ candidate tasks }}
# GOAL #
[…]
# REQUIREMENTS #
1. The keys of the generated score dictionary must align with the provided candidate tasks, and you should output scores for all candidate tasks.
2. The “score” field denotes a concrete score that assesses whether each task can solve the given step’s demand. The score should be in the range of $[1,2,3,4,5]$, where a higher score indicates better task-solving and matching abilities.
3. Carefully consider the user’s intention in # STEP # to assign the score. If the # STEP # contains a candidate task, its score should be >= 3.
# EXAMPLE #
{{ in-context learning examples }}
# STEP #
{{ step description }}
Now please generate your result in a strict JSON format: # RESULT #
Path
Selection
# GOAL #
[…]
# REQUIREMENTS #
1. Carefully analyze both the user’s request and previously inferred task steps. Select the best solution that can perfectly follow the inferred steps and solve user’s request. Do not change their corresponding sequences.
2. Make sure that each task in the final solution list exists in the valid # TASK LIST # {{ task list }}.
# USER REQUEST #
{{ user request }}
# STEPS #
{{ steps }}
# SOLUTION LIST #
{{ list of searched solutions }}
Now please generate your result in a strict JSON format: # RESULT #
D.2 Implementation of Training-free Methods
In this subsection, we present the implementation details of two training-free baselines, including LLM’s direct inference and GraphSearch. Method illustrations are shown in Figure 5.
Figure 5: Illustration of LLM’s Direct Inference and GraphSearch Method.
Figure 6: Illustration of our GNN-enhanced Task Planning. We first prompt LLM to decompose complex and implicit user request into several manageable sub-tasks. We then use either a training-free or training-required GNN to sequentially align each step description with a suitable task, ultimately generating the invocation path.
LLM’s Direct Inference: During experiments, we uniformly apply 1-shot in context learning for LLM’s direct inference of task invocation path. For open-sourced LLMs, the temperature parameter is set to 0.2.
GraphSearch: This algorithm conducts an iterative search on the task graph to identify an optimal task invocation path that can best satisfy a given request. In each iteration, the neighbors of the last selected task are considered as candidates. These candidates are evaluated by LLM for their suitability for the current step. The search process follows a depth-first approach. After the task assessment in the final step, a set of potential invocation paths is generated. Subsequently, LLM is prompted to select the most appropriate path from these options. The GraphSearch algorithm is implemented in three distinct variants, each employing a unique task selection strategy:
  • GreedySearch consistently selects the task node with the highest score at each step. Although fast and simple, this approach can lead to cascading errors, resulting in degraded performance.
  • AdaptiveSearch selects tasks with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode. During experiments, we empirically set the score threshold to 3.
  • BeamSearch retains the top-$k$ tasks based on the LLM’s assessment scores within candidates. Beam search can expand the search space but slightly reduces the efficiency. We uniformly set the beam width to 2.
SGC: For the implementation of SGC, the parameter $\alpha$ is uniformly set to 0.7. Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC on HuggingFace dataset, Roberta-large-355M [26] serves as the text encoder. For all other datasets and LLMs, the e5-355M [44] configuration is employed.
D.3 Implementation of Training-based GNNs
LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M [44] as the LM backbone. For the graph encoder, our setup includes a single layer with a hidden dimension of 1024. During the model training, we set the batch size to 512 and run for 20 epochs with a learning rate of 1e-3. We use the Adam optimizer [17] and implement an early stopping mechanism with a patience of 5 epochs to prevent over-fitting. All experiments are conducted on a single NVIDIA A100-80G GPU.
Training Data Preparation: From each dataset in TaskBench, we randomly select $3,000$ samples to create the trainset. The original data includes specific task steps and corresponding task invocation paths. Therefore, we first employ a topological sort to align each task step accurately with corresponding task, forming “<step, ground-truth task>” pairs. Then, for each pair, we randomly sample 2 negative tasks to constitute the “<step, positive task, negative task>” triplets for model training. These negative samples are selected based on how textually similar they are to the positive one, creating a robust differentiation challenge for the model.
Choices of Different Configurations: Our training-based method offers two configuration options: training only the GNN while keeping the LM frozen, or co-training both the LM and GNN. The first configuration is designed to explore GNN’s capability in task retrieval. The latter leverages LM’s dataset-specific semantic embeddings to enhance performance. Illustrations of these configurations are provided in Figure 7. For the co-training setup, we use a learning rate of 2e-5 and a training duration of 10 epochs.
Figure 7: LM+GNN Configuration. We offer two configurations: only training the GNN while keeping the LM frozen (Table 8), or co-training both the LM and GNN (Table 9).
D.4 Fine-tuning LLMs
To explore the effectiveness of our proposed framework on fine-tuned LLMs, we employ Supervised Fine-Tuning (SFT) on the CodeLlama-7B and Vicuna-13B models. We use the same set of $3,000$ samples from GNN training as the LLMs’ fine-tuning data. In fine-tuning the LLM with LoRA [13], we set the lora_r parameter (dimension for LoRA update matrices) to 8 and the lora_alpha (scaling factor) to 16. The dropout ratio is set to 0.1, the batch size to 2, and we conduct training over 2 epochs with a learning rate of 1e-5. For HuggingFace dataset, the maximum input length is set to 800, while the maximum output length is 400. For Multimedia and Daily Life datasets, which contain a larger number of tasks and require longer textual inputs, we set the maximum input and output length to 1000 and 500, respectively. We utilize 2 NVIDIA A100-80G GPUs for fine-tuning the LLMs. Each dataset undergoes specific fine-tuning for both the CodeLlama-7B and Vicuna-13B models.
Appendix E Datasets
E.1 Overview
We provide the statistics of experimental datasets from two task planning benchmarks in Table 5. The illustrative examples from each dataset are shown in Figure 8. For datasets from TaskBench, each sample consists of original user request, corresponding decomposed task steps, and ground-truth task invocation path. As RestBench includes only user requests and corresponding API invocation sequences, we prompt GPT-4 to infer decomposed task steps aligned with each invoked API, thereby finalizing the dataset.
E.2 Reformatting Details of RestBench
The TMDB dataset from RestBench, focuses on movie-related searching and recommending functions. To align RestBench with our experiments, we have implemented the following processing steps: Reformatting original APIs by assigning unique task names and descriptions: APIs in RestBench were represented by request paths, such as “GET /movie/top_rated”, referring to the API that retrieves top-rated movies on TMDB. To enhance semantic differentiation among APIs, we first prompt GPT-4 to assign a unique name and a detailed functional description to each API. These names and descriptions were then manually verified and refined. For example, the API previously mentioned is renamed “Get Top-Rated Movies” with the description “This API retrieves a list of the highest-rated movies.” Note that though the original TMDB dataset contains 54 APIs, some were never invoked in any data examples. Therefore, we focus only on those APIs that appeared in at least one user request, resulting in a refined set of 46 APIs.
Constructing a Task Graph: Each API is regarded as a unique task node, and we depict their relationships from two aspects: (1) categorical association, and (2) resource dependencies. For instance, APIs that provide movie-related functionalities, such as retrieving movie details or recommending films, are grouped under the movie category. Conversely, APIs focused on person-related functions, like searching for actors, are classified under the person category. Additionally, if two APIs share a common parameter like move_id, we establish a link between them to indicate a resource dependency.
[…]
Table 5: Statistics of Experimental Datasets
TaskBench RestBench
Type Statistic HuggingFace Multimedia Daily Life TMDB
Task Graph # Nodes 23 40 40 46
# Links 225 449 1560 979
Link Type Resource Resource Temporal Resource / Category
All Data # Samples 7,546 5,584 4,320 100
Test Set # Samples 500 500 500 94
# Avg. Nodes 3.81 3.92 4.05 2.33
# Avg. Links 2.81 2.92 3.05 1.33
Figure 8: Illustrative examples of four experimental datasets.
Appendix F Supplementary Experiments
F.1 Full Results of Training-based GNNs
In Table 8, we present comprehensive results for all training-based GNNs, including GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32]. To highlight the improvements brought by GNNs, we also include results from the strongest baseline, BeamSearch.
From the results, it is obvious that all GNN encoders significantly enhance the task planning abilities. For instance, when applied to Vicuna-13B on HuggingFace dataset, the introduction of GCN results in a performance improvement of $17.83\%$, GAT contributes to a $17.48\%$ increase, GraphSAGE leads to a $22.59\%$ boost, and TransformerConv improves predictions by $18.0\%$. This conclusion can be generalized across various LLMs and datasets, demonstrating the robust capabilities of GNNs.
F.2 Performance of LM+GNN Co-trained Mode
In this subsection, we conduct a supplementary study where the parameters of a pre-trained LM are also tuned along with GNN during model training. The model configuration is illustrated in Figure 7, and the results are presented in Table 9.
The results demonstrate that, compared to the GNN-only tunable mode, co-training LM+GNN can lead to further performance improvements. This enhancement occurs because the language model acquires task-specific semantics, which makes the representations more discriminative and boosts the GNN’s effectiveness in task retrieval. Additionally, it is noted that under the co-training setup, the differences between various GNN encoders are relatively minor, with performance variations across GNNs for a specific LLM on any dataset remaining within $2\%$.
For LM+GNN co-trained mode, as the parameters of the LM backbone, e5-355M, are also trained, more computational resources are required. We list the number of trainable parameters and the required training times for this architecture in Table 7. In application scenarios, the choice between GNN-only and LM+GNN co-trained modes can be decided based on the availability of resources and performance requirements.
F.3 Full Efficiency Study
In this subsection, we present a comprehensive efficiency study on various experimental setups: inference times across different training-free modes, time required for training GNN or LM+GNN, and the resources needed for fine-tuning LLMs. Results are shown in Tables 6 and 7.
Table 6: Efficiency Study of Training-free Methods. Due to space constraints in the table, some LLMs are abbreviated such as “GPT-3.5” for “GPT-3.5-turbo”. For Vicuna-13B and Mistral-7B, some discrepancies in reported times may be attributed to variations in deployment across different A100 devices. SGC maintains efficiency comparable to LLM’s direct inference.
Inference Times Comparison on HuggingFace (Seconds)
Mode Baichuan Vicuna CodeLlama-7B Mistral CodeLlama-13B GPT-3.5 GPT-4
Direct 6.0 3.6 10.9 4.5 9.7 2.7 26.1
GreedySearch 30.7 45.7 23.1 109.8 29.1 7.4 55.6
AdaptiveSearch 50.2 79.4 27.2 28.2 52.4 9.4 87.0
BeamSearch 102.0 55.0 60.8 85.5 92.3 14.9 270.2
SGC 6.1 3.7 10.7 4.6 9.5 3.0 24.4
Inference Times Comparison on Multimedia (Seconds)
Direct 9.7 3.3 13.2 4.5 14.6 2.9 25.1
GreedySearch 52.3 54.3 37.0 109.7 9.9 8.8 52.2
AdaptiveSearch 98.1 142.5 25.7 41.6 25.6 9.6 84.2
BeamSearch 122.3 69.8 103.0 92.0 84.4 15.0 70.9
SGC 9.5 3.4 12.9 4.5 14.1 3.1 23.5
Inference Times Comparison on Daily Life (Seconds)
Direct 10.0 6.5 19.4 5.6 18.6 3.4 31.0
GreedySearch 62.7 49.8 29.3 198.3 37.6 13.2 124.3
AdaptiveSearch 133.3 97.6 30.5 69.1 45.5 16.5 209.0
BeamSearch 196.9 54.1 106.9 195.9 64.9 89.4 161.7
SGC 9.9 6.5 18.6 5.7 17.9 3.6 29.5
Table 7: Efficiency Study of Training-based Methods. We present total training times for both GNN and LM+GNN co-trained modes, and resources needed for fine-tuning LLMs.
Training GNNs
Time (Seconds)
Mode Configuration # Param HuggingFace Multimedia Daily Life
GNN-only GCN 1,049,600 136.5 136.1 237.7
GAT 1,051,648 136.9 151.8 237.8
GraphSAGE 2,098,176 134.2 149.8 233.8
GIN 2,099,200 134.2 134.9 233.6
TransformerConv 4,198,400 135.0 150.2 233.7
LM+GNN Co-trained LM+SGC 335,141,889 743.1 323.3 384.7
LM+GCN 336,191,488 482.8 567.8 384.2
LM+GAT 336,193,536 741.3 812.3 384.4
LM+GraphSAGE 337,240,064 741.2 406.8 381.7
LM+GIN 337,241,088 735.6 361.8 382.9
LM+TransformerConv 339,340,288 741.0 362.7 405.8
Fine-tuning LLMs
Device & Time (Hours)
LLM # Param (# Tunable Param) HugggingFace Multimedia Daily Life
CodeLlama-7B 6,742,740,992 (4,194,304) 2$\times$×A100 7.0 1$\times$×A100 13.8 2$\times$×A100 9.5
Vicuna-13B 13,022,417,920 (6,553,600) 1$\times$×A100 17.8 2$\times$×A100 10.3 2$\times$×A100 19.0
Efficiency of Training-free Modes: Open-sourced LLMs were deployed as local API services using the FastChat framework on a single A100-80G GPU. This configuration enables faster and parallel inference. Under this setup, LLM’s direct inference requires 3-15 seconds per request. GPT-3.5-turbo and GPT-4-turbo are accessed via API, with the latter generally requiring more time. GraphSearch requires several minutes to complete a request due to its exhaustive search on the task graph, impacting the efficiency. In contrast, SGC achieves comparable efficiency to LLM’s direct inference, as it requires only a single LLM query and both LM and SGC’s forward propagation processes are extremely efficient. Note that some discrepancies in reported times, such as Mistral-7B’s GreedySearch taking longer than other modes, may be attributed to variations in the deployment across different A100 GPUs.
Efficiency of Training GNNs: During experiments, each dataset shares the same GNN configuration: 1 single layer with a hidden dimension of 1024. Therefore, for each GNN, the number of parameters remains consistent across datasets. The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request. For LM+GNN co-trained mode, where e5-355M serves as the LM backbone, training times increase to approximately 6-12 minutes. In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
Efficiency of Fine-tuning LLMs: Fine-tuning a LLM with $3,000$ training samples over 2 epochs requires huge time, typically 10-20 hours on one or two A100-80G GPU devices.
F.4 Case Study
Effectiveness of GNN (Qualitative Result): We show two cases in Figure 9 where the results of LLM’s direct inference, BeamSearch, and GraphSAGE are compared. Due to space constraints and issues such as LLM’s output content decoding errors or invalid paths, we present only the first four valid paths searched by BeamSearch on the task graph. From the cases, we can conclude that BeamSearch relies on LLM’s inherent reasoning abilities. Although LLM can explore the ground-truth invocation path on the task graph, their final solutions are usually not the optimal as containing either the hallucination or wrongly invoked tasks due to limited instruction following and reasoning abilities. On the contrary, GNN can effectively align decomposed steps with suitable tasks, accurately hitting the ground-truth result and enhancing task planning.
Figure 9: Case Study of GNN’s Effectiveness. Nodes colored in pink and red denote wrongly predicted task or hallucinated task, respectively. Due to space limitations, we only show the first four valid searched paths of BeamSearch for illustration. Even though LLM can explore ground-truth paths during searching, they lack certain instruction-following and reasoning abilities to consistently choose the optimal path. On the contrary, GNN can correctly align decomposed steps with suitable tasks, hitting the ground-truth result.
Failure Cases of GNNs: We also present the failure cases where GNN performance deteriorates compared to direct inference to provide a comprehensive discussion of our method. Our conclusion is that the method is sensitive to the quality of decomposed task steps, as an ambiguous step may mislead GNN to wrongly select the task, and such errors can cascade due to the sequential selection of tasks on the task graph. As the case shown in Figure 10, step 2 is ambiguously described as “Segment the image and identify the tabular data”, which actually incorporates two distinct steps. This ambiguty causes GNN to choose the unsuitable Tabular Classification instead of the correct Image Segmentation. Since tasks are selected sequentially on the task graph, where the next task is a neighbor of the current selection, such an error can prevent the exploration of the next appropriate task, as it may not be a neighbor of the current, incorrect selection. We also present the BeamSearch explored paths and its final solution, where it hits the ground-truth result.
Figure 10: Failure cases of GNN. Our framework relies heavily on the quality of decomposed task steps. Ambiguous steps (step 2 actually incorporates two steps) may mislead GNN to select the wrong task.
Diagnosing GraphSearch with GPT-4-turbo: In our experiments, GraphSearch brings marginal or even decreased performance for GPT-4-turbo across most experimental datasets, contradicting current research [23, 22] which suggests that exhaustive search strategies can enhance the performance of more powerful LLMs. To provide a detailed analysis, we show three types of cases in Figure 11: Successful cases, Failure cases, as well as Maintaining cases:
  • Successful Cases: As shown in the figure, despite inaccuracies in task decomposition (GPT-4-turbo predicts an extra step compared to the three-step ground truth), its inherent reasoning abilities and the knowledge explored along the task graph can fix these mistakes, hitting the ground-truth.
  • Failure Cases: In the failure case, although GPT-4-turbo identifies the ground-truth solution during searching process, the final decision includes an incorrect task. This occurs because the final solution selection demands complex reasoning abilities of LLM, as the context contains long textual information from different aspects: the full task graph, user request, all searched paths, and related instructions (Prompt template is shown in Table 4). The demanding context and reasoning challenge exceed GPT-4’s capabilities, leading to errors.
  • Maintaining Cases: These occur when GraphSearch merely replicates the result of GPT-4-turbo’s direct inference, indicating no added benefit from exploring the task graph. In these instances, despite navigating the graph, GPT-4 fails to self-correct inaccuracies due to inherent reasoning limitations.
We emphasize that the results of GraphSearch with GPT-4-turbo, tend to “maintaining cases”. Under such conditions, even a few failures can degrade overall performance, explaining why GraphSearch does not consistently enhance performance for GPT-4-turbo.
Figure 11: Diagnosing GraphSearch with GPT-4-turbo. We provide three types of cases: (1) Successful Cases where GPT-4 can leverage inherent reasoning abilities to select the optimal invocation path. (2) Failure Cases where GPT-4 miss in the extremely long context, containing the whole task graph, all searched paths, and instructions, selecting an unsatisfactory path. (3) Maintaining Cases where the searched result is the same as direct inference result, and those wrongly predicted tasks can still not be refined even under exhaustive search.
Table 8: Performance of training-based GNNs. We also presents the results of BeamSeaerch, the strongest variant from GraphSearch method to provide a comprehensive comparison. All GNNs consistently enhance task planning performance across diverse LLMs, showing the effectiveness.
HuggingFace Tools Multimedia Tools Daily Life APIs
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
BeamSearch 41.06 9.59 32.24 9.09 36.18 13.18
GCN 57.67 31.47 55.51 30.16 62.11 37.05
GAT 57.74 31.87 54.95 29.24 62.11 37.05
GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
GIN 57.38 31.17 55.08 30.04 62.11 37.05
TransformerConv 59.15 34.36 56.06 31.12 64.52 40.14
Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
BeamSearch 56.64 26.93 54.09 26.19 54.55 23.60
GCN 59.46 33.14 62.48 38.89 83.05 62.95
GAT 59.28 33.39 62.96 39.24 83.05 62.95
GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
GIN 59.14 32.33 62.61 38.82 83.05 62.95
TransformerConv 59.57 33.47 63.32 39.38 85.41 66.28
CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
BeamSearch 60.34 31.36 64.12 38.99 83.25 63.48
GCN 65.07 40.50 67.46 45.84 87.23 69.27
GAT 65.20 40.93 67.41 46.46 87.23 69.27
GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
GIN 65.52 40.98 66.89 45.41 87.23 69.27
TransformerConv 65.83 42.58 68.90 47.20 88.36 71.72
Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
BeamSearch 67.13 36.73 73.55 47.12 85.87 61.53
GCN 66.54 40.74 73.34 50.76 86.39 65.49
GAT 66.77 40.74 73.36 50.20 86.39 65.49
GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
GIN 66.69 40.79 72.89 50.44 86.39 65.49
TransformerConv 68.26 43.08 73.80 51.45 88.25 67.84
Direct 57.55 28.88 68.57 41.79 91.20 76.07
BeamSearch 62.65 34.31 69.53 43.35 91.74 76.60
GCN 66.22 41.05 72.99 52.18 91.83 77.88
GAT 66.29 41.28 74.08 53.56 91.83 77.88
GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
GIN 66.40 40.89 73.62 53.15 91.83 77.88
CodeLlama-13B TransformerConv 66.70 42.07 74.72 54.10 93.81 80.44
GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
BeamSearch 75.51 49.62 83.57 64.50 95.66 82.72
GCN 76.93 51.43 84.92 65.05 96.38 86.15
GAT 75.63 49.36 84.77 65.48 96.38 86.15
GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
GIN 76.86 51.00 84.14 64.30 96.38 86.15
TransformerConv 77.61 52.30 84.21 64.32 96.38 86.19
GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
BeamSearch 77.56 52.54 88.16 70.39 97.35 85.78
GCN 77.01 50.49 88.56 69.60 97.10 85.22
GAT 76.41 49.66 88.43 69.52 97.10 85.22
GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
GIN 77.74 51.02 88.05 69.13 97.36 84.58
TransformerConv 78.47 52.17 88.07 68.71 97.32 85.57
Table 9: Performance of Training-based GNNs under the LM+GNN Co-trained Mode. Simultaneous training of LM and GNN yields significant performance improvements.
HuggingFace Tools Multimedia Tools Daily Life APIs
LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
SGC 60.97 36.12 56.02 31.36 64.84 40.00
GCN 60.68 36.31 57.82 32.87 64.73 38.92
GAT 60.39 35.37 57.24 32.19 64.46 40.14
GraphSAGE 59.76 35.59 57.97 33.29 63.21 38.10
GIN 60.31 35.82 56.62 31.53 63.55 38.19
TransformerConv 60.76 36.82 56.82 31.40 64.88 40.23
Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
SGC 64.40 38.97 65.12 41.63 84.74 65.90
GCN 62.06 35.49 65.02 40.63 85.22 66.93
GAT 63.06 36.97 64.58 40.30 85.63 67.11
GraphSAGE 62.82 37.04 65.89 42.18 84.23 65.44
GIN 62.09 35.33 64.44 40.67 85.31 66.83
TransformerConv 62.11 36.01 64.57 40.17 85.42 66.55
CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
SGC 67.47 43.58 69.61 48.24 87.98 70.63
GCN 67.03 43.24 69.33 47.60 87.88 70.40
GAT 67.12 42.96 68.62 46.17 88.59 71.64
GraphSAGE 67.19 42.94 70.00 48.28 87.81 70.20
GIN 66.62 42.34 69.00 47.72 88.45 71.53
TransformerConv 67.12 43.08 69.27 47.96 88.43 71.59
Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
SGC 69.04 44.22 76.09 54.91 87.58 66.70
GCN 67.72 43.02 76.79 54.90 87.87 67.13
GAT 67.54 43.56 76.26 53.94 87.86 67.30
GraphSAGE 67.61 43.14 76.96 55.46 87.61 66.75
GIN 68.95 43.97 76.47 54.95 87.75 67.07
TransformerConv 67.94 43.52 77.06 55.39 87.76 67.00
CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
SGC 70.14 45.20 75.65 55.45 93.45 79.89
GCN 69.39 45.18 76.03 55.22 93.38 79.74
GAT 69.68 45.57 75.24 54.99 94.06 80.96
GraphSAGE 68.92 44.85 76.28 55.41 93.30 79.51
GIN 69.01 44.76 74.72 53.91 94.24 81.23
TransformerConv 69.52 45.68 75.46 55.14 93.98 81.06
GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
SGC 77.87 52.86 85.95 66.95 96.39 86.16
GCN 77.72 52.58 85.84 66.92 96.33 86.06
GAT 77.49 52.30 85.81 66.97 96.38 86.15
GraphSAGE 77.87 53.04 85.51 66.56 96.34 86.09
GIN 77.73 52.36 85.63 66.69 96.38 86.19
TransformerConv 77.78 52.79 86.09 67.26 96.33 86.06
GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
SGC 78.44 52.84 89.09 70.52 97.38 85.85
GCN 78.33 52.75 89.00 70.24 97.34 85.67
GAT 78.37 52.43 88.99 70.48 97.32 85.56
GraphSAGE 78.49 52.62 88.86 70.25 97.42 85.80
GIN 78.45 53.07 88.74 69.84 97.42 85.80
TransformerConv 78.30 52.27 88.90 70.24 97.42 85.80

### v2

@@
-Can Graph Learning Improve Task Planning?
+Can Graph Learning Improve Planning in LLM-based Agents?
 Abstract
-Task planning is emerging as an important research topic alongside the development of large language models (LLMs).
-It aims to break down complex user requests into solvable sub-tasks, thereby fulfilling the original requests.
+Task planning in language agents is emerging as an important research topic alongside the development of large language models (LLMs).
+It aims to break down complex user requests in natural language into solvable sub-tasks, thereby fulfilling the original requests.
 In this context, the sub-tasks can be naturally viewed as a graph, where the nodes represent the sub-tasks, and the edges denote the dependencies among them.
@@
 Extensive experiments demonstrate that GNN-based methods surpass existing solutions even without training, and minimal training can further enhance their performance.
-Additionally, our approach complements prompt engineering and fine-tuning techniques, with performance further enhanced by improved prompts or a fine-tuned model 111 The code and datasets are available at https://github.com/WxxShirley/GNN4TaskPlan.
+The performance gain increases with a larger task graph size.
+111 The code and datasets are available at https://github.com/WxxShirley/GNN4TaskPlan
 1 Introduction
-LLM-empowered autonomous agents have recently emerged as a rapidly growing field of research and are considered a significant step towards artificial general intelligence (AGI) [43, 4].
-These agents have achieved remarkable successes across a variety of domains, as evidenced by their ability to address complex AI challenges (e.g., HuggingGPT [31] and Visual ChatGPT [46]), excel in gaming environments (e.g., Voyager [40]), and drive innovation in chemical research (e.g., [3]).
-Within this burgeoning field, task planning emerges as a critical area of study.
-It involves LLMs autonomously interpreting user instructions, breaking these instructions down into concrete and solvable sub-tasks, and then fulfilling the user’s request by executing each sub-task [29, 31, 30].
+LLM-based agents have recently emerged as a rapidly growing field of research and are considered a significant step towards artificial general intelligence (AGI) [60, 7].
+These agents have achieved remarkable successes across a variety of domains, as evidenced by their ability to address complex AI challenges (e.g., HuggingGPT [45]), excel in gaming environments (e.g., Voyager [57]), and drive innovation in chemical research (e.g., [5]).
+Within this burgeoning field, task planning in language agents emerges as a critical area of study.
+It involves LLMs autonomously interpreting user instructions, breaking user’s instructions in natural language into concrete and solvable sub-tasks, and then fulfilling the user’s request by executing each sub-task [42, 45, 44].
 For instance, in the case of HuggingGPT [45], task planning involves invoking expert AI models from the HuggingFace website to solve complex AI tasks beyond the capabilities of GPT alone.
@@
 This finding presents intriguing questions from both theoretical and empirical perspectives.
-Theoretically, it initiates a discussion on the inherent limitations of LLMs in processing graph-based tasks.
+Theoretically, it initiates a discussion on the inherent limitations of LLMs in processing task graphs.
 Empirically, it highlights the urgent need for developing effective and efficient strategies to mitigate this deficiency and improve task planning performance.
-For the theoretical question, we first investigate the expressiveness of Transformer architectures when applied to graph tasks with sequential graph input, such as edge list representations.
+For the theoretical question, we first investigate the expressiveness of Transformer architectures when applied to graph tasks with sequential graph input, such as edge list representations, which is the graph input format for task planning.
 Our initial hypothesis is that the format of sequential graph input might not align with the inductive bias inherent to graph structures, potentially reducing expressiveness.
@@
 Furthermore, the performance is further enhanced by improved prompts or a fine-tuned model.
-2 Preliminaries and Formulations
-In this section, we introduce the task planning and GNNs.
-More importantly, we connect them with the classic concepts, i.e., decision-making problems and dynamic programming.
-This connection will play a pivotal role in the theoretical analysis and algorithm development.
-2.1 Task Planning and Decision Making
-In task planning, the input comprises an ambiguous user request along with a pool of available sub-tasks.
-The output is a selection of specific sub-tasks to be invoked and the manner in which they are to be executed.
-An illustrative example within HuggingGPT is depicted in Figure 1.
-The sub-tasks form a task graph $G=(V,E)$ [30], where the nodes $V=\{v_{1},\cdots,v_{n}\}$ represent a set of solvable tasks, and an edge from $v_{i}$ to $v_{j}$ is established if the sub-task $v_{j}$ is dependent on the sub-task $v_{i}$.
-The goal of task planning is to identify a path or a connected sub-graph within the task graph that fulfills the users’ requests.
-In the context of HuggingGPT [31], the nodes correspond to expert AI models from the HuggingFace website.
-An edge is drawn from $v_{i}$ to $v_{j}$ if the input format of $v_{j}$ matches the output format of $v_{i}$.
-Task planning involves invoking a sequence of AI models (i.e., a path).
+2 Preliminaries
+In this section, we introduce task planning in language agents and the current LLM-based solutions.
+2.1 Task Planning in Language Agents
+We start with the definition of task planning with a concrete example of HuggingGPT [45].
+In task planning, there is a pool of pre-defined tasks.
+Task planning inputs include this task pool and a user request.
+The user request is expressed in natural language, which is ambiguous and could encompass multiple complex tasks.
+The output is a sequence of tasks and the order of their invocation to address the user’s request.
+[…]
+2.2 Current LLM-based Solution to Task Planning
+The current solution of task planning is purely based on LLMs and involves two stages [44, 45].
+The first stage involves request decomposition, where a user’s ambiguous request is broken down into concrete steps via LLMs.
+For instance, the request illustrated in Figure 1 is decomposed into the following steps: (1) analyze the pose of the boy; (2) take that pose and generate a new image; (3) generate the caption for newly generated image; (4) convert the generated text into audio.
+The second stage is task retrieval.
+For each decomposed step, LLMs are employed to retrieve an appropriate task from the task pool and execute them in sequence.
+For example, “Analyze the pose of the boy” corresponds to Pose detection.
+The output tasks should be (1) Pose detection; (2) Pose-to-Image; (3) Image-to-Text; (4) Text-to-Speech.
+Figure 6 illustrates this procedure.
+Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT)
+3 Graph Formulation and Insights
+3.1 Graph Formulation of Task Planning
+In this subsection, we formulate the task planning as a decision-making problem on the task graph.
+The task graph is a special kind of text-attributed graphs and we define it as $G=(V,E,T)$.
+Each node $v\in V$ represents a pre-defined task in the task pool, associated with a text $t_{v}\in T$ describing its function (e.g., “Translation.
+Translation is the task of converting text from one language to another.”).
+Each edge $(u,v)\in E$ indicates a dependency between tasks (e.g., the output format of task $u$ matches the input format of task $v$).
+Task planning is to select a path or connected sub-graph on the task graph.
 Viewed from this angle, task planning bears resemblance to traditional decision-making problems on graphs, such as planning for the shortest path.
-The key distinction from classic decision-making problems lies in the specification of the objective value: it is explicitly defined in the former, while in task planning, it is implicitly inferred from the users’ requests.
-Figure 1: Illustration of Task Planning in HuggingGPT.
-2.2 GNNs and Dynamic Programming
-GNNs are a class of neural networks preserving equivariance under graph isomorphism.
-The most popular GNNs are based on message passing scheme, where the node representation is iteratively updated by aggregating the representations from its neighbors.
-Formally, the $k$-th layer of a GNN is given by
-$\bm{m}_{u}^{(k)}=\text{Aggregate}^{(k)}(\{[\bm{h}_{v}^{(k-1)},\bm{e}_{u,v}]|v\in\mathcal{N}(u)\}),\quad\bm{h}_{u}^{(k)}=\text{Combine}^{(k)}(\bm{h}_{u}^{(k-1)},\bm{m}^{(k)}_{u}),$ (1)
-where $\mathcal{N}(u)$ denotes the set of in-neighbors of node $u$, $\bm{e}_{u,v}$ is the edge feature, and $\bm{h}_{u}^{(k)}$ is the representation of node $u$ with $\bm{h}_{u}^{(0)}$ representing initial node features.
-The aggregation function $\text{Aggregate}^{(k)}$ and update function $\text{Combine}^{(k)}$ are parameterized by neural networks and the choice of these two functions leads to different types of GNNs.
-DP [1] presents a general framework for solving decision-making problems.
-It breaks a complex problem into simpler sub-tasks, each of which can be solved based on the results of previous sub-tasks.
-The general form of DP is expressed as
-$\text{Answer}[k][i]=f\left(\square_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]\oplus c[i][j])\right),$ (2)
-where $\text{Answer}[k][i]$ is the solution to state $i$ in the $k$-th iteration, $f$ is a task-specific update function, $c[i][j]$ is a cost associated with state $i$ and $j$, $\mathcal{T}(i)$ is the set of state can be transited to $i$, $\square_{j\in\mathcal{T}(i)}$ is an aggregation function such as MAX or SUM, $\oplus$ is an operator such as + or XOR.
-We give the formulation of typical DP algorithms in Appendix C.1.
-DP updates in (2) can be expressed by GNNs’ updates in Eq.
-(1) if we set the states as the nodes, $\mathcal{N}(u)=\mathcal{T}(u)$, and $\bm{e}_{u,v}=c[u][v]$.
-More formal relationships are established in [50, 7], showing that GNNs are well-suited for solving DP problems.
-Figure 2: Illustration of performance and hallucination ratio across LLMs.
-3 Diagnosing Task Planning in LLMs
-3.1 Empirical Observations
+Compared with traditional planning, task planning in language agents involves diverse and open-ended goals due to the varied users’ personal requests.
+For example, on platforms like HuggingFace, users’ intentions span across video, text, and image domains.
+On the contrary, classic planning has a fixed goal for a given domain, which is often explicitly expressed by mathematical formulas [54, 35].
+(a) Performance and Hallucination ratios
+(b) Hallucination ratios across datasets
+Figure 2: Illustration of (a) LLMs’ planning performance and hallucination in HuggingGPT, and (b) hallucination in relation to task graph size.
+3.2 Failures of LLMs in Planning: Empirical Findings
+With the task graph at hand, we diagnose LLMs in task planning in Figure 2.
 We adopt the experimental settings as outlined in the work of HuggingGPT [45], where the prompts are specifically optimized for task planning on HuggingFace.
@@
 Additionally, we report two task-graph-related metrics: the node hallucination ratio and the edge hallucination ratio.
-These metrics measure the frequency of non-existent nodes (i.e., tasks) and edges (i.e., relations) outputted by LLMs, respectively, indicative of the models’ misinterpretation of the graph input.
+These metrics measure the frequency of non-existent nodes (i.e., tasks) and edges (i.e., dependencies) outputted by LLMs, respectively, indicative of the models’ misinterpretation of the graph input.
 Our empirical findings reveal that (1) LLMs exhibit a certain hallucination ratio, and (2) there is a strong correlation between the hallucination ratio and planning performance.
 This suggests that LLMs struggle to accurately interpret the task graph while the task graph is the key to the performance.
-3.2 Theoretical Insights
-Besides the example mentioned above, existing studies demonstrate that LLMs struggle with understanding graph-structured inputs [41, 10].
-Despite these empirical evaluations, there remains a limited theoretical understanding of this phenomenon.
-In this subsection, we explore the performance of LLMs in graph decision-making problems, including task planning.
+We further explore whether the incidence of hallucinations correlates with the number of sub-tasks.
+The HuggingGPT dataset contains 23 sub-tasks, and our analysis is expanded to incorporate the UltraTool dataset [19], which consists of 260 sub-tasks.
+Figure 2(b) illustrates that the hallucination ratio increases with a larger task graph size.
+3.3 Failures of LLMs in Planning: Theoretical Insights
+In this subsection, we provide theoretical insights into the limitations of LLMs in processing task graphs.
 In contrast to previous graph learning approaches for graph decision-making problems, LLMs process the graph input by flattening it into a sequence and are trained using an auto-regressive loss.
 We will then examine the impact of these two factors.
-Sequential Graph Input: We consider general graph decision-making problems that can be resolved using DP as described in (2). The input comprises the edge list and initial states:
+How does sequential graph input impact the expressiveness?
+We consider general graph decision-making problems that can be resolved using dynamic programming (DP) as described in (2).
+The input comprises the edge list and initial states:
 $\underbrace{u_{1}\ v_{1}\ c[u_{1}][v_{1}]\ u_{1}\ v_{2}\ c[u_{1}][v_{2}]\ \ldots}_{\text{edge list}}\ \underbrace{u_{1}\ \text{Answer}[0][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[0][u_{n}]}_{\text{initial states}}$ (1)
 The intended output format is $u_{1}\ \text{Answer}[k][u_{1}]\ \ldots\ u_{n}\ \text{Answer}[k][u_{n}]$.
+In existing studies, task graphs are often presented in (1), where the edge list is described by natural language and the initial states are the task descriptions of each task node, detailed in Appendix A.9 of [44].
+As discussed in the previous subsection, task planning is a decision-making problem on the task graph.
+The decision-making problems on graphs are often solved by DP [3] and its general formulation is given by
+$\text{Answer}[k][i]=f\left(\square_{j\in\mathcal{T}(i)}g(\text{Answer}[k-1][j],c[i][j])\right),$ (2)
+where $\text{Answer}[k][i]$ is the solution to state $i$ in the $k$-th iteration, $c[i][j]$ is a cost associated with state $i$ and $j$, $\mathcal{T}(i)$ is the set of state can be transited to $i$, $\square_{j\in\mathcal{T}(i)}$ is an aggregation function such as MAX or SUM, and $f,g$ are task-specific update functions.
+We give the formulation of typical DP algorithms in Appendix D.1 including some NPC problems.
+For the decision-making problems on the text-attributed graphs (e.g., task planning), one may conceptualize them as DP in the feature space, as discussed in [67].
 To our surprise, although the edge list input does not directly reflect the geometric structures of graphs, it enables Transformers to simulate DP efficiently, in terms of expressiveness, as demonstrated by the following theorem.
 Theorem 1.
-(Expressiveness) Assume the input format is given in (3) and $f,\oplus,\square$ in DP update (2) satisfy the assumptions 1 and 2.
+(LLMs have enough expressiveness) Assume the input format is given in (1) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2 in Appendix.
 There exists a log-precision constant-depth and constant-width Transformer that simulates one step of DP update in (2).
 As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
-The proof is presented in Appendix B.1.
+The proof is presented in Appendix D.2.
 However, certain aspects of the proof’s constructions are challenging to be realized in Transformers that have been pretrained on natural language.
@@
 This invariance property does not align with the inductive biases inherent in natural language, making it difficult to achieve.
-Consequently, if an LLM can accurately produce the correct answer for a specific ordering of nodes, it might not maintain this accuracy after the nodes have been reordered (experiments given in Appendix C.2).
+Consequently, if an LLM can accurately produce the correct answer for a specific ordering of nodes, it might not maintain this accuracy after the nodes have been reordered (experiments given in Appendix D.3).
 Second, each token needs to synchronize its hidden states with all other tokens sharing the same token ID, which is of order $O(|V|)$.
 In practice, the attention trained from natural language is typically sparse [65], leading to intractability issues.
-The formal lower bound is provided in the following proposition and the proof is given in Appendix B.2.
+The formal lower bound is provided in the following proposition and the proof is given in Appendix D.4.
 Proposition 1.
-Assume the input format is described (3) and that the attention mechanism is limited to attending to a constant number of tokens.
+(Inductive bias of language hinders expressiveness) Assume the input format is described (1) and that the attention mechanism is limited to attending to a constant number of tokens.
 There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
-Auto-regressive Loss: Our investigation next focuses on the auto-regressive loss and considers the following scenario: given a fixed task graph, user data is collected to perform instruction tuning with next-token-prediction loss.
-A question arises concerning whether auto-regressive loss might introduce spurious correlations harmful to planning problems.
-We conceptualize this issue through the lens of a path-planning problem.
-The training dataset comprises input sequences of the form $s\ t\ s\ v_{1}\ v_{2}\ \cdots\ t$, where $s$ represents the source node, $t$ the target node, and the sequence $s\ v_{1}\ v_{2}\ \cdots\ t$ is a path that adheres to specified constraints.
+How does auto-regressive Loss impact the generalization?
+Our investigation next focuses on the auto-regressive loss and considers the following scenario: given a fixed task graph, user data is collected to perform instruction tuning with next-token-prediction loss.
+For a tractable theoretical study, we conceptualize this issue as a path-planning problem, since the output of task planning is essentially a path.
+We consider the training dataset comprises input sequences of the form $s\ t\ s\ v_{1}\ v_{2}\ \cdots\ t$, where $s$ represents the source node, $t$ the target node, and the sequence $s\ v_{1}\ v_{2}\ \cdots\ t$ is a path that adheres to specified constraints.
 During testing, given the initial and target nodes $s$ and $t$, the model is expected to generate a path with the same constraint.
-Our findings indicate that auto-regressive loss can lead to the emergence of a frequency-based spurious correlation, as substantiated by the following theorem and the proof is given in Appendix B.3.
+Our findings indicate that auto-regressive loss can lead to the emergence of a frequency-based spurious correlation, as substantiated by the following theorem and the proof is given in Appendix D.5.
 Theorem 2.
@@
 This results in the model’s inability to predict the next node of $a$ when given $a$ as the source node and $d$ as the target node.
-For a human, finding a path from $a$ to $d$ simply involves concatenating the paths $a\ b\ c$ and $b\ c\ d$.
+To a human, finding a path from $a$ to $d$ simply involves concatenating the paths $a\ b\ c$ and $b\ c\ d$.
 However, auto-regressive loss fails under such circumstances.
-In task planning datasets, we indeed observe that the performance of finetuned LLMs is inferior to that of GNNs trained on the same dataset, as shown in Figure 3(b).
-4 A Simple Fix: Introducing GNNs
+In task planning datasets, we indeed observe that the performance of fine-tuned LLMs is inferior to that of GNNs trained on the same dataset, as shown in Figure 3(b).
+4 Integrating GNNs and LLMs for Planning
+4.1 Motivations
+In the last section, we find that a considerable portion of planning failures can be ascribed to the LLMs’ inefficacy in accurately discerning the structure of the task graph, due to the hallucination, the inductive bias of the attention, and next-token prediction loss.
 In contrast to LLMs, GNNs can strictly operate on the task graph, thereby avoiding hallucinations.
@@
 As a result, the simplest fix is to integrate GNNs into the task-planning algorithm.
-In this section, we first briefly introduce the current LLM-based solutions and then illustrate how GNNs can be integrated into them.
-We propose both training-free and training-based approaches to enhance performance.
-Training-free methods are necessary when the available tasks are continuously changing, or new tasks are emerging consistently.
+In the following subsections, we propose both training-free and training-based approaches to enhance performance.
+Training-free methods are necessary when the available tasks are continuously changing, or new tasks are emerging constantly.
 This scenario is common when the task planning module is deployed in a new system.
 Once the task planning module has been deployed for a period, it becomes possible to collect users’ requests and label a small proportion of the data, enabling lightweight training-based methods.
-4.1 Purely LLM-based Task Planning
-We outline a two-stage pipeline for task planning adopted in HuggingGPT [30, 31].
-The first stage involves request decomposition, where a user’s ambiguous request is broken down into concrete steps via LLMs.
-For instance, the request illustrated in Figure 1 is decomposed into the following steps: (1) analyze the pose of the boy; (2) take that pose and generate a new image; (3) generate the caption for newly generated image; (4) convert the generated text into audio.
-The second stage is task retrieval.
-For each decomposed step, LLMs are employed to retrieve an appropriate task from the task pool and execute them in sequence.
 4.2 A Training-free GNN-based Approach
-We introduce GNNs into the task retrieval stage by formulating it as a node classification problem.
-The node embedding is the embedding of this node’s description.
-For example, the node description in HuggingFace can be “Translation.
-Translation is the task of converting text from one language to another”.
-Given that the node description is usually simple, node embedding is obtained by e5-355M [44] for efficiency.
-Then we use the parameter-free GNN (i.e., SGC [47]) to refine the node embeddings.
-The update is given by $\bm{H}^{(k)}=\alpha\bm{H}^{(k-1)}+(1-\alpha)\hat{\bm{A}}\bm{H}^{(k-1)}$, where $\bm{H}^{(k)}$ is the embedding after $k$-layer SGC, $\bm{H}^{(0)}$ is the initial node feature, $\hat{\bm{A}}$ is the normalized adjacency matrix of the task graph, and $\alpha$ is a hyperparameter.
-When selecting the task node for a step $s_{i}$, e5-355M is used to embed the LLMs’ generated step as $\bm{x}_{i}^{\text{step}}$.
-Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the node embedding.
+As we discussed in Section 2.2, the current solution to task planning involves two stages.
+The first stage requires the ability to understand users’ requests in natural language and break them down into concrete instructions, which is the unique ability of LLMs.
+The second stage is to select a path on the task graph, where each node corresponds to a decomposed step.
+Thus, we can integrate GNNs in this stage.
+The illustration of our framework is shown in Figure 7 in Appendix.
+For each decomposed step outputted by the first stage, we use a GNN to select a corresponding node within the task graph.
+Suppose we are selecting the node for the $i$-th decomposed step.
+First, we utilize a small pre-trained language model, e5-355M [61], to embed the $i$-th decomposed step.
+The resulting embedding is denoted as $\bm{x}_{i}^{\text{step}}$.
+Second, for the task graph, we first use the same pre-trained language model e5-355M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node.
+Then we adopt a $K$-layer SGC [64] to compute the final node embeddings, resulting in $\bm{h}_{v}=\bm{h}_{v}^{(K)}$.
+Given a sequence of previously selected task nodes $\{v_{1},\cdots,v_{i-1}\}$, the next node $v_{i}$ is chosen according to $v_{i}=\text{argmax}_{v\in\mathcal{N}(v_{i-1})}\;\langle\bm{h}_{v},\bm{x}_{i}^{\text{step}}\rangle$, where $\bm{h}_{v}$ is the final node embedding, and $\mathcal{N}(v_{i-1})$ denotes the neighbors of node $v_{i-1}$ in the task graph.
 Particularly, $v_{1}$ can be selected from the whole graph.
-The one-by-one node selection approach is inspired by the traditional research of GNNs for decision-making problems like combinatorial optimization [16].
+As e5-355M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
 4.3 A Training-required GNN-based Approach
@@
 Data Preparation: We assume that each entry in the task planning dataset comprises a user request, a sequence of decomposed steps, and the corresponding ground-truth tasks, denoted as $(\text{request},\{s_{1},\ldots,s_{n}\},\{v_{1},\ldots,v_{n}\})$.
-If the dataset does not adhere to this format, we reformat it accordingly using GPT-4, with details provided in Appendix E.2.
+If the dataset does not adhere to this format, we reformat it accordingly using GPT-4, with details provided in Appendix C.2.
 It is important to note that there is a one-to-one correspondence between the steps and tasks in the dataset.
-The node classification dataset is represented as $\{s_{i},v_{i}\}_{i=1}^{n}$, where $s_{i}$ is a step described in natural language, and $v_{i}$ is its corresponding task.
-Training Loss: Inspired by GNN for recommendation [12], we adopt the Bayesian Personalized Ranking (BPR) loss [27] designed for recommendation with binary rankings. The loss function is given by
-$\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle),$
-where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is ground-truth task, and $v^{\prime}$ is a negative task.
+Therefore, the training dataset can be represented as $\{(s_{i},v_{i})\}_{i=1}^{n}$, where $s_{i}$ is a step described in natural language, and $v_{i}$ is its corresponding invoked task.
+Training Loss: The problem in the dataset can be viewed as a binary ranking problem, where the labeled node is $1$ and the other nodes are 00.
+Therefore, we adopt the Bayesian Personalized Ranking (BPR) loss [40] designed for recommendation with binary rankings.
+The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task.
 We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task.
-The trainable parameters may merely include GNNs or both GNNs and e5-355M.
+The trainable parameters may merely include GNNs or both GNNs and e5-355M with illustrative configurations shown in Figure 8 in Appendix.
 5 Experiments and Analysis
@@
 We adopt the evaluation metric in TaskBench [44] and HuggingGPT [45], i.e., Node F1-Score (n-F1) and Link F1-Score (l-F1), which measure the accuracy of invoked tasks and invoked dependencies, respectively.
+Besides, the Accuracy (Acc) can measure the success rate from task level.
 We also measure the token consumption (# tok) as the efficiency metric.
 Choices of LLMs: We consider close-sourced LLMs, i.e., GPT-3.5-turbo and GPT-4-turbo, as well as open-sourced LLMs with different parameter scales, including CodeLlama-13B(or 7B)-Instruct-hf [41], Mistral-7B-Instruct-v0.2 [22], Vicuna-13B-v1.5 [71], and Baichuan2-13B-Chat [68].
-Choices of GNNs: To comprehensively investigate the effectiveness of different graph learning methods in enhancing task planning, we consider a wide range of graph neural networks, including SGC [47], GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32].
+Choices of GNNs: To comprehensively investigate the effectiveness of different graph learning methods for task planning, we consider a wide range of graph neural networks, including SGC [64], GCN [25], GAT [56], GraphSAGE [15], GIN [66], and Graph Transformers [46].
 5.2 Performance of the Training-free Approach
 We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 2.2.
-(2) GraphSearch [23, 36, 22] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score to the selection.
+(2) GraphSearch [32, 51, 31] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score for node selection.
 Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores.
-(3) SGC [47]: Employ a training-free SGC for task retrieval based on decomposed task steps.
-Illustrations of these methods are shown in Figures 5 and 6 in Appendix.
-Table 1 shows both the overall performance and token consumption costs.
+(3) SGC [64] employs a training-free SGC for task retrieval based on decomposed task steps.
+The details of baselines are given in Section E.1 and illustrated in Figure 6.
+Table 1 shows both the overall performance and token consumption costs, with results of Accuracy (Acc) moved to Table 9 in Appendix.
 […]
@@
 LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ #Tok $\downarrow$
-Baichuan2 13B Direct 45.85 19.00 2.43 47.57 4.08 2.59 33.45 9.52 3.72 30.87 9.92 1.96
-GreedySearch 30.58 4.89 6.42 18.74 4.45 5.69 15.60 1.61 5.91 22.52 2.98 3.62
-AdaptiveSearch 39.30 10.41 10.81 33.24 9.22 11.17 34.39 12.73 16.71 30.33 10.00 8.71
-BeamSearch 41.06 9.59 24.69 32.24 9.09 21.60 36.18 13.18 23.83 30.97 7.61 9.08
-SGC 56.53 29.94 2.28 56.75 31.62 2.43 62.31 36.69 3.53 32.97 9.11 1.84
 Vicuna 13B Direct 50.46 21.27 2.50 53.57 23.19 2.64 73.70 45.80 3.82 44.66 14.01 2.02
@@
 SGC 59.62 31.98 2.31 61.78 37.60 2.43 83.33 63.77 3.82 48.79 15.99 1.89
-CodeLlama 7B Direct 58.06 29.39 2.44 59.44 30.83 2.57 84.12 62.89 3.82 65.67 41.99 1.94
-GreedySearch 58.71 31.56 5.84 62.83 38.12 5.35 82.51 63.83 7.08 65.51 42.60 3.12
-AdaptiveSearch 60.42 33.18 6.84 62.32 36.81 5.50 83.42 64.15 7.83 65.37 40.64 5.00
-BeamSearch 60.34 31.36 17.95 64.12 38.99 21.48 83.25 63.48 24.48 64.60 40.50 5.78
-SGC 63.98 39.27 2.30 67.04 45.04 2.43 87.73 70.49 3.59 66.15 42.62 1.88
 Mistral 7B Direct 60.60 30.23 2.49 69.83 39.85 2.64 84.26 53.63 3.77 62.23 22.02 1.96
@@
 SGC 77.79 52.20 2.03 88.54 69.83 2.10 97.35 85.76 3.16 82.27 56.37 1.62
-Compared with direct inference, integrating SGC consistently improves performance, underscoring the effectiveness of the proposed method.
+Compared with direct inference, integrating an SGC consistently improves performance, underscoring the effectiveness of the proposed method.
 GraphSearch-type methods rely on beam search to identify paths and employ LLMs for evaluation, where longer processing times generally lead to better outcomes.
-Notably, our proposed method achieves comparable or superior performance to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 6).
-The case studies are provided in Appendix F.4.
+Notably, our proposed method achieves comparable or superior performance (Table 1 and Table 9) to BeamSearch while requiring $5$-$10$ times fewer tokens (Table 1) and inference time (Table 10).
+The case studies are provided in Appendix H.
 However, we observed only marginal improvements with GPT-4-turbo.
 A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets.
-In addition, the language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output.
-A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11.
+In addition, the pre-trained language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output.
+A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11 in Appendix.
 5.3 Performance of the Training-based Approaches
@@
 The TMDB dataset is excluded due to its limited sample size.
-Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM, whose role is to generate node embeddings derived from task names and descriptions.
-Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 8 and Table 9.
+Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-355M), whose role is to generate node embeddings derived from task names and descriptions.
+Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 11 and Table 12 in Appendix.
+[…]
 Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method.
@@
 This pattern indicates that the task’s challenge may not lie in the expressiveness of the models but rather in their ability to generalize.
-Efficiency: The training cost is remarkably low because we use e5-355M [44] as the text embedding model for GNNs.
+Regarding baselines, TAPE is unsuitable for task planning as its classification approach simplifies task planning, overlooking task dependencies.
+While GraphToken demonstrates superior performance over LLMs’ direct inference, we have noted instances of minor hallucination.
+This observation suggests that GraphToken’s understanding of the task graph is not yet perfect.
+In addition, GraphToken is limited to open-sourced LLMs.
+The proposed approaches also greatly improve the parameter prediction performance, as given in Appendix G (e.g.
+$9\%$ improvement to to GPT-3.5 and $3\%$ improvement to GPT-4).
+Efficiency: The details of the training time are given in Table 13.
+The training cost is remarkably low because we use e5-355M [61] as the text embedding model for GNNs.
 If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes.
@@
 This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs.
-The details of the training time are given in Table 7.
-Table 2: Performance of training-based GNN. The performance of all other GNNs are given in Table 8 and Table 9 in the Appendix.
+[…]
 HuggingFace Multimedia Daily Life
-LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
-Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
-GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
-GraphSAGE (co-train) 59.76 35.59 57.97 33.29 63.21 38.10
-Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
-GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
-GraphSAGE (co-train) 62.82 37.04 65.89 42.18 84.23 65.44
-CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
-GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
-GraphSAGE (co-train) 67.19 42.94 70.00 48.28 87.81 70.20
-Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
-GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
-GraphSAGE (co-train) 67.61 43.14 76.96 55.46 87.61 66.75
-CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
-GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
-GraphSAGE (co-train) 68.92 44.85 76.28 55.41 93.30 79.51
-GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
-GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
-GraphSAGE (co-train) 77.87 53.04 85.51 66.56 96.34 86.09
-GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
-GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
-GraphSAGE (co-train) 78.49 52.62 88.86 70.25 97.42 85.80
+LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$
+Vicuna-13B Direct 50.46 21.27 8.72 53.57 23.19 11.20 73.70 45.80 24.43
+TAPE 59.47 NA 5.07 54.97 NA 2.07 73.26 NA 12.50
+GraphToken 63.37 31.54 15.61 65.40 36.38 19.87 81.65 48.29 43.37
+GraphSAGE 61.86 35.68 20.08 63.71 39.88 21.37 86.07 67.63 48.64
+[…]
+Mistral-7B Direct 60.60 30.23 16.36 69.83 39.85 25.05 84.26 53.63 44.52
+TAPE 61.82 NA 6.13 58.92 NA 3.29 76.40 NA 16.44
+GraphToken 64.42 32.04 18.60 72.31 42.60 30.31 86.82 57.06 53.99
+GraphSAGE 68.12 43.09 25.77 75.51 52.94 34.29 87.51 66.57 52.74
+[…]
+CodeLlama-13B Direct 57.55 28.88 14.29 68.57 41.79 24.10 91.20 76.07 66.40
+TAPE 64.03 NA 8.05 58.27 NA 2.01 77.74 NA 17.37
+GraphToken 62.15 32.55 20.08 74.57 47.60 35.06 92.50 73.57 69.42
+GraphSAGE 67.30 42.41 26.56 74.93 54.52 38.55 93.84 80.38 73.60
+[…]
+Direct 73.85 45.73 28.95 82.85 62.07 47.96 96.09 83.65 81.30
+TAPE 68.00 NA 8.83 62.43 NA 3.87 70.67 NA 8.92
+GPT-3.5-turbo GraphSAGE 77.90 52.68 35.11 85.29 65.80 53.55 96.43 86.26 83.13
+[…]
+Direct 77.60 52.18 33.68 88.29 69.38 60.56 97.36 84.58 86.77
+TAPE 68.82 NA 9.50 63.94 NA 4.02 71.51 NA 9.40
+GPT-4-turbo GraphSAGE 78.76 52.53 34.09 88.63 69.65 60.36 97.34 85.67 86.97
+[…]
+Table 3: Performance Comparison of SGC and GraphSAGE on the UltraTool Benchmark [19]. Integrating GNNs can lead to more significant improvements in LLMs on larger task graphs.
+0-shot 1-shot
+LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ Acc $\uparrow$ # Tok $\downarrow$
+Direct 38.88 16.42 13.58 10,535 57.64 30.44 26.25 10,737
+BeamSearch 49.71 22.51 17.08 26,008 64.93 36.23 33.47 23,023
+SGC 61.07 37.61 25.31 10,456 71.64 44.00 39.68 10,658
+CodeLlama-13B GraphSAGE 63.78 39.91 27.98 10,456 72.81 45.26 43.49 10,658
+Direct 54.35 21.35 18.33 8,462 63.58 30.85 25.00 8,614
+BeamSearch 55.40 28.02 19.76 21,979 63.41 34.05 26.28 20,813
+SGC 59.80 37.82 25.87 8,352 64.96 37.96 29.70 8,504
+GPT-3.5-turbo GraphSAGE 63.97 42.26 30.35 8,352 70.49 47.79 39.74 8,504
+Direct 68.63 40.01 27.20 8,513 69.54 41.79 28.17 8,693
+BeamSearch 71.29 43.99 30.40 18,793 71.99 44.54 31.62 20,515
+SGC 70.87 44.01 31.60 8,346 70.46 44.82 33.00 8,504
+GPT-4-turbo GraphSAGE 70.67 43.83 34.40 8,346 70.75 47.68 37.22 8,504
+5.4 Scaling to Large Task Graphs
+To demonstrate the scalability of our method to larger task graphs, we conducted a supplementary experiment on the newly released planning benchmark UltraTool [19], which features a relatively large task graph with $260$ nodes.
+Details on processing this dataset are provided in Appendix C.3.
+We present a performance comparison of GNN models (training-free SGC and training-required GraphSAGE) against strong baselines like BeamSearch in Table 3.
+Among the metrics, accuracy (Acc) is calculated based on whether the predicted tasks match the ground-truth tasks, measuring the success rate at each case level.
+In such conditions, integrating a GNN significantly enhances performance and mitigates planning failures, e.g., GPT-4-turbo undergoes a 9.05% accuracy improvement with the introduction of GraphSAGE.
+The results indicate that (1) LLMs’ performance is vulnerable to the scale of task graphs; (2) The performance gain of the proposed method increases with a larger task graph.
 (a) Different Prompts
 (b) Fine-tuned LLMs
-Figure 3: Orthogonal Effectiveness of GNN on either different prompt templates (on HuggingFace dataset) or fine-tuned LLMs.
+Figure 3: Orthogonal Effectiveness to both Improved Prompts and Fine-tuned LLMs
 5.5 Improved Prompts and Fine-tuned LLMs
-In this subsection, we show that the proposed method is orthogonal to two dominant methods, i.e., the prompt engineering and finetuning.
+In this subsection, we show that the proposed method is orthogonal to two dominant methods, i.e., prompt engineering and fine-tuning.
 Orthogonal to Improved Prompts: We investigate GNN’s effectiveness when applied to improved prompt templates, i.e., strategically designed prompts that enhance the task planning abilities of LLMs.
 Specifically, we consider two types of prompts: (1) In-context Learning with Increased Examples [44] During main experiments, we maintain the consistent 1-shot in-context learning example for LLM’s direct inference.
-To realize further improvements, we increase the number of examples to $2$, and results under this setting are denoted as “2-shot Prompt”; (2) Plan like a Graph (PlaG) [20] We adopt the prompt in [20] to encourage LLM to think and plan in a graph-like manner, we convert the entire task graph into plain text and then integrate PlaG instructions to enhance LLM’s task planning.
+To realize further improvements, we increase the number of examples to $2$, and results under this setting are denoted as “2-shot Prompt”; (2) Plan like a Graph (PlaG) [29] We adopt the prompt in [29] to encourage LLM to think and plan in a graph-like manner.
+Specifically, we convert the entire task graph into plain text and then integrate PlaG instructions to enhance LLM’s planning.
 Results under this prompt template are denoted as “PlaG Prompt”.
-From the results shown in Figure 3(a), where we apply three different prompts to CodeLlama-13B and Mistral-7B on HuggingFace, it is clear that applying a GNN to improved prompts, where task steps are more concisely decomposed and predictions are more accurate, can also boost performance.
+From the results shown in Figure 3(a), where we apply three different prompts to CodeLlama-13B and Mistral-7B on HuggingFace, it is clear that applying GraphSAGE to improved prompts, where task steps are more concisely decomposed and predictions are more accurate, can also boost performance.
 Orthogonal to LLMs’ Fine-tuning: To explore whether our framework maintains effectiveness on fine-tuned LLMs, which have acquired dataset-specific task planning capabilities, we conduct further experiments.
 For each dataset, we use LoRA [18] to fine-tune two LLMs of different parameter scales, including CodeLlama-7B and Vicuna-13B, based on the same training data as GNNs.
-Details of fine-tuning process are provided in Appendix D.4.
+Details of fine-tuning process are provided in Appendix F.3.
 The finetuned model is named as “FT-CodeLlama” and “FT-Vicuna” in Figure 3(b).
 The results depicted in Figure 3(b) demonstrate that fine-tuning markedly enhances the task-planning capabilities of LLMs.
-Furthermore, applying GraphSAGE to the decomposed tasks of LLMs further improves the accuracy of the task planning.
-6 Related Works
-Due to space limitations, this section focuses on related works of task planning, while discussions of other relevant areas are deferred to Appendix A.
+Furthermore, applying GraphSAGE to the decomposed tasks of LLMs further improves the accuracy of task planning.
+6 Conclusions
+This paper presents an initial exploration into graph-learning-based approaches for task planning in language agents.
+Through theoretical analysis, we demonstrate the inductive bias of the attention mechanism and the utility of auto-regressive loss impedes their effectiveness in task planning.
+We propose to integrate GNNs for task graph analysis, which yields performance improvements across a range of LLMs and planning benchmarks.
+Limitations: Despite the encouraging performance, there are limitations that highlight significant opportunities for enhancement.
+Firstly, our proposed method, while effective, is straightforward; more sophisticated graph-learning-based decision-making algorithms could potentially offer further improvements.
+Secondly, the construction of the task graph currently requires manual effort.
+Investigating automated graph generation techniques for diverse applications is another promising direction for future work.
+Acknowledgements
+This work is partly supported by grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14217622).
+Appendix A Related Works and Discussions
+A.1 Planning Algorithms in LLMs
 The existing studies of task planning approaches can be categorized into several directions, including task decomposition, multi-plan selection, the use of external planners, reflection, and memory-aided planning [21].
@@
 Subsequently, LLMs evaluate these plans to select the most effective one.
-The GraphSearch methods used in our baselines fall into this category.
-External planner approaches [21] formalize tasks using the Planning Domain Definition Language (PDDL) and employ classic solvers to address the planning problem, which is closely related to the proposed approaches in this paper.
-The task planning considered here is more flexible and cannot be directly translated into formal languages.
-However, we demonstrate that GNNs can serve as an effective external planner.
+The “GraphSearch” methods used in our baselines fall into this category.
+External planner approaches [30] use LLMs to convert the problem into Planning Domain Definition Language (PDDL) and then employ classic solvers to address the planning problem.
+PDDL requires a pre-defined goal, for example, moving the blocks from one state to another state.
+However, the goal of task planning investigated in language agents deals with more flexible and personal goals, spanning personal needs in video, text, and image processing.
+Translating these goals into formal PDDL is very difficult.
+We instead demonstrate that GNNs can serve as an effective external planner in this application.
 Reflection-based methods [47] focus on reflecting upon experiences to refine the plan, while memory-aided planning approaches [72] utilize external experiences, such as those from search engines.
 These approaches are deployed in interactive environments and orthogonal to this paper.
-7 Conclusions
-This paper presents an initial exploration into graph-learning-based approaches for task planning.
-Through theoretical analysis, we demonstrate the inductive bias of the attention mechanism and the utility of auto-regressive loss impedes their effectiveness in task planning.
-We propose to integrate GNNs for task graph analysis, which yields performance improvements across a range of LLMs and datasets.
-Limitations: Despite the encouraging performance, there are limitations that highlight significant opportunities for enhancement.
-Firstly, our proposed method, while effective, is straightforward; more sophisticated GNN-based decision-making algorithms could potentially offer further improvements.
-Secondly, in our current framework, GNNs function as an external module with limited interaction with LLMs.
-Exploring the synergies between GNNs and LLMs, particularly by incorporating GNN outputs as tokens within LLMs, presents an interesting avenue for research.
-Thirdly, the construction of the task graph currently requires manual effort.
-Investigating automated graph generation techniques for this application is another promising direction for future work.
-Appendix A More Related Works
-A.1 LLMs for Graphs
-With the breakthroughs in LLMs, there has been a surge of interest in applying LLMs to graph-related problems.
+A.2 Task Planning in Traditional AI
+Apart from task planning in language agents, there is also a domain in traditional AI called task planning [1, 6, 17, 54, 43, 50, 34, 10].
+Task planning in traditional AI is defined as $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C},\mathcal{G},s_{0})$, where $\mathcal{S}$ is the states, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, a cost function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$, a set of goal states $\mathcal{G}\subseteq\mathcal{S}$, and an initial state $s_{0}$.
+An agent following a policy $\pi:\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ will start in state $s_{0}$, then repeatedly choose an action $a\sim\pi(a|s)$ and execute it to reach a new state $s^{\prime}\sim\mathcal{T}(s^{\prime}|s,a)$, incurring a cost $\mathcal{C}(s,a)$ along its way.
+An optimal policy $\pi^{*}(a|s)$ is one that reaches the goal state with probability $1$ while minimizing the total expected cost.
+Traditionally, task planning is solved by reinforcement learning approaches [6] and heuristic A approaches [17].
+The neural network-based approaches are employed to accelerate the computation and improve the performance [54, 43, 50, 34, 10].
+The task planning in language agents is a different application.
+It cannot be solved as a constraint satisfaction problem, since both the features of task graph and user request are expressed in natural language.
+Compared with traditional planning, task planning for language agents involves diverse and open-ended goals due to the varied personal requirements users have.
+For example, on platforms like Hugging Face, users’ intentions span across video, text, and image domains.
+On the contrary, classic planning has a fixed goal for a given domain, e.g., in the n-puzzle [54], the goal is formally as placing the tiles in numerical order.
+Within this new application domain, while existing research primarily focuses on prompt design for pre-trained LLMs, our work underscores the importance of traditional planning methods, such as GNNs, in complementing LLMs.
+A.3 Planning in Agents and Neuroscience
+Planning is a pivotal topic in both agents and neuroscience, where graphs play an indispensable role. We believe the concepts and insights presented in this paper are useful to these fields.
+The TaskBench, RestBench, and UltraTool dataset used in this paper belongs to the tool agents.
+The Math agent, AlphaGeometry, employs LLMs to generate auxiliary constructions in geometry [55].
+Considering lemmas as nodes and their interdependencies as edges, the endeavor to prove a theorem resembles the task of identifying a route to the theorem node within the graph constituted by potential lemma nodes and the edges that represent their interdependencies.
+There are no explicit task graphs in game agents [57], embodied agents [20], and code agents [47].
+The core strategy in these domains is to employ verbal reinforcement learning within LLMs.
+The state and transitions in reinforcement learning can be modeled as nodes and edges in the graph.
+In addition, there are case-by-case graph models in these agent applications.
+For example, in code agents, one can view the class as the nodes and dependencies as the edges.
+In the embodied agents, the objects in the environment can be viewed as the nodes.
+In the neuroscience, animal planning is often assessed through path planning in mazes [2, 63].
+Inspired by these animal experiments, planning testbenches have been developed for LLMs [36].
+A computational model known as the Tolman-Eichenbaum Machine (TEM) has been proposed to decipher the mechanisms of general planning in animals across various environments, such as mazes [63].
+The TEM model posits that hippocampal cells, including place and landmark cells, remap between environments, while entorhinal cells exhibit a range of properties that mirror spatial responses, including grid, band, border, and object-vector cells.
+In essence, hippocampal cells map sensory inputs onto locations in abstract graphs and remap, and entorhinal cells execute graph operations.
+A.4 LLMs for Graphs
+With the breakthroughs in LLMs, there has been a surge of interest in applying LLMs to graph-related problems [26, 27].
 GPT4Graph [14] and NLGraph [58] are two prominent benchmarks designed to evaluate the performance of LLMs in the context of graph tasks.
@@
 The analysis in Section 3.3 aims to shed light on the empirical observations reported in these studies.
-Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [37] and GraphLLM [6].
+Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [52], GraphLLM [9], and GraphToken [37].
 These approaches have demonstrated significant improvements in performance on GNN-related tasks.
@@
 A promising future direction involves using task planning data generated by GPT to fine-tune graph foundation models, such as GraphGPT [52], and applying them to task planning.
-A.2 Theoretical Analysis of Reasoning
+This paper proposes to use task planning as a new benchmark for this line of research.
+A.5 Theoretical Analysis of Reasoning
 Reasoning is closely related to task planning and decision-making.
@@
 These theoretical contributions are novel and promise to be valuable for general reasoning and planning tasks.
-A.3 GNNs for Decision-making
+A.6 GNNs and GraphSearch for Combinatorial Optimization
 GNNs are popular approaches for solving decision-making problems on graphs.
@@
 Given this connection, a promising future direction involves repurposing GNNs for decision-making approaches in task-planning applications.
-Appendix B Proofs
-B.1 Proof of Theorem 1
-Assumption 1.
-Each function $f,\oplus$ in (2) can be approximated by constant size MLP.
-Assumption 2.
-The aggregation function $\square$ in (2) is one of min, max, sum, mean.
-The first assumption is mild as MLPs are universal approximators. The second assumption is mild because these are the most commonly used aggregation functions.
-Theorem 3.
-(Expressiveness) Assume the input format is given in (3) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2.
-There exists a log-precision constant-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
-As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
-Proof.
-Token Embedding and Positional Embedding: The three-dimensional token embedding includes the token type $e^{\text{type1}}$ (00 for answer; $1$ for node; $2$ for edge cost), refined token type $e^{\text{type2}}$ (00 for answer, $1$ for the node tokens in initial states, $2$ for the target node tokens in the edge list, $3$ for the source node tokens in the edge list, $4$ for edge cost), and the token id $e^{\text{token}}$ (from 00 to $|V|-1$).
-The two-dimensional positional embedding includes the embedding for initial state tokens $e^{\text{pos1}}$ (00 for edge list tokens, $1$ for the first two elements of initial states, $2$ for the second two elements of initial states, etc.), embedding for edge list tokens $e^{\text{pos2}}$ (00 for initial state tokens, $1$ for the first three elements of the edge list, $2$ for the second three elements of the edge list, etc.).
-There are also constant-dimensional placeholders to put the states of DP.
-Block 1 - Initial State Broadcast: The goal of the first block is to broadcast the initial states from the initial state token to node tokens.
-(1) Use MLPs to recover the digits of the answer tokens and put them in the first placeholder if $e^{type}_{k}==0$; (2) Copy the first placeholder from answer token to its previous node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos1}}_{k}-e^{\text{pos1}}_{j})^{2}<\delta\}$; (3) Broadcast the first placeholder with MEAN in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{type1}}_{k}-e^{\text{type1}}_{j})^{2}+(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}<\delta\}$.
-Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}]]$.
-Block 2 - Edge Feature Operations: The goal of the second block is to copy the edge features from the edge feature token to the corresponding node tokens.
-(1) Use MLPs to recover the digits of the edge feature tokens and put them in the second placeholder if $e^{\text{type1}}==2$; (2) Copy the second placeholder from the edge feature token to the node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos2}}_{k}-e^{\text{pos2}}_{j})^{2}<\delta\}$.
-Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}]]$.
-Block 3 - Message Preparation: (1) Use MLPs to compute $\oplus$ and place the results in the third placeholder.
-Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}],\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$; (2) Use MLPs to clean up the first and second placeholder.
-Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos1}},\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{i}]]$
-Block 4 - Message Passing: The goal of the fourth block is to compute $\square$ and $f$.
-(1) Use one or two attention heads (one for max, min, mean aggregations, and two for sum aggregations) to perform the aggregation operation.
-This is achieved by using MEAN or MAX or SUM for the first placeholder and setting $\mathcal{S}_{k}=\{j|(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}+(e^{\text{type2}}_{k}-e^{\text{type2}}_{j})^{2}<\delta\}$.
-Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\square_{v_{j}\in\mathcal{T}(u_{i})}\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]]$; (2) Use MLPs to compute $f$.
-Now the state for every node token is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},f(\square_{v_{j}\in\mathcal{T}(i)}(\text{Answer}[0][u_{i}]\oplus c[u_{i}][v_{j}]))]$.
-After four blocks, the final state for every node token $u_{i}$ is given by $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[1][u_{i}]]$.
-$\text{Answer}[k][u_{i}]$ can be obtained by repeating the above four blocks $k$ times.
-∎
-Lemma 1.
-[8] Let $n\in\mathbb{N}$ be an integer and $\bm{x}_{1},\cdots,\bm{x}_{n}$ be a sequence of vectors where $\bm{x}_{i}=(\tilde{\bm{x}}_{i},r_{i},1)\in[-M,M]^{d+2}$ where $M$ is a large constant.
-Let $\bm{K},\bm{Q},\bm{V}\in\mathbb{R}^{d^{\prime}\times(d+2)}$ be any matrices with $\|\bm{V}\|_{\infty}\leq 1$ and let $0<\rho,\delta<M$ be any real numbers.
-Denote $\bm{q}_{i}=\bm{Q}\bm{x}_{i},\bm{k}_{j}=\bm{K}\bm{x}_{i},\bm{v}_{j}=\bm{V}\bm{x}_{j}$.
-Define a matching set $\mathcal{S}=\{j||\bm{q}_{i}^{T}\bm{k}_{j}|\leq\rho\}$.
-Define two following operations
-  • COPY: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$ with $\bm{u}_{i}=\bm{v}_{\text{pos}(i)}$, where $\text{pos}(i)=\arg\max_{j\in\mathcal{S}_{i}}r_{j}$.
-  • MEAN, MAX, SUM: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$, where $\bm{u}_{i}=\square_{j\in\mathcal{S}_{i}}\bm{v}_{j}$ and $\square$ is min or max or sum or mean.
-Specifically, for any sequence of vectors $\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}$, denote the corresponding output of the attention layer as $\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}$.
-Then, we have $\|\bm{u}_{i}-\bm{o}_{i}\|_{\infty}\leq\epsilon$ for all $i\in[n]$ and $\mathcal{S}\neq\emptyset$.
-B.2 Proof of Proposition 2
-Proposition 2.
-Assume the input format is as described in Equation (3) and that the attention mechanism is limited to attending to a constant number of tokens.
-There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
-Proof.
-We present a proof by contradiction.
-Assume that a token in a Transformer with a constant depth, constant width, and log-precision can attend to only a constant number of nodes.
-Under this assumption, the total information accessible to the token in such a Transformer architecture amounts to $O(\log n)$ bits.
-However, for a graph with $|V|$ nodes, the number of possible outcomes from executing one-step DP is $O(e^{|V|})$, necessitating $\Theta(|V|)$ bits for representation.
-By the pigeonhole principle, this scenario inevitably leads to at least two distinct DP outcomes being represented by the same output sequence generated by the model, thereby constituting a contradiction.
-∎
-B.3 Proof of Theorem 2
-Theorem 4.
-(Spurious correlations of auto-regressive loss) Assume (1) the loss employed is a next-token-prediction loss utilizing cross-entropy, applied to the sub-sequence $v_{1}\ v_{2}\ \cdots\ t$ during training; (2) the output logits are determined by target node $t$ and the current node $v_{i-1}$.
-Let $N_{t,v_{i-1},u}$ be the number of times in the training dataset such that $t$ is the target node, $v_{i-1}$ is the current node and $v_{i}=u$ is the next node.
-The optimal logits for predicting the next node $u$ from current node $v_{i-1}$ towards target node $t$ is given by $\hat{\bm{v}}_{i}[u]=\frac{N_{t,v_{i-1},u}}{\sum_{u}N_{t,v_{i-1},u}}$ if $\sum_{u}N_{t,v_{i-1},u}>0$.
-If $\sum_{u}N_{t,v_{i-1},u}=0$, $\hat{\bm{v}}_{i}[u]$ can be any non-negative number subject to $\sum_{u}\hat{\bm{v}}_{i}[u]=1$.
-Proof.
-We denote $\mathcal{D}$ as the training dataset, $L_{i}$ as the sequence length of the $i$-th sequence in the dataset, $\bm{v}_{i,j}$ as the one hot embedding of the $j$-th token in the $i$-th training sequence, and $\hat{\bm{v}}_{i,j,u}$ as the $u$-th logit at the $j$-th token in the $i$-th sequence.
-The cross-entropy loss is given by
-$-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\bm{v}_{i,j,u}\log\hat{\bm{v}}_{i,j,u}=-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\mathbb{I}_{u=v_{i,j}}\log\hat{\bm{v}}_{i,j,u}\overset{(a)}{=}-\sum_{t,v_{j-1},u}N_{t,v_{j-1},u}\log\hat{\bm{v}}_{i,j,u}$
-=$\overset{(b)}{=}$ $-\sum_{t,v_{j-1},u}\left(\sum_{u}N_{t,v_{j-1},u}\right)\left[\left(\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}\right)\log\hat{\bm{v}}_{i,j,u}\right],$
-where (a) uses the assumption that the output logits are determined by target node $t$ and the current node $v_{i-1}$.
-In (b), we assume that $\sum_{u}N_{t,v_{i-1},u}\neq 0$.
-If $\sum_{u}N_{t,v_{i-1},u}=0$, then the corresponding logits will not affect the loss function and $\hat{\bm{v}}_{i,j,u}$ can take any number.
-The cross-entropy is minimized when $\hat{\bm{v}}_{i,j,u}=\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}$.
-∎
-Appendix C More Discussions
-C.1 Examples of Dynamic Programming
-Longest Increasing Subsequence: The Longest Increasing Subsequence (LIS) problem is a classic dynamic programming problem that involves finding the length of the longest subsequence within a given array arr where the elements are in strictly increasing order.
-The state transition function for the LIS problem can be expressed as:
-$\text{Answer}[k][i]=\max_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+(\mathbb{I}(j\neq i)\times 1)),$
-where $\mathcal{T}(i)=\{i\}\cup\{j\;|j<i\;\text{and}\;\text{arr}[j]<\text{arr}[i]\}$ denotes the set of states that can transfer to state $i$, the aggregation function $\square$ is implemented as $\max(\cdot)$, and the cost $c[i][j]$ is $1$ for those candidate states that are not equal to state $i$ as adding the element leads to a longer subsequence while 00 for the state itself.
-Bellman-Ford Algorithm: The Bellman-Ford algorithm is also a classic dynamic programming algorithm used to find the shortest path from a single source vertex to all other vertices in a weighted graph.
-The core idea behind the Bellman-Ford algorithm is that the distance from the source vertex to a target vertex can be computed as the minimum distance from the source to any of the target’s neighboring vertices, plus the weight of the edge connecting the neighbor to the target.
-Therefore, the state transition function for the Bellman-Ford algorithm can be expressed as:
-$\text{Answer}[k][i]=\min_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+w[j][i]),$
-where $\text{Answer}[k][i]$ represents the length of the shortest path from the source vertex to node $i$ at the $k$-th iteration, $\mathcal{T}(i)=\mathcal{N}^{-}(i)$ denotes the set of in-neighbors of node $i$, and $w[j][i]$ denotes the weight from node $j$ to $i$.
-The aggregation function $\square$ is implemented as $\min(\cdot)$ as we try to find the shortest path.
-Travelling Salesman Problem: This problem is defined as, given a set of cities and the distances between every pair of cities, finding the shortest possible route that visits every city exactly once and returns to the starting point.
-If we regard the set of already visited city $\mathcal{S}$ ending at $i$-th city as the current state, then states that can transfer to current state are those that ending city can reach $k$.
-Therefore, the state transition function for the TSP problem can be expressed as:
-$\text{Answer}[k][i][\mathcal{S}]=\min_{j\in\mathcal{S},j\neq i}(\text{Answer}[k-1][j][\mathcal{S}\setminus\{i\}]+w[j][i]),$
-where $\text{Answer}[k][i][\mathcal{S}]$ represents the cost of the shortest tour that visits all the cities in the set $\mathcal{S}$ and ends at the $i$-th city, the aggregation function is still implemented as $\min(\cdot)$ since we aim to find the shortest path, and $w[j][i]$ denotes the distance from city $j$ to city $i$.
-C.2 Permutation Invariance Test of LLMs
-Figure 4: Illustrative Examples of LLMs Failure to Solve Graph Problems under Node-ID Permutation. Experiments were repeated for 30 times.
-Appendix D Implementation Details
-D.1 Details of Prompt Templates
-To facilitate reproducibility, we provide prompt templates for all LLM-related queries, including direct inference and GraphSearch in Tables 3 and 4, respectively.
-GraphSearch consists of two types of LLM queries: Task Assessment, where LLM evaluates the suitability of each candidate task to address current step’s requirement during the iterative search, and Path Selection, where LLM selects the optimal path to fulfill user’s request after the search.
+Appendix B Prompts
 Table 4: Prompt template for LLM’s direct inference [44]
@@
 Now please generate your result in a strict JSON format: # RESULT #
-D.2 Implementation of Training-free Methods
-In this subsection, we present the implementation details of two training-free baselines, including LLM’s direct inference and GraphSearch. Method illustrations are shown in Figure 5.
-Figure 5: Illustration of LLM’s Direct Inference and GraphSearch Method.
-Figure 6: Illustration of our GNN-enhanced Task Planning.
-We first prompt LLM to decompose complex and implicit user request into several manageable sub-tasks.
-We then use either a training-free or training-required GNN to sequentially align each step description with a suitable task, ultimately generating the invocation path.
-LLM’s Direct Inference: During experiments, we uniformly apply 1-shot in context learning for LLM’s direct inference of task invocation path.
-For open-sourced LLMs, the temperature parameter is set to 0.2.
-GraphSearch: This algorithm conducts an iterative search on the task graph to identify an optimal task invocation path that can best satisfy a given request.
-In each iteration, the neighbors of the last selected task are considered as candidates.
-These candidates are evaluated by LLM for their suitability for the current step.
-The search process follows a depth-first approach.
-After the task assessment in the final step, a set of potential invocation paths is generated.
-Subsequently, LLM is prompted to select the most appropriate path from these options.
-The GraphSearch algorithm is implemented in three distinct variants, each employing a unique task selection strategy:
-  • GreedySearch consistently selects the task node with the highest score at each step. Although fast and simple, this approach can lead to cascading errors, resulting in degraded performance.
-  • AdaptiveSearch selects tasks with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode. During experiments, we empirically set the score threshold to 3.
-  • BeamSearch retains the top-$k$ tasks based on the LLM’s assessment scores within candidates.
-Beam search can expand the search space but slightly reduces the efficiency.
-We uniformly set the beam width to 2.
-SGC: For the implementation of SGC, the parameter $\alpha$ is uniformly set to 0.7.
-Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC on HuggingFace dataset, Roberta-large-355M [26] serves as the text encoder.
-For all other datasets and LLMs, the e5-355M [44] configuration is employed.
-D.3 Implementation of Training-based GNNs
-LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M [44] as the LM backbone.
-For the graph encoder, our setup includes a single layer with a hidden dimension of 1024.
-During the model training, we set the batch size to 512 and run for 20 epochs with a learning rate of 1e-3.
-We use the Adam optimizer [17] and implement an early stopping mechanism with a patience of 5 epochs to prevent over-fitting.
-All experiments are conducted on a single NVIDIA A100-80G GPU.
-Training Data Preparation: From each dataset in TaskBench, we randomly select $3,000$ samples to create the trainset.
-The original data includes specific task steps and corresponding task invocation paths.
-Therefore, we first employ a topological sort to align each task step accurately with corresponding task, forming “<step, ground-truth task>” pairs.
-Then, for each pair, we randomly sample 2 negative tasks to constitute the “<step, positive task, negative task>” triplets for model training.
-These negative samples are selected based on how textually similar they are to the positive one, creating a robust differentiation challenge for the model.
-Choices of Different Configurations: Our training-based method offers two configuration options: training only the GNN while keeping the LM frozen, or co-training both the LM and GNN.
-The first configuration is designed to explore GNN’s capability in task retrieval.
-The latter leverages LM’s dataset-specific semantic embeddings to enhance performance.
-Illustrations of these configurations are provided in Figure 7.
-For the co-training setup, we use a learning rate of 2e-5 and a training duration of 10 epochs.
-Figure 7: LM+GNN Configuration. We offer two configurations: only training the GNN while keeping the LM frozen (Table 8), or co-training both the LM and GNN (Table 9).
-D.4 Fine-tuning LLMs
-To explore the effectiveness of our proposed framework on fine-tuned LLMs, we employ Supervised Fine-Tuning (SFT) on the CodeLlama-7B and Vicuna-13B models.
-We use the same set of $3,000$ samples from GNN training as the LLMs’ fine-tuning data.
-In fine-tuning the LLM with LoRA [13], we set the lora_r parameter (dimension for LoRA update matrices) to 8 and the lora_alpha (scaling factor) to 16.
-The dropout ratio is set to 0.1, the batch size to 2, and we conduct training over 2 epochs with a learning rate of 1e-5.
-For HuggingFace dataset, the maximum input length is set to 800, while the maximum output length is 400.
-For Multimedia and Daily Life datasets, which contain a larger number of tasks and require longer textual inputs, we set the maximum input and output length to 1000 and 500, respectively.
-We utilize 2 NVIDIA A100-80G GPUs for fine-tuning the LLMs.
-Each dataset undergoes specific fine-tuning for both the CodeLlama-7B and Vicuna-13B models.
-Appendix E Datasets
-E.1 Overview
-We provide the statistics of experimental datasets from two task planning benchmarks in Table 5.
+Table 6: Prompt template for LLM’s filling in invocation parameters [44]
+# GOAL #
+Given a # USER REQUEST # and # PLANNED TASKS # to be invoked in sequence to solve this request, please fill up each invoked task’s invocation parameters.
+The format must be in strict JSON format, like:
+[…]
+[…]
+[…]
+[…]
+# REQUIREMENTS #
+1. Consider each task’s input and output requirements, and carefully fill in the arguments for each task.
+2. Analyze the resource dependencies, keeping in mind that these tasks are invoked sequentially to address the original request.
+3. The number of predicted task_nodes must strictly align with the provided tasks.
+# USER REQUEST #
+{{ user request }}
+# PLANNED TASKS #
+{{ a list of previously GNN retrieved tasks }}
+# DETAILS OF TASKS #
+{{ details, i.e., input and output requirements of each planned task }}
+# EXAMPLE #
+{{ in-context learning examples }}
+Now, please generate your response in a strict JSON format: # RESULT #
+Appendix C Datasets
+C.1 Overview
+We provide the statistics of experimental datasets from three task planning benchmarks in Table 7.
 The illustrative examples from each dataset are shown in Figure 4.
-For datasets from TaskBench, each sample consists of original user request, corresponding decomposed task steps, and ground-truth task invocation path.
-As RestBench includes only user requests and corresponding API invocation sequences, we prompt GPT-4 to infer decomposed task steps aligned with each invoked API, thereby finalizing the dataset.
-E.2 Reformatting Details of RestBench
-The TMDB dataset from RestBench, focuses on movie-related searching and recommending functions.
-To align RestBench with our experiments, we have implemented the following processing steps: Reformatting original APIs by assigning unique task names and descriptions: APIs in RestBench were represented by request paths, such as “GET /movie/top_rated”, referring to the API that retrieves top-rated movies on TMDB.
+For datasets from TaskBench [44], each sample consists of original user request, corresponding decomposed task steps, and ground-truth task invocation path.
+As RestBench [49] and UltraTool [19] include only user requests and corresponding API invocation sequences, we prompt GPT-4 to infer decomposed task steps aligned with each invoked API, thereby finalizing the dataset.
+C.2 Reformatting Details of RestBench
+The TMDB dataset from RestBench, focuses on movie-related searching and recommending functions. To align RestBench with our experiments, we have implemented the following processing steps:
+Reformatting original APIs by assigning unique task names and descriptions: APIs in RestBench were represented by request paths, such as “GET /movie/top_rated”, referring to the API that retrieves top-rated movies on TMDB.
 To enhance semantic differentiation among APIs, we first prompt GPT-4 to assign a unique name and a detailed functional description to each API.
 These names and descriptions were then manually verified and refined.
-For example, the API previously mentioned is renamed “Get Top-Rated Movies” with the description “This API retrieves a list of the highest-rated movies.” Note that though the original TMDB dataset contains 54 APIs, some were never invoked in any data examples.
+For example, the API previously mentioned is renamed “Get Top-Rated Movies” with the description: “This API retrieves a list of the highest-rated movies.” Note that though the original TMDB dataset contains $54$ APIs, some were never invoked in any data examples.
 Therefore, we focus only on those APIs that appeared in at least one user request, resulting in a refined set of $46$ APIs.
@@
 TaskBench RestBench
-Type Statistic HuggingFace Multimedia Daily Life TMDB
-Task Graph # Nodes 23 40 40 46
-# Links 225 449 1560 979
-Link Type Resource Resource Temporal Resource / Category
-All Data # Samples 7,546 5,584 4,320 100
-Test Set # Samples 500 500 500 94
-# Avg. Nodes 3.81 3.92 4.05 2.33
-# Avg. Links 2.81 2.92 3.05 1.33
-Figure 8: Illustrative examples of four experimental datasets.
-Appendix F Supplementary Experiments
-F.1 Full Results of Training-based GNNs
-In Table 8, we present comprehensive results for all training-based GNNs, including GCN [18], GAT [39], GraphSAGE [11], GIN [49], and TransformerConv [32].
-To highlight the improvements brought by GNNs, we also include results from the strongest baseline, BeamSearch.
-From the results, it is obvious that all GNN encoders significantly enhance the task planning abilities.
-For instance, when applied to Vicuna-13B on HuggingFace dataset, the introduction of GCN results in a performance improvement of $17.83\%$, GAT contributes to a $17.48\%$ increase, GraphSAGE leads to a $22.59\%$ boost, and TransformerConv improves predictions by $18.0\%$.
-This conclusion can be generalized across various LLMs and datasets, demonstrating the robust capabilities of GNNs.
-F.2 Performance of LM+GNN Co-trained Mode
-In this subsection, we conduct a supplementary study where the parameters of a pre-trained LM are also tuned along with GNN during model training.
-The model configuration is illustrated in Figure 7, and the results are presented in Table 9.
-The results demonstrate that, compared to the GNN-only tunable mode, co-training LM+GNN can lead to further performance improvements.
-This enhancement occurs because the language model acquires task-specific semantics, which makes the representations more discriminative and boosts the GNN’s effectiveness in task retrieval.
-Additionally, it is noted that under the co-training setup, the differences between various GNN encoders are relatively minor, with performance variations across GNNs for a specific LLM on any dataset remaining within $2\%$.
-For LM+GNN co-trained mode, as the parameters of the LM backbone, e5-355M, are also trained, more computational resources are required.
-We list the number of trainable parameters and the required training times for this architecture in Table 7.
-In application scenarios, the choice between GNN-only and LM+GNN co-trained modes can be decided based on the availability of resources and performance requirements.
-F.3 Full Efficiency Study
-In this subsection, we present a comprehensive efficiency study on various experimental setups: inference times across different training-free modes, time required for training GNN or LM+GNN, and the resources needed for fine-tuning LLMs.
-Results are shown in Tables 6 and 7.
-Table 6: Efficiency Study of Training-free Methods.
-Due to space constraints in the table, some LLMs are abbreviated such as “GPT-3.5” for “GPT-3.5-turbo”.
-For Vicuna-13B and Mistral-7B, some discrepancies in reported times may be attributed to variations in deployment across different A100 devices.
-SGC maintains efficiency comparable to LLM’s direct inference.
+Type Statistic HuggingFace Multimedia Daily Life TMDB UltraTool
+Task Graph # Node 23 40 40 46 260
+# Links 225 449 1560 979 611
+Link Type Resource Resource Temporal Resource / Category Resource
+All Data # Samples 7,546 5,584 4,320 100 3,527
+Test Set # Samples 500 500 500 94 500
+# Avg Nodes 3.81 3.92 4.05 2.33 2.38
+# Avg Links 2.81 2.92 3.05 1.33 1.38
+Figure 4: Illustrative details of experimental datasets.
+C.3 Reformatting Details of UltraTool
+Motivation: In our main experiments using TaskBench [44] and RestBench [49], we observed that strong LLMs like GPT-4-turbo already perform well.
+This may be attributed to two factors: (1) the training of GPT-4-turbo likely included knowledge relevent to these benchmarks, as both were released before GPT-4 and utilize popular platforms such as HuggingFace and TMDB; (2) the task graphs are relatively small, containing no more than $50$ tasks, which falls within the capabilities of LLMs.
+Therefore, we aim to evacuate planning performance as well as GNN’s effectiveness in a more challenging scenario, i.e., one that requires specific knowledge beyond the LLM’s training and involves larger task graphs that possibly exceed its memory and reasoning capabilities.
+Constructing a Task Graph: We utilized UltraTool [19], released in March 2024, which features complex planning scenarios across thousands of tasks, including travel, tourism, and other daily life domains.
+The original UltraTool contains $5,824$ samples with $2,032$ available tasks across $22$ distinct domains.
+However, we observed that some tasks appear in only one sample, making them quite rare.
+Therefore, we filtered this dataset as follows: first, we considered only tasks that appeared more than $5$ times across all samples, focusing on more common tasks that cater to daily life.
+Next, we retained samples that incorporated these filtered tasks, ensuring that the number of invoked tasks exceeded 2 to satisfy the multi-task planning scenario.
+After this filtering, we identified $260$ distinct tasks.
+We then constructed the task graph by treating each task as a node and adding links between tasks that were invoked sequentially in the dataset.
+For each task, we further verified and refined its functional descriptions to ensure semantic suitability.
+Reformatting Raw Data Examples: We allocated $500$ samples for testing and $3,000$ samples for GNN training.
+Although the original UltraTool provided decomposed steps, we found them too coarse-grained, making it difficult to align step descriptions with suitable tasks.
+Therefore, we employed a similar strategy by prompting GPT-4 to infer decomposed task steps that align with each invoked task.
+As a result, each data sample consists of the user request, corresponding steps, and ground-truth invoked tasks, ensuring high quality for GNN training.
+Appendix D Supplementary Materials for Theoretical Results
+D.1 Dynamic Programming
+Longest Increasing Subsequence: The Longest Increasing Subsequence (LIS) problem is a classic dynamic programming problem that involves finding the length of the longest subsequence within a given array arr where the elements are in strictly increasing order.
+The state transition function for the LIS problem can be expressed as:
+$\text{Answer}[k][i]=\max_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+(\mathbb{I}(j\neq i)\times 1)),$
+where $\mathcal{T}(i)=\{i\}\cup\{j\;|j<i\;\text{and}\;\texttt{arr}[j]<\texttt{arr}[i]\}$ denotes the set of states that can transfer to state $i$, the aggregation function $g(x,y)$ is implemented as $\max(x,y)$, and the cost $c[i][j]$ is $1$ for those candidate states that are not equal to state $i$ as adding the element leads to a longer subsequence while 00 for the state itself.
+Bellman-Ford Algorithm: The Bellman-Ford algorithm is also a classic dynamic programming algorithm used to find the shortest paths from a single source vertex to all other vertices in a weighted graph.
+The core idea behind the Bellman-Ford algorithm is that the distance from the source vertex to a target vertex can be computed as the minimum distance from the source to any of the target’s neighboring vertices, plus the weight of the edge connecting the neighbor to the target.
+Therefore, the state transition function for the Bellman-Ford algorithm can be expressed as:
+$\text{Answer}[k][i]=\min_{j\in\mathcal{T}(i)}(\text{Answer}[k-1][j]+w[j][i]),$
+where $\text{Answer}[k][i]$ represents the length of the shortest path from the source vertex to node $i$ at the $k$-th iteration, $\mathcal{T}(i)=\mathcal{N}^{-}(i)$ denotes the set of in-neighbors of node $i$, and $w[j][i]$ denotes the weight from node $j$ to $i$.
+The aggregation function $g(x,y)$ is implemented as $\min(x,y)$ as we try to find the shortest path.
+Travelling Salesman Problem (TSP): This problem is defined as, given a set of cities and the distances between every pair of cities, finding the shortest possible route that visits every city exactly once and returns to the starting point.
+If we regard the set of already visited city $\mathcal{S}$ ending at $i$-th city as the current state, then states that can transfer to current state are those that ending city can reach $k$.
+Therefore, the state transition function for the TSP problem can be expressed as:
+$\text{Answer}[k][i][\mathcal{S}]=\min_{j\in\mathcal{S},j\neq i}(\text{Answer}[k-1][j][\mathcal{S}\setminus\{i\}]+w[j][i]),$
+where $\text{Answer}[k][i][\mathcal{S}]$ represents the cost of the shortest tour that visits all the cities in the set $\mathcal{S}$ and ends at the $i$-th city, the aggregation function $g(x,y)$ is still implemented as $\min(x,y)$ since we aim to find the shortest path, and $w[j][i]$ denotes the distance from city $j$ to city $i$.
+D.2 Proof of Theorem 1
+Assumption 1.
+Each function $f,g$ in (2) can be approximated by constant size MLP.
+Assumption 2.
+The aggregation function $\square$ in (2) is one of min, max, sum, mean.
+The first assumption is mild as MLPs are universal approximators. The second assumption is mild because these are the most commonly used aggregation functions.
+Theorem 3.
+(Expressiveness) Assume the input format is given in (1) and $f,g,\square$ in DP update (2) satisfy the assumptions 1 and 2.
+There exists a log-precision constant-depth and constant-width Transformer that simulates $1$ steps of DP update in (2).
+As a consequence, there exists a log-precision $O(k)$-depth and constant-width Transformer that simulates $k$ steps of DP update in (2).
+Proof.
+Token Embedding and Positional Embedding: The three-dimensional token embedding includes the token type $e^{\text{type1}}$ (00 for answer; $1$ for node; $2$ for edge cost), refined token type $e^{\text{type2}}$ (00 for answer, $1$ for the node tokens in initial states, $2$ for the target node tokens in the edge list, $3$ for the source node tokens in the edge list, $4$ for edge cost), and the token id $e^{\text{token}}$ (from 00 to $|V|-1$).
+The two-dimensional positional embedding includes the embedding for initial state tokens $e^{\text{pos1}}$ (00 for edge list tokens, $1$ for the first two elements of initial states, $2$ for the second two elements of initial states, etc.), embedding for edge list tokens $e^{\text{pos2}}$ (00 for initial state tokens, $1$ for the first three elements of the edge list, $2$ for the second three elements of the edge list, etc.).
+There are also constant-dimensional placeholders to put the states of DP.
+Block 1 - Initial State Broadcast: The goal of the first block is to broadcast the initial states from the initial state token to node tokens.
+(1) Use MLPs to recover the digits of the answer tokens and put them in the first placeholder if $e^{type}_{k}==0$; (2) Copy the first placeholder from answer token to its previous node token by using COPY in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos1}}_{k}-e^{\text{pos1}}_{j})^{2}<\delta\}$; (3) Broadcast the first placeholder with SUM in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{type1}}_{k}-e^{\text{type1}}_{j})^{2}+(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}<\delta\}$.
+Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}]]$.
+Block 2 - Edge Feature Operations: The goal of the second block is to copy the edge features from the edge feature token to the corresponding node tokens.
+(1) Use MLPs to recover the digits of the edge feature tokens and put them in the second placeholder if $e^{\text{type1}}==2$; (2) Copy the second placeholder from the edge feature token to the node token by using SUM in Lemma 1 and setting $\mathcal{S}_{k}=\{j|(e^{\text{pos2}}_{k}-e^{\text{pos2}}_{j})^{2}<\delta\}$.
+Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}]]$.
+Block 3 - Message Preparation: (1) Use MLPs to compute $g$ and place the results in the third placeholder.
+Now the state for every node token $u_{i}$ of the $i$-th edge is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[0][u_{i}],c[u_{i}][v_{i}],g(\text{Answer}[0][u_{i}],c[u_{i}][v_{i}])]$; (2) Use MLPs to clean up the first and second placeholder.
+Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},g(\text{Answer}[0][u_{i}],c[u_{i}][v_{i}])]$
+Block 4 - Message Passing: The goal of the fourth block is to compute $\square$ and $f$.
+(1) Use one or two attention heads (one for max, min, mean aggregations, and two for sum aggregations) to perform the aggregation operation.
+This is achieved by using MEAN or MAX or SUM for the first placeholder and setting $\mathcal{S}_{k}=\{j|(e^{\text{token}}_{k}-e^{\text{token}}_{j})^{2}+(e^{\text{type2}}_{k}-e^{\text{type2}}_{j})^{2}<\delta\}$.
+Now the state for every node token $u_{i}$ is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\square_{v_{j}\in\mathcal{T}(u_{i})}g(\text{Answer}[0][u_{i}],c[u_{i}][v_{j}])]$; (2) Use MLPs to compute $f$.
+Now the state for every node token is $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},f(\square_{v_{j}\in\mathcal{T}(i)}g(\text{Answer}[0][u_{i}],c[u_{i}][v_{j}]))]$.
+After four blocks, the final state for every node token $u_{i}$ is given by $[e^{\text{type1}},e^{\text{type2}},e^{\text{token}},e^{\text{pos1}},e^{\text{pos2}},\text{Answer}[1][u_{i}]]$.
+$\text{Answer}[k][u_{i}]$ can be obtained by repeating the above four blocks $k$ times.
+Lemma 1.
+[12] Let $n\in\mathbb{N}$ be an integer and $\bm{x}_{1},\cdots,\bm{x}_{n}$ be a sequence of vectors where $\bm{x}_{i}=(\tilde{\bm{x}}_{i},r_{i},1)\in[-M,M]^{d+2}$ where $M$ is a large constant.
+Let $\bm{K},\bm{Q},\bm{V}\in\mathbb{R}^{d^{\prime}\times(d+2)}$ be any matrices with $\|\bm{V}\|_{\infty}\leq 1$ and let $0<\rho,\delta<M$ be any real numbers.
+Denote $\bm{q}_{i}=\bm{Q}\bm{x}_{i},\bm{k}_{j}=\bm{K}\bm{x}_{i},\bm{v}_{j}=\bm{V}\bm{x}_{j}$.
+Define a matching set $\mathcal{S}=\{j||\bm{q}_{i}^{T}\bm{k}_{j}|\leq\rho\}$.
+Define two following operations
+  • COPY: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$ with $\bm{u}_{i}=\bm{v}_{\text{pos}(i)}$, where $\text{pos}(i)=\arg\max_{j\in\mathcal{S}_{i}}r_{j}$.
+  • MEAN, MAX, SUM: The output is a sequence of vectors $\bm{u}_{1},\cdots,\bm{u}_{n}$, where $\bm{u}_{i}=\square_{j\in\mathcal{S}_{i}}\bm{v}_{j}$ and $\square$ is min or max or sum or mean.
+Specifically, for any sequence of vectors $\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}$, denote the corresponding output of the attention layer as $\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}$.
+Then, we have $\|\bm{u}_{i}-\bm{o}_{i}\|_{\infty}\leq\epsilon$ for all $i\in[n]$ and $\mathcal{S}\neq\emptyset$.
+D.3 Permutation Invariance Test of LLMs
+We test whether LLMs respect the permutation invariance property in graph problems and the results are given in Figure 5.
+Figure 5: Illustrative Examples of LLMs Failure to Solve Graph Computational Problems under Permutation (i.e., node re-odering). Experiments were conducted for 30 times.
+D.4 Proof of Proposition 1
+Proposition 2.
+Assume the input format is as described in Equation (1) and that the attention mechanism is limited to attending to a constant number of tokens.
+There exists at least one instance of one-step DP update such that no log-precision constant-width constant-depth transformer can simulate.
+Proof.
+We present a proof by contradiction.
+Assume that a token in a Transformer with a constant depth, constant width, and log-precision can attend to only a constant number of nodes.
+Under this assumption, the total information accessible to the token in such a Transformer architecture amounts to $O(\log n)$ bits.
+However, for a graph with $|V|$ nodes, the number of possible outcomes from executing one-step DP is $O(e^{|V|})$, necessitating $\Theta(|V|)$ bits for representation.
+By the pigeonhole principle, this scenario inevitably leads to at least two distinct DP outcomes being represented by the same output sequence generated by the model, thereby constituting a contradiction.
+∎
+D.5 Proof of Theorem 2
+Theorem 4.
+(Spurious correlations of auto-regressive loss) Assume (1) the loss employed is a next-token-prediction loss utilizing cross-entropy, applied to the sub-sequence $v_{1}\ v_{2}\ \cdots\ t$ during training; (2) the output logits are determined by target node $t$ and the current node $v_{i-1}$.
+Let $N_{t,v_{i-1},u}$ be the number of times in the training dataset such that $t$ is the target node, $v_{i-1}$ is the current node and $v_{i}=u$ is the next node.
+The optimal logits for predicting the next node $u$ from current node $v_{i-1}$ towards target node $t$ is given by $\hat{\bm{v}}_{i}[u]=\frac{N_{t,v_{i-1},u}}{\sum_{u}N_{t,v_{i-1},u}}$ if $\sum_{u}N_{t,v_{i-1},u}>0$.
+If $\sum_{u}N_{t,v_{i-1},u}=0$, $\hat{\bm{v}}_{i}[u]$ can be any non-negative number subject to $\sum_{u}\hat{\bm{v}}_{i}[u]=1$.
+Proof.
+We denote $\mathcal{D}$ as the training dataset, $L_{i}$ as the sequence length of the $i$-th sequence in the dataset, $\bm{v}_{i,j}$ as the one hot embedding of the $j$-th token in the $i$-th training sequence, and $\hat{\bm{v}}_{i,j,u}$ as the $u$-th logit at the $j$-th token in the $i$-th sequence.
+The cross-entropy loss is given by
+$-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\bm{v}_{i,j,u}\log\hat{\bm{v}}_{i,j,u}=-\sum_{i\in[|\mathcal{D}|]}\sum_{j=4}^{L_{i}}\mathbb{I}_{u=v_{i,j}}\log\hat{\bm{v}}_{i,j,u}\overset{(a)}{=}-\sum_{t,v_{j-1}}\sum_{u}N_{t,v_{j-1},u}\log\hat{\bm{v}}_{i,j,u}$
+=$\overset{(b)}{=}$ $-\sum_{t,v_{j-1},u}\left(\sum_{u}N_{t,v_{j-1},u}\right)\left[\left(\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}\right)\log\hat{\bm{v}}_{i,j,u}\right],$
+where (a) uses the assumption that the output logits are determined by target node $t$ and the current node $v_{i-1}$.
+In (b), we assume that $\sum_{u}N_{t,v_{i-1},u}\neq 0$.
+The cross-entropy is minimized when $\hat{\bm{v}}_{i,j,u}=\frac{N_{t,v_{j-1},u}}{\sum_{u}N_{t,v_{j-1},u}}$.
+If $\sum_{u}N_{t,v_{i-1},u}=0$, then the corresponding logits will not affect the loss function and $\hat{\bm{v}}_{i,j,u}$ can take any number.
+∎
+Appendix E Supplementary Materials for Training-free Methods
+E.1 Implementation of Baselines
+In this subsection, we present the implementation details of training-free baselines, including LLM’s direct inference, GraphSearch, and ours SGC. Method illustrations are shown in Figure 6.
+Figure 6: Illustration of LLM’s Direct Inference and GraphSearch Method.
+LLM’s Direct Inference: The prompt template for LLM’s direct inference is given in Table 4.
+During experiments, we uniformly apply 1-shot in context learning for LLM’s direct inference of task invocation path.
+For open-sourced LLMs, the temperature parameter is set to $0.2$.
+GraphSearch: The prompt template for GraphSearch is given in Table 5.
+This algorithm conducts an iterative search on the task graph to identify an optimal task invocation path that can best satisfy a given request.
+In each iteration, the neighbors of the last selected task are considered as candidates.
+These candidates are evaluated by LLM for their suitability for the current step (Task Assessment).
+The search process follows a depth-first approach.
+After the task assessment in the final step, a set of potential invocation paths is generated.
+Subsequently, LLM is prompted to select the most appropriate path from these options (Path Selection).
+The GraphSearch algorithm is implemented in three distinct variants, each employing a unique task selection strategy:
+  • GreedySearch consistently selects the task node with the highest score at each step. Although fast and simple, this approach can lead to cascading errors, resulting in degraded performance.
+  • AdaptiveSearch selects tasks with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode. During experiments, we empirically set the score threshold to $3$.
+  • BeamSearch retains the top-$k$ tasks based on the LLM’s assessment scores within candidates.
+Beam search can expand the search space but slightly reduces the efficiency.
+We uniformly set the beam width to $2$.
+Ours SGC: Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC, the Roberta-355M [39] serves as the text encoder.
+For all other datasets and LLMs, the e5-355M [61] configuration is employed.
+E.2 Results of All LLMs
+Table 8 supplements Table 1 with other LLMs. The proposed methods perform consistently better.
+[…]
+TaskBench RestBench
+HuggingFace Multimedia Daily Life TMDB
+LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ # Tok $\downarrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ #Tok $\downarrow$
+Baichuan2 13B Direct 45.85 19.00 2.43 47.57 4.08 2.59 33.45 9.52 3.72 30.87 9.92 1.96
+GreedySearch 30.58 4.89 6.42 18.74 4.45 5.69 15.60 1.61 5.91 22.52 2.98 3.62
+AdaptiveSearch 39.30 10.41 10.81 33.24 9.22 11.17 34.39 12.73 16.71 30.33 10.00 8.71
+BeamSearch 41.06 9.59 24.69 32.24 9.09 21.60 36.18 13.18 23.83 30.97 7.61 9.08
+SGC 56.53 29.94 2.28 56.75 31.62 2.43 62.31 36.69 3.53 32.97 9.11 1.84
+Vicuna 13B Direct 50.46 21.27 2.50 53.57 23.19 2.64 73.70 45.80 3.82 44.66 14.01 2.02
+GreedySearch 52.94 25.73 6.23 46.99 23.11 5.55 42.98 13.33 7.18 45.22 13.69 3.42
+AdaptiveSearch 54.36 25.67 9.81 51.24 24.32 11.25 62.71 31.15 13.92 41.32 7.02 6.51
+BeamSearch 56.64 26.93 24.11 54.09 26.19 25.42 54.55 23.60 24.86 46.91 15.41 7.79
+SGC 59.62 31.98 2.31 61.78 37.60 2.43 83.33 63.77 3.82 48.79 15.99 1.89
+CodeLlama 7B Direct 58.06 29.39 2.44 59.44 30.83 2.57 84.12 62.89 3.82 65.67 41.99 1.94
+GreedySearch 58.71 31.56 5.84 62.83 38.12 5.35 82.51 63.83 7.08 65.51 42.60 3.12
+AdaptiveSearch 60.42 33.18 6.84 62.32 36.81 5.50 83.42 64.15 7.83 65.37 40.64 5.00
+BeamSearch 60.34 31.36 17.95 64.12 38.99 21.48 83.25 63.48 24.48 64.60 40.50 5.78
+SGC 63.98 39.27 2.30 67.04 45.04 2.43 87.73 70.49 3.59 66.15 42.62 1.88
+Mistral 7B Direct 60.60 30.23 2.49 69.83 39.85 2.64 84.26 53.63 3.77 62.23 22.02 1.96
+GreedySearch 65.91 38.13 6.52 58.92 34.72 6.26 75.18 49.47 8.27 60.64 23.18 4.38
+AdaptiveSearch 67.30 38.90 7.68 71.59 44.84 10.66 86.39 63.65 10.92 54.04 21.35 9.99
+BeamSearch 67.13 36.73 25.66 73.55 47.12 31.10 85.87 61.53 39.16 63.41 26.79 11.26
+SGC 67.43 42.08 2.32 74.07 49.90 2.43 87.13 66.49 3.54 64.72 25.67 1.89
+CodeLlama 13B Direct 57.55 28.88 2.45 68.57 41.79 2.59 91.20 76.07 3.88 68.91 43.74 2.02
+GreedySearch 61.67 34.02 5.95 67.98 42.04 4.95 91.50 76.56 5.54 66.67 42.16 3.81
+AdaptiveSearch 60.85 31.66 11.10 68.14 41.71 6.77 91.34 76.09 7.18 63.74 37.17 8.16
+BeamSearch 62.65 34.31 20.14 69.53 43.35 19.51 91.74 76.60 19.19 68.08 42.92 8.88
+SGC 65.51 39.44 2.31 73.32 53.28 2.43 92.96 79.57 3.64 71.40 47.55 1.90
+GPT- 3.5-turbo Direct 73.85 45.73 2.14 82.85 62.07 2.26 96.09 83.65 3.36 81.70 57.52 1.67
+GreedySearch 67.75 43.88 5.29 81.11 63.02 4.92 93.77 81.26 7.36 76.19 50.11 3.06
+AdaptiveSearch 72.18 47.55 7.47 81.86 62.71 5.71 93.79 81.41 8.53 77.57 53.65 5.89
+BeamSearch 75.51 49.62 14.22 83.57 64.50 12.91 95.66 82.72 22.05 81.24 57.98 6.42
+SGC 76.37 50.04 2.02 83.65 63.65 2.09 96.38 86.19 3.16 82.63 59.15 1.61
+GPT- 4-turbo Direct 77.60 52.18 2.19 88.29 69.38 2.28 97.36 84.58 3.37 82.56 56.67 1.75
+GreedySearch 74.75 50.44 5.78 86.81 69.80 5.52 97.36 85.78 7.37 75.34 49.95 3.73
+AdaptiveSearch 76.17 51.30 8.94 88.02 69.99 7.14 97.30 85.80 9.04 81.78 55.15 6.35
+BeamSearch 77.56 52.54 8.98 88.16 70.39 6.90 97.35 85.78 8.99 80.11 51.00 5.18
+SGC 77.79 52.20 2.03 88.54 69.83 2.10 97.35 85.76 3.16 82.27 56.37 1.62
+E.3 Accuracy Results of Training-free Methods
+[…]
+LLM Method HuggingFace Multimedia DailyLife LLM Method HuggingFace Multimedia DailyLife
+Vicuna 13B Direct 8.72 11.20 24.43 CodeLlama 7B Direct 15.00 15.19 47.69
+GreedySearch 10.95 9.34 3.76 GreedySearch 16.20 20.04 45.07
+AdaptiveSearch 10.55 10.37 13.15 AdaptiveSearch 18.79 19.41 46.48
+BeamSearch 12.58 12.03 11.06 BeamSearch 17.00 21.10 45.67
+SGC 16.02 20.12 42.17 SGC 21.20 29.32 55.33
+Mistral 7B Direct 16.36 25.05 44.52 CodeLlama 13B Direct 14.29 24.10 66.40
+GreedySearch 20.45 16.02 29.22 GreedySearch 19.11 24.90 67.00
+AdaptiveSearch 21.88 26.90 49.32 AdaptiveSearch 17.30 24.10 66.80
+BeamSearch 20.45 29.36 45.89 BeamSearch 19.92 25.70 67.20
+SGC 25.15 33.68 52.28 SGC 22.54 36.75 70.80
+GPT- 3.5-turbo Direct 28.95 47.96 81.30 GPT- 4-turbo Direct 33.68 60.56 86.77
+GreedySearch 26.90 52.47 73.17 GreedySearch 33.68 61.37 86.77
+AdaptiveSearch 29.36 51.61 74.59 AdaptiveSearch 33.47 61.17 86.77
+BeamSearch 32.03 52.47 80.87 BeamSearch 33.26 61.57 86.57
+SGC 32.44 51.61 83.13 SGC 34.09 60.97 86.77
+Due to space limitations, we present only the Node-F1 and Link-F1 scores for training-free methods in the main text.
+Here, we provide the Accuracy results in Table 9.
+These results show that integrating SGC significantly enhances accuracy across different LLMs on all datasets, making previously unsolvable planning scenarios manageable and successful.
+E.4 Computational Cost Analysis
+In this subsection, we present a comprehensive efficiency study on inference time of training-free methods and results are shown in Table 10.
+Table 10: Computational Cost Analysis of Training-free Methods. Due to space constraints in the table, some LLMs are abbreviated such as “GPT-3.5” for “GPT-3.5-turbo”.
 Inference Times Comparison on HuggingFace (Seconds)
-Mode Baichuan Vicuna CodeLlama-7B Mistral CodeLlama-13B GPT-3.5 GPT-4
+Method Baichuan Vicuna CodeLlama-7B Mistral CodeLlama-13B GPT-3.5 GPT-4
 Direct 6.0 3.6 10.9 4.5 9.7 2.7 26.1
@@
 SGC 9.9 6.5 18.6 5.7 17.9 3.6 29.5
-Table 7: Efficiency Study of Training-based Methods. We present total training times for both GNN and LM+GNN co-trained modes, and resources needed for fine-tuning LLMs.
+Open-sourced LLMs were deployed as local API services using the FastChat framework [^2] https://github.com/lm-sys/FastChat on a single A100-80G GPU.
+This configuration enables faster and parallel inference.
+Under this setup, LLM’s direct inference requires 3-15 seconds per request.
+GPT-3.5-turbo and GPT-4-turbo are accessed via API, with the latter generally requiring more time.
+GraphSearch requires several minutes to complete a request due to its exhaustive search on the task graph, impacting the efficiency.
+In contrast, SGC achieves comparable efficiency to LLM’s direct inference, as it requires only a single LLM query and both LM and SGC’s forward propagation processes are extremely efficient (typically completing within seconds).
+Note that some discrepancies in reported times, such as Mistral-7B’s GreedySearch taking longer than other modes, may be attributed to variations in the deployment across different A100 services.
+Appendix F Supplementary Materials for Training-based Methods
+F.1 Implementation of Training-based GNNs
+Figure 7: Illustration of our GNN-enhanced Task Planning.
+First, LLMs interpret user request into several manageable steps.
+Then, we leverage GNNs for task retrieval, sequentially matching each step description to a suitable task, finally generating the invocation path.
+LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M model [61] as the LM backbone.
+For the graph encoder, our setup includes a single layer with a hidden dimension of $1024$.
+During the model training, we set the batch size to $512$ and run for $20$ epochs with a learning rate of $31e-$ 3.
+We use the Adam optimizer [24] and implement an early stopping mechanism with a patience of $5$ epochs to prevent over-fitting.
+All experiments are conducted on a single NVIDIA A100-80G GPU.
+Training Data Preparation: From each dataset in TaskBench, we randomly select $3,000$ samples to create the trainset.
+The original data includes specific task steps and corresponding task invocation paths.
+Therefore, we first employ a topological sort to align each task step accurately with corresponding task, forming “<step, ground-truth task>” pairs.
+Then, for each pair, we randomly sample two negative tasks to constitute the “<step, positive task, negative task>” triplets for model training.
+These negative samples are selected based on how textually similar they are to the positive one, creating a robust differentiation challenge for the model.
+Choices of Different Configurations: Our training-based model offers two configuration options: training only the GNN while keeping the LM frozen, or co-training both the LM and GNN.
+Illustrations of these configurations are provided in Figure 8.
+The first configuration is designed to explore the GNN’s capability in task retrieval.
+The latter leverages the LM’s dataset-specific semantic embeddings to enhance performance.
+For the co-training setup, we use a learning rate of $52e-$ 5 and a training duration of $10$ epochs.
+Figure 8: LM+GNN Configuration. We offer two configurations: only training the GNN while keeping the LM frozen (Table 11), or co-training both the LM and GNN (Table 12).
+F.2 Implementation of TAPE and GraphToken
+We adapt TAPE [16] and GraphToken [37] as training-required baselines for task planning. Here, we detail the adaptation processes for each method.
+TAPE: To adapt TAPE for task planning, we reformulate the planning task as a node classification problem, aiming to classify a user request into the appropriate task labels within the task graph.
+Firstly, LLMs interpret user requests by generating high-quality explanatory text, i.e., chain-of-thought reasoning, to understand each request.
+Then, a fine-tuned LM encodes these chain-of-thought texts into latent embeddings.
+The fine-tuning process is based on pairs of textual descriptions and their corresponding ground-truth tasks.
+Finally, GNNs select the suitable tasks by leveraging both the generated text embeddings and the task embeddings.
+For a fair performance comparison, we fine-tune the LM as e5-355M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$.
+The training data for both the LM and GNNs are consistent, utilizing the same split of the training dataset.
+GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-355M.
+Then, a GNN transforms the initial node embedding matrix into $H$.
+The GNN is configured as a $2$-layer GraphSAGE with a hidden dimension of $1024$.
+A mean pooling operation is further applied to the node embeddings to obtain the graph’s overall representation as $\bm{h}_{G}=\text{Mean-Pooling}(H)$.
+Finally, both $\bm{h}_{G}$ and the text embeddings of input instruction and user request are concatenated and fed into the LLM to generate the output planning result.
+During this process, the backbone LLMs are frozen, and only the GNN’s parameters are tuned.
+For each dataset, we train over $4$ epochs using the same $3,000$ training samples consisting of <user request, ground-truth planning> pairs.
+The batch size is set to $16$, with a maximum input length of $512$ and a maximum output length of $300$ for HuggingFace and Multimedia datasets, and $600$ for DailyLife dataset.
+The learning rate is configured to $51e-$ 5.
+F.3 Implementation of Fine-tuning LLMs
+To explore the effectiveness of our proposed framework on fine-tuned LLMs, we employ Supervised Fine-Tuning (SFT) on the CodeLlama-7B and Vicuna-13B models.
+We use the same set of $3,000$ samples from GNN training as the LLMs’ fine-tuning data.
+In fine-tuning the LLM with LoRA [18], we set the lora_r parameter (dimension for LoRA update matrices) to $8$ and the lora_alpha (scaling factor) to $16$.
+The dropout ratio is set to $0.1$, the batch size to $2$, and we conduct training over 2 epochs with a learning rate of $51e-$ 5.
+For HuggingFace dataset, the maximum input length is set to 800, while the maximum output length is 400.
+For Multimedia and Daily Life datasets, which contain a larger number of tasks and require longer textual inputs, we set the maximum input and output length to 1000 and 500, respectively.
+We utilize 2 NVIDIA A100-80G GPUs for fine-tuning the LLMs.
+F.4 Full Results of Training-based GNNs
+In Table 11, we present comprehensive results for all training-based GNNs, including GCN [25], GAT [56], GraphSAGE [15], GIN [66], and Graph Transformer [46].
+To highlight the improvements brought by GNNs, we also include results from the strongest baseline, BeamSearch.
+From the results, it is obvious that all GNN encoders significantly enhance the task planning abilities.
+For instance, when applied to Vicuna-13B on HuggingFace dataset, the introduction of GCN results in a performance improvement of $17.83\%$, GAT contributes to a $17.48\%$ increase, GraphSAGE leads to a $22.59\%$ boost, and GraphTransformer improves predictions by $18.0\%$.
+This conclusion can be generalized across various LLMs and datasets, demonstrating the robust capabilities of GNNs.
+Table 11: Performance of training-based GNNs.
+We also presents the results of BeamSeaerch, the strongest variant from GraphSearch method to provide a comprehensive comparison.
+All GNNs consistently enhance task planning performance across diverse LLMs, showing the effectiveness.
+HuggingFace Tools Multimedia Tools Daily Life APIs
+LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
+Baichuan-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
+BeamSearch 41.06 9.59 32.24 9.09 36.18 13.18
+GCN 57.67 31.47 55.51 30.16 62.11 37.05
+GAT 57.74 31.87 54.95 29.24 62.11 37.05
+GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
+GIN 57.38 31.17 55.08 30.04 62.11 37.05
+GraphTransformer 59.15 34.36 56.06 31.12 64.52 40.14
+Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
+BeamSearch 56.64 26.93 54.09 26.19 54.55 23.60
+GCN 59.46 33.14 62.48 38.89 83.05 62.95
+GAT 59.28 33.39 62.96 39.24 83.05 62.95
+GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
+GIN 59.14 32.33 62.61 38.82 83.05 62.95
+GraphTransformer 59.57 33.47 63.32 39.38 85.41 66.28
+CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
+BeamSearch 60.34 31.36 64.12 38.99 83.25 63.48
+GCN 65.07 40.50 67.46 45.84 87.23 69.27
+GAT 65.20 40.93 67.41 46.46 87.23 69.27
+GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
+GIN 65.52 40.98 66.89 45.41 87.23 69.27
+GraphTransformer 65.83 42.58 68.90 47.20 88.36 71.72
+Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
+BeamSearch 67.13 36.73 73.55 47.12 85.87 61.53
+GCN 66.54 40.74 73.34 50.76 86.39 65.49
+GAT 66.77 40.74 73.36 50.20 86.39 65.49
+GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
+GIN 66.69 40.79 72.89 50.44 86.39 65.49
+GraphTransformer 68.26 43.08 73.80 51.45 88.25 67.84
+Direct 57.55 28.88 68.57 41.79 91.20 76.07
+BeamSearch 62.65 34.31 69.53 43.35 91.74 76.60
+GCN 66.22 41.05 72.99 52.18 91.83 77.88
+GAT 66.29 41.28 74.08 53.56 91.83 77.88
+GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
+GIN 66.40 40.89 73.62 53.15 91.83 77.88
+CodeLlama-13B GraphTransformer 66.70 42.07 74.72 54.10 93.81 80.44
+GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
+BeamSearch 75.51 49.62 83.57 64.50 95.66 82.72
+GCN 76.93 51.43 84.92 65.05 96.38 86.15
+GAT 75.63 49.36 84.77 65.48 96.38 86.15
+GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
+GIN 76.86 51.00 84.14 64.30 96.38 86.15
+GraphTransformer 77.61 52.30 84.21 64.32 96.38 86.19
+GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
+BeamSearch 77.56 52.54 88.16 70.39 97.35 85.78
+GCN 77.01 50.49 88.56 69.60 97.10 85.22
+GAT 76.41 49.66 88.43 69.52 97.10 85.22
+GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
+GIN 77.74 51.02 88.05 69.13 97.36 84.58
+GraphTransformer 78.47 52.17 88.07 68.71 97.32 85.57
+Table 12: Performance of Training-based GNNs under the LM+GNN Co-trained Mode. Simultaneous training of LM and GNN yields significant performance improvements.
+HuggingFace Tools Multimedia Tools Daily Life APIs
+LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
+Baichuan-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
+SGC 60.97 36.12 56.02 31.36 64.84 40.00
+GCN 60.68 36.31 57.82 32.87 64.73 38.92
+GAT 60.39 35.37 57.24 32.19 64.46 40.14
+GraphSAGE 59.76 35.59 57.97 33.29 63.21 38.10
+GIN 60.31 35.82 56.62 31.53 63.55 38.19
+GraphTransformer 60.76 36.82 56.82 31.40 64.88 40.23
+Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
+SGC 64.40 38.97 65.12 41.63 84.74 65.90
+GCN 62.06 35.49 65.02 40.63 85.22 66.93
+GAT 63.06 36.97 64.58 40.30 85.63 67.11
+GraphSAGE 62.82 37.04 65.89 42.18 84.23 65.44
+GIN 62.09 35.33 64.44 40.67 85.31 66.83
+GraphTransformer 62.11 36.01 64.57 40.17 85.42 66.55
+CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
+SGC 67.47 43.58 69.61 48.24 87.98 70.63
+GCN 67.03 43.24 69.33 47.60 87.88 70.40
+GAT 67.12 42.96 68.62 46.17 88.59 71.64
+GraphSAGE 67.19 42.94 70.00 48.28 87.81 70.20
+GIN 66.62 42.34 69.00 47.72 88.45 71.53
+GraphTransformer 67.12 43.08 69.27 47.96 88.43 71.59
+Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
+SGC 69.04 44.22 76.09 54.91 87.58 66.70
+GCN 67.72 43.02 76.79 54.90 87.87 67.13
+GAT 67.54 43.56 76.26 53.94 87.86 67.30
+GraphSAGE 67.61 43.14 76.96 55.46 87.61 66.75
+GIN 68.95 43.97 76.47 54.95 87.75 67.07
+GraphTransformer 67.94 43.52 77.06 55.39 87.76 67.00
+CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
+SGC 70.14 45.20 75.65 55.45 93.45 79.89
+GCN 69.39 45.18 76.03 55.22 93.38 79.74
+GAT 69.68 45.57 75.24 54.99 94.06 80.96
+GraphSAGE 68.92 44.85 76.28 55.41 93.30 79.51
+GIN 69.01 44.76 74.72 53.91 94.24 81.23
+GraphTransformer 69.52 45.68 75.46 55.14 93.98 81.06
+GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
+SGC 77.87 52.86 85.95 66.95 96.39 86.16
+GCN 77.72 52.58 85.84 66.92 96.33 86.06
+GAT 77.49 52.30 85.81 66.97 96.38 86.15
+GraphSAGE 77.87 53.04 85.51 66.56 96.34 86.09
+GIN 77.73 52.36 85.63 66.69 96.38 86.19
+GraphTransformer 77.78 52.79 86.09 67.26 96.33 86.06
+GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
+SGC 78.44 52.84 89.09 70.52 97.38 85.85
+GCN 78.33 52.75 89.00 70.24 97.34 85.67
+GAT 78.37 52.43 88.99 70.48 97.32 85.56
+GraphSAGE 78.49 52.62 88.86 70.25 97.42 85.80
+GIN 78.45 53.07 88.74 69.84 97.42 85.80
+GraphTransformer 78.30 52.27 88.90 70.24 97.42 85.80
+F.5 Performance of LM+GNN Co-trained Mode
+During our main experiments, the LM backbone remains frozen while only the GNN is trained to automatically learn the alignment between implicit step descriptions and suitable tasks, facilitating task retrieval.
+In this subsection, we conduct a supplementary study where the parameters of a pre-trained LM are also tuned along with GNN during model training.
+The model configuration is illustrated in Figure 8, and the results are presented in Table 12.
+The results demonstrate that, compared to the GNN-only tunable mode, co-training LM+GNN can lead to further performance improvements.
+This enhancement occurs because the language model acquires task-specific semantics, which makes the representations more discriminative and boosts the GNN’s effectiveness in task retrieval.
+Additionally, it is noted that under the co-training setup, the differences between various GNN encoders are relatively minor, with performance variations across GNNs for a specific LLM on any dataset remaining within $2\%$.
+F.6 Computational Cost Analysis
+In this subsection, we provide the computational costs of training-based methods: training time for GNN or LM+GNN co-trained, and resources needed for fine-tuning LLMs. Results are shown in Table 13.
+Table 13: Computational Cost Analysis of Training-based Methods. We present total training times for both GNN and LM+GNN co-trained modes, and resources needed for fine-tuning LLMs.
 Training GNNs
 Time (Seconds)
-Mode Configuration # Param HuggingFace Multimedia Daily Life
+Mode Configuration # Parameters HuggingFace Multimedia Daily Life
 GNN-only GCN 1,049,600 136.5 136.1 237.7
@@
 GIN 2,099,200 134.2 134.9 233.6
-TransformerConv 4,198,400 135.0 150.2 233.7
+GraphTransformer 4,198,400 135.0 150.2 233.7
 LM+GNN Co-trained LM+SGC 335,141,889 743.1 323.3 384.7
@@
 LM+GIN 337,241,088 735.6 361.8 382.9
-LM+TransformerConv 339,340,288 741.0 362.7 405.8
+LM+GraphTransformer 339,340,288 741.0 362.7 405.8
 Fine-tuning LLMs
 Device & Time (Hours)
-LLM # Param (# Tunable Param) HugggingFace Multimedia Daily Life
+LLM # Param (# Trainable Param) HuggingFace Multimedia Daily Life
 CodeLlama-7B 6,742,740,992 (4,194,304) 2$\times$×A100 7.0 1$\times$×A100 13.8 2$\times$×A100 9.5
 Vicuna-13B 13,022,417,920 (6,553,600) 1$\times$×A100 17.8 2$\times$×A100 10.3 2$\times$×A100 19.0
-Efficiency of Training-free Modes: Open-sourced LLMs were deployed as local API services using the FastChat framework on a single A100-80G GPU.
-This configuration enables faster and parallel inference.
-Under this setup, LLM’s direct inference requires 3-15 seconds per request.
-GPT-3.5-turbo and GPT-4-turbo are accessed via API, with the latter generally requiring more time.
-GraphSearch requires several minutes to complete a request due to its exhaustive search on the task graph, impacting the efficiency.
-In contrast, SGC achieves comparable efficiency to LLM’s direct inference, as it requires only a single LLM query and both LM and SGC’s forward propagation processes are extremely efficient.
-Note that some discrepancies in reported times, such as Mistral-7B’s GreedySearch taking longer than other modes, may be attributed to variations in the deployment across different A100 GPUs.
 Efficiency of Training GNNs: During experiments, each dataset shares the same GNN configuration: $1$ single layer with a hidden dimension of $1024$.
-Therefore, for each GNN, the number of parameters remains consistent across datasets.
+Therefore, for each GNN, its number of parameters remains consistent across datasets.
 The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request.
-For LM+GNN co-trained mode, where e5-355M serves as the LM backbone, training times increase to approximately 6-12 minutes.
+For LM+GNN co-trained mode, where e5-355M [61] serves as the LM backbone, training times increase to approximately 6-12 minutes.
 In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
 Efficiency of Fine-tuning LLMs: Fine-tuning a LLM with $3,000$ training samples over $2$ epochs requires huge time, typically 10-20 hours on one or two A100-80G GPU devices.
-F.4 Case Study
-Effectiveness of GNN (Qualitative Result): We show two cases in Figure 9 where the results of LLM’s direct inference, BeamSearch, and GraphSAGE are compared.
+Appendix G Experiments on Task Parameter Prediction
+To assess the quality of task planning, we primarily focus on the predicted tasks and their dependencies, leaving the parameters for invoking these tasks undiscussed.
+As LLMs’ direct inference for solely predicting tasks has proven unsatisfactory (Figure 2), relying on LLMs alone to directly predict these parameters is unreliable.
+In this section, we will first demonstrate that it is quite straightforward for LLMs to fill in the parameters given a planned task sequence, which is a simple extension of our framework.
+Furthermore, we will empirically show that with more accurately planned tasks, i.e., those retrieved by GNNs, LLMs can intelligently fill in the parameters.
+G.1 Prompting LLMs to Fill in Parameters
+Process: Recall that our framework enables more accurate invoked task sequences, e.g., $\{v_{1},\ldots,v_{n}\}$, for a specific user request.
+To finalize the invocation sequence, just adding an additional LLM query can complete the invocation parameters for each task.
+Specifically, by providing the original user request, planned tasks, and detailed descriptions including each task’s input and output requirements, LLMs can be intelligently prompted to fill in the invocation parameters for each planned task, resulting in a well-structured invocation sequence ready for execution.
+Prompts are provided in Table 6.
+[…]
+G.2 Empirical Results of LLMs Predicted Parameters
+Experimental Setup: We conduct this supplementary experiment on the HuggingFace dataset from TaskBench [44].
+Specifically, we compare the directly predicted invocation parameters from LLMs (denoted as Direct) with the parameters filled automatically by LLMs using our GNN-retrieved tasks (querying prompt as shown in Table 6).
+For GNNs, we consider both training-free SGC and training-required GraphSAGE.
+Notably, with SGC, since the query for completing parameters is also training-free, the entire pipeline can be deployed without any extensive model training or labeled data.
+We adopt evaluation metrics from TaskBench, including Parameter-Type F1-Score (Param t-F1) and Parameter-Value F1-Score (Param v-F1).
+These metrics measure the accuracy of the predicted parameter types and concrete values, respectively.
+For example, when filling in the invocation for Pose Detection, the ground-truth parameter type is Image, and the ground-truth value is ‘example.jpg’.
+[…]
+LLM Method t-F1 $\uparrow$ v-F1 $\uparrow$ LLM Method t-F1 $\uparrow$ v-F1 $\uparrow$
+Direct 38.77 18.56 Direct 44.62 33.24
+SGC 58.13${}_{\textbf{+19.36}}$ 39.64${}_{\textbf{+21.08}}$ SGC 57.74${}_{\textbf{+13.21}}$ 43.21${}_{\textbf{+9.97}}$
+Mistral-7B GraphSAGE 59.07${}_{\textbf{+20.30}}$ 41.40${}_{\textbf{+22.84}}$ CodeLlama-13B GraphSAGE 59.49${}_{\textbf{+14.87}}$ 45.54${}_{\textbf{+12.30}}$
+Direct 62.42 48.27 Direct 70.73 55.54
+SGC 68.13${}_{\textbf{+5.71}}$ 54.34${}_{\textbf{+6.07}}$ SGC 72.91${}_{\textbf{+2.18}}$ 58.02${}_{\textbf{+2.48}}$
+GPT-3.5-turbo GraphSAGE 71.19${}_{\textbf{+8.77}}$ ${}_{\textbf{+8.56}}$ GPT-4-turbo GraphSAGE 73.09${}_{\textbf{+2.36}}$ 58.20${}_{\textbf{+2.66}}$
+Observation: From the results shown in Table 14, we can conclude that: (1) LLMs’ direct inference of invocation arguments is unsatisfactory.
+Even for the strongest LLM, GPT-4-turbo, the Parameter-Value F1 score is only $55\%$, which is far from expectations.
+(2) With accurate planning provided by GNNs, LLMs can leverage their inherent reasoning abilities to analyze the context and correctly fill in the parameters.
+Across four LLMs, improvements of $2\%$ to $20\%$ can be observed, highlighting the advantages of GNN-enhanced planning.
+(3) Our method is well-suited for training-free scenarios, as the tasks retrieved by the training-free SGC already enable LLMs to infer the parameters effectively, with only a minor performance gap compared to the training-required GraphSAGE.
+Appendix H Case Studies
+Effectiveness of GNNs: We show two cases in Figure 9 where the results of LLM’s direct inference, BeamSearch, and GraphSAGE are compared.
 Due to space constraints and issues such as LLM’s output content decoding errors or invalid paths, we present only the first four valid paths searched by BeamSearch on the task graph.
 From the cases, we can conclude that BeamSearch relies on LLM’s inherent reasoning abilities.
-Although LLM can explore the ground-truth invocation path on the task graph, their final solutions are usually not the optimal as containing either the hallucination or wrongly invoked tasks due to limited instruction following and reasoning abilities.
-On the contrary, GNN can effectively align decomposed steps with suitable tasks, accurately hitting the ground-truth result and enhancing task planning.
+Although LLM can explore the ground-truth invocation path on the task graph, their final solutions are usually not the optimal as either containing the hallucination or wrongly invoked tasks due to limited instruction following and reasoning abilities.
+On the contrary, GNN can effectively align decomposed steps with suitable tasks, accurately achieving the ground-truth result and enhancing task planning.
 Figure 9: Case Study of GNN’s Effectiveness.
@@
 As the case shown in Figure 10, step 2 is ambiguously described as “Segment the image and identify the tabular data”, which actually incorporates two distinct steps.
-This ambiguty causes GNN to choose the unsuitable Tabular Classification instead of the correct Image Segmentation.
+This ambiguity causes GNN to choose the unsuitable Tabular Classification instead of the correct Image Segmentation.
 Since tasks are selected sequentially on the task graph, where the next task is a neighbor of the current selection, such an error can prevent the exploration of the next appropriate task, as it may not be a neighbor of the current, incorrect selection.
@@
 Our framework relies heavily on the quality of decomposed task steps.
-Ambiguous steps (step 2 actually incorporates two steps) may mislead GNN to select the wrong task.
+Ambiguous steps (step 2 which actually incorporates two steps) may mislead GNN to select the wrong task.
 Diagnosing GraphSearch with GPT-4-turbo: In our experiments, GraphSearch brings marginal or even decreased performance for GPT-4-turbo across most experimental datasets, contradicting current research [32, 31] which suggests that exhaustive search strategies can enhance the performance of more powerful LLMs.
 To provide a detailed analysis, we show three types of cases in Figure 11: Successful cases, Failure cases, as well as Maintaining cases:
-  • Successful Cases: As shown in the figure, despite inaccuracies in task decomposition (GPT-4-turbo predicts an extra step compared to the three-step ground truth), its inherent reasoning abilities and the knowledge explored along the task graph can fix these mistakes, hitting the ground-truth.
-  • Failure Cases: In the failure case, although GPT-4-turbo identifies the ground-truth solution during searching process, the final decision includes an incorrect task.
-This occurs because the final solution selection demands complex reasoning abilities of LLM, as the context contains long textual information from different aspects: the full task graph, user request, all searched paths, and related instructions (Prompt template is shown in Table 4).
+  • Successful Cases: As shown in the figure, despite inaccuracies in task decomposition (GPT-4 predicts an extra step compared to the three-step ground truth), its inherent reasoning abilities and the knowledge explored along the task graph can fix these mistakes, hitting the ground-truth.
+  • Failure Cases: In the failure case, although GPT-4 identifies the ground-truth solution during searching process, the final decision includes an incorrect task.
+This occurs because the final solution selection demands complex reasoning abilities of LLM, as the context contains long textual information from different aspects: the full task graph, user request, all searched paths, and related instructions.
 The demanding context and reasoning challenge exceed GPT-4’s capabilities, leading to errors.
-  • Maintaining Cases: These occur when GraphSearch merely replicates the result of GPT-4-turbo’s direct inference, indicating no added benefit from exploring the task graph.
-In these instances, despite navigating the graph, GPT-4 fails to self-correct inaccuracies due to inherent reasoning limitations.
-We emphasize that the results of GraphSearch with GPT-4-turbo, tend to “maintaining cases”.
+  • Maintaining Cases: These occur when GraphSearch merely replicates the result of LLM’s direct inference, indicating no added benefit from exploring the task graph.
+In these instances, despite navigating the graph, the LLM fails to self-correct inaccuracies due to inherent reasoning limitations.
+We emphasize that the results of GraphSearch with GPT-4-turbo, tend to “maintaining” cases.
 Under such conditions, even a few failures can degrade overall performance, explaining why GraphSearch does not consistently enhance performance for GPT-4-turbo.
-Figure 11: Diagnosing GraphSearch with GPT-4-turbo.
-We provide three types of cases: (1) Successful Cases where GPT-4 can leverage inherent reasoning abilities to select the optimal invocation path.
-(2) Failure Cases where GPT-4 miss in the extremely long context, containing the whole task graph, all searched paths, and instructions, selecting an unsatisfactory path.
-(3) Maintaining Cases where the searched result is the same as direct inference result, and those wrongly predicted tasks can still not be refined even under exhaustive search.
-Table 8: Performance of training-based GNNs.
-We also presents the results of BeamSeaerch, the strongest variant from GraphSearch method to provide a comprehensive comparison.
-All GNNs consistently enhance task planning performance across diverse LLMs, showing the effectiveness.
-HuggingFace Tools Multimedia Tools Daily Life APIs
-LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
-Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
-BeamSearch 41.06 9.59 32.24 9.09 36.18 13.18
-GCN 57.67 31.47 55.51 30.16 62.11 37.05
-GAT 57.74 31.87 54.95 29.24 62.11 37.05
-GraphSAGE 59.32 34.36 56.15 31.60 65.18 40.49
-GIN 57.38 31.17 55.08 30.04 62.11 37.05
-TransformerConv 59.15 34.36 56.06 31.12 64.52 40.14
-Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
-BeamSearch 56.64 26.93 54.09 26.19 54.55 23.60
-GCN 59.46 33.14 62.48 38.89 83.05 62.95
-GAT 59.28 33.39 62.96 39.24 83.05 62.95
-GraphSAGE 61.86 35.68 63.71 39.88 86.07 67.63
-GIN 59.14 32.33 62.61 38.82 83.05 62.95
-TransformerConv 59.57 33.47 63.32 39.38 85.41 66.28
-CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
-BeamSearch 60.34 31.36 64.12 38.99 83.25 63.48
-GCN 65.07 40.50 67.46 45.84 87.23 69.27
-GAT 65.20 40.93 67.41 46.46 87.23 69.27
-GraphSAGE 66.67 43.03 67.97 46.31 88.53 72.02
-GIN 65.52 40.98 66.89 45.41 87.23 69.27
-TransformerConv 65.83 42.58 68.90 47.20 88.36 71.72
-Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
-BeamSearch 67.13 36.73 73.55 47.12 85.87 61.53
-GCN 66.54 40.74 73.34 50.76 86.39 65.49
-GAT 66.77 40.74 73.36 50.20 86.39 65.49
-GraphSAGE 68.12 43.09 75.51 52.94 87.51 66.57
-GIN 66.69 40.79 72.89 50.44 86.39 65.49
-TransformerConv 68.26 43.08 73.80 51.45 88.25 67.84
-Direct 57.55 28.88 68.57 41.79 91.20 76.07
-BeamSearch 62.65 34.31 69.53 43.35 91.74 76.60
-GCN 66.22 41.05 72.99 52.18 91.83 77.88
-GAT 66.29 41.28 74.08 53.56 91.83 77.88
-GraphSAGE 67.30 42.41 74.93 54.52 93.84 80.38
-GIN 66.40 40.89 73.62 53.15 91.83 77.88
-CodeLlama-13B TransformerConv 66.70 42.07 74.72 54.10 93.81 80.44
-GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
-BeamSearch 75.51 49.62 83.57 64.50 95.66 82.72
-GCN 76.93 51.43 84.92 65.05 96.38 86.15
-GAT 75.63 49.36 84.77 65.48 96.38 86.15
-GraphSAGE 77.90 52.68 85.29 65.80 96.43 86.26
-GIN 76.86 51.00 84.14 64.30 96.38 86.15
-TransformerConv 77.61 52.30 84.21 64.32 96.38 86.19
-GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
-BeamSearch 77.56 52.54 88.16 70.39 97.35 85.78
-GCN 77.01 50.49 88.56 69.60 97.10 85.22
-GAT 76.41 49.66 88.43 69.52 97.10 85.22
-GraphSAGE 78.76 52.53 88.63 69.65 97.34 85.67
-GIN 77.74 51.02 88.05 69.13 97.36 84.58
-TransformerConv 78.47 52.17 88.07 68.71 97.32 85.57
-Table 9: Performance of Training-based GNNs under the LM+GNN Co-trained Mode. Simultaneous training of LM and GNN yields significant performance improvements.
-HuggingFace Tools Multimedia Tools Daily Life APIs
-LLM Method n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$ n-F1 $\uparrow$ l-F1 $\uparrow$
-Baichuan2-13B Direct 45.85 19.00 47.57 4.08 33.45 9.52
-SGC 60.97 36.12 56.02 31.36 64.84 40.00
-GCN 60.68 36.31 57.82 32.87 64.73 38.92
-GAT 60.39 35.37 57.24 32.19 64.46 40.14
-GraphSAGE 59.76 35.59 57.97 33.29 63.21 38.10
-GIN 60.31 35.82 56.62 31.53 63.55 38.19
-TransformerConv 60.76 36.82 56.82 31.40 64.88 40.23
-Vicuna-13B Direct 50.46 21.27 53.57 23.19 73.70 45.80
-SGC 64.40 38.97 65.12 41.63 84.74 65.90
-GCN 62.06 35.49 65.02 40.63 85.22 66.93
-GAT 63.06 36.97 64.58 40.30 85.63 67.11
-GraphSAGE 62.82 37.04 65.89 42.18 84.23 65.44
-GIN 62.09 35.33 64.44 40.67 85.31 66.83
-TransformerConv 62.11 36.01 64.57 40.17 85.42 66.55
-CodeLlama-7B Direct 58.06 29.39 59.44 30.83 84.12 62.89
-SGC 67.47 43.58 69.61 48.24 87.98 70.63
-GCN 67.03 43.24 69.33 47.60 87.88 70.40
-GAT 67.12 42.96 68.62 46.17 88.59 71.64
-GraphSAGE 67.19 42.94 70.00 48.28 87.81 70.20
-GIN 66.62 42.34 69.00 47.72 88.45 71.53
-TransformerConv 67.12 43.08 69.27 47.96 88.43 71.59
-Mistral-7B Direct 60.60 30.23 69.83 39.85 84.26 53.63
-SGC 69.04 44.22 76.09 54.91 87.58 66.70
-GCN 67.72 43.02 76.79 54.90 87.87 67.13
-GAT 67.54 43.56 76.26 53.94 87.86 67.30
-GraphSAGE 67.61 43.14 76.96 55.46 87.61 66.75
-GIN 68.95 43.97 76.47 54.95 87.75 67.07
-TransformerConv 67.94 43.52 77.06 55.39 87.76 67.00
-CodeLlama-13B Direct 57.55 28.88 68.57 41.79 91.20 76.07
-SGC 70.14 45.20 75.65 55.45 93.45 79.89
-GCN 69.39 45.18 76.03 55.22 93.38 79.74
-GAT 69.68 45.57 75.24 54.99 94.06 80.96
-GraphSAGE 68.92 44.85 76.28 55.41 93.30 79.51
-GIN 69.01 44.76 74.72 53.91 94.24 81.23
-TransformerConv 69.52 45.68 75.46 55.14 93.98 81.06
-GPT-3.5-turbo Direct 73.85 45.73 82.85 62.07 96.09 83.65
-SGC 77.87 52.86 85.95 66.95 96.39 86.16
-GCN 77.72 52.58 85.84 66.92 96.33 86.06
-GAT 77.49 52.30 85.81 66.97 96.38 86.15
-GraphSAGE 77.87 53.04 85.51 66.56 96.34 86.09
-GIN 77.73 52.36 85.63 66.69 96.38 86.19
-TransformerConv 77.78 52.79 86.09 67.26 96.33 86.06
-GPT-4-turbo Direct 77.60 52.18 88.29 69.38 97.36 84.58
-SGC 78.44 52.84 89.09 70.52 97.38 85.85
-GCN 78.33 52.75 89.00 70.24 97.34 85.67
-GAT 78.37 52.43 88.99 70.48 97.32 85.56
-GraphSAGE 78.49 52.62 88.86 70.25 97.42 85.80
-GIN 78.45 53.07 88.74 69.84 97.42 85.80
-TransformerConv 78.30 52.27 88.90 70.24 97.42 85.80
+Figure 11: Diagnosing GraphSearch for GPT-4.
+We provide three types of cases: Success Cases where GPT-4 can leverage inherent reasoning abilities to select the optimal invocation path, which may even fix wrongly decomposed task steps.
+Failure Cases where GPT-4 miss in the extremely long context, containing the whole task graph, all searched paths, and instructions, selecting an unsatisfactory path.
+Maintain Cases where the searched result is the same as direct inference result, and those wrongly predicted tasks can still not be refined even under exhaustive search.

### v3

@@
 Figure 6 illustrates this procedure.
-Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT)
+Figure 1: Illustration of Task Planning in Language Agents (e.g., HuggingGPT [46])
 3 Graph Formulation and Insights
@@
 Suppose we are selecting the node for the $i$-th decomposed step.
-First, we utilize a small pre-trained language model, e5-355M [61], to embed the $i$-th decomposed step.
+First, we utilize a small pre-trained language model, e5-335M [62], to embed the $i$-th decomposed step.
 The resulting embedding is denoted as $\bm{x}_{i}^{\text{step}}$.
-Second, for the task graph, we first use the same pre-trained language model e5-355M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node.
+Second, for the task graph, we first use the same pre-trained language model e5-335M to convert each node’s description into embeddings, denoted as the node feature $\bm{h}^{0}_{v}$, where the superscript indicates the layer and the subscript represents the node.
 Then we adopt a $K$-layer SGC [65] to compute the final node embeddings, resulting in $\bm{h}_{v}=\bm{h}_{v}^{(K)}$.
@@
 Particularly, $v_{1}$ can be selected from the whole graph.
-As e5-355M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
+As e5-335M is pre-trained and SGC is parameter-free, the proposed method requires no additional training and can be effectively applied in a zero-shot manner.
 4.3 A Training-required GNN-based Approach
@@
 Therefore, we adopt the Bayesian Personalized Ranking (BPR) loss [41] designed for recommendation with binary rankings.
-The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-355M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task.
+The loss function is given by $\ell=\sum_{(\bm{x}^{\text{step}},v,v^{\prime})}-\log\;\sigma(\langle\bm{h}_{v},\bm{x}^{\text{step}}\rangle-\langle\bm{h}_{v^{\prime}},\bm{x}^{\text{step}}\rangle)$, where $\bm{x}^{\text{step}}$ represents the embedding of the step’s textual description generated by e5-335M, $v$ is the ground-truth task, and $v^{\prime}$ is a negative task.
 We select negative tasks that are textually similar to the positive task, and for computational efficiency, we limit our selection to $2$ negative tasks per positive task.
-The trainable parameters may merely include GNNs or both GNNs and e5-355M with illustrative configurations shown in Figure 8 in Appendix.
+The trainable parameters may merely include GNNs or both GNNs and e5-335M with illustrative configurations shown in Figure 8 in Appendix.
 5 Experiments and Analysis
@@
 We compare the performance across three training-free methods: (1) LLM’s Direct Inference is introduced in Section 2.2.
-(2) GraphSearch [32, 51, 31] leverages the classic graph search method to generate the candidate nodes and use LLMs to give a score for node selection.
+(2) GraphSearch [33, 52, 32] leverages the classic graph search method to generate the candidate nodes and uses LLMs to give a score for node selection.
 Given a step, GreedySearch consistently selects the node with the highest score and adjacent to the previous task node; AdaptiveSearch selects the nodes with scores above a fixed threshold, adjusting the breadth of the search space in an adaptive mode; BeamSearch retains the $k$ nodes with highest scores.
@@
 A unique feature of GPT-4-turbo is its ability to manage ChatGPT-plugins, and it may have been specially trained on task planning datasets.
-In addition, the pre-trained language model used for feature extraction in SGC is e5-355M, which may not be sufficiently powerful to effectively analyze GPT-4’s output.
+In addition, the pre-trained language model used for feature extraction in SGC is e5-335M, which may not be sufficiently powerful to effectively analyze GPT-4’s output.
 A detailed diagnostic analysis of cases involving GPT-4 is provided in Figure 11 in Appendix.
@@
 The TMDB dataset is excluded due to its limited sample size.
-Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-355M), whose role is to generate node embeddings derived from task names and descriptions.
+Throughout our experiments, we trained a spectrum of GNN variants, both with and without co-training the small LM (i.e., e5-335M), whose role is to generate node embeddings derived from task names and descriptions.
 Owing to space constraints, we only show the performance of GraphSAGE in the main text, relegating a detailed comparison of all the situations to Table 11 and Table 12 in Appendix.
@@
 Observations: From Table 2, we observe a significant improvement in performance when employing a training-based GraphSAGE approach over the training-free method.
-However, the co-training of GNNs with e5-355M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance.
+However, the co-training of GNNs with e5-335M does not yield a marked improvement, suggesting that message passing is the crucial element for enhancing performance.
 Further analysis across a broad spectrum of GNNs (as shown in Table 11 and Table 12) reveals that powerful GNNs, such as GINs, perform similarly to networks perceived as less complex, like GCNs, and even underperform compared to GraphSAGE.
@@
 In addition, GraphToken is limited to open-sourced LLMs.
-The proposed approaches also greatly improve the parameter prediction performance, as given in Appendix G (e.g.
-$9\%$ improvement to to GPT-3.5 and $3\%$ improvement to GPT-4).
+Besides, our proposed approaches also greatly boost the parameter prediction performance, as given in Appendix G (e.g.
+$9\%$ improvement to GPT-3.5-turbo and $3\%$ improvement to GPT-4-turbo).
 Efficiency: The details of the training time are given in Table 13.
-The training cost is remarkably low because we use e5-355M [61] as the text embedding model for GNNs.
+The training cost is remarkably low because we use e5-335M [62] as the text embedding model for GNNs.
 If the trainable parameters are limited to the GNNs alone, training typically concludes within just 3 minutes.
-Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-355M model.
+Furthermore, the training duration extends to only 15 minutes when GNNs are jointly trained with e5-335M model.
 This efficiency stands in stark contrast to the 10-20 hours required for tuning open-sourced LLMs.
@@
 The analysis in Section 3.3 aims to shed light on the empirical observations reported in these studies.
-Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [52], GraphLLM [9], and GraphToken [37].
+Considering these negative results, a new line of research has emerged that utilizes the output of GNNs as tokens for LLMs, as seen in GraphGPT [53], GraphLLM [9], GraphToken [38], and G-Retriever [17].
 These approaches have demonstrated significant improvements in performance on GNN-related tasks.
@@
 Ours SGC: Regarding the choices of LM backbones, for integrating GPT-3.5-turbo and GPT-4-turbo with SGC, the Roberta-355M [40] serves as the text encoder.
-For all other datasets and LLMs, the e5-355M [61] configuration is employed.
+For all other datasets and LLMs, the e5-335M [62] configuration is employed.
 E.2 Results of All LLMs
@@
 Then, we leverage GNNs for task retrieval, sequentially matching each step description to a suitable task, finally generating the invocation path.
-LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-355M model [61] as the LM backbone.
+LM and GNN Configuration: For training-based GNNs, we uniformly use the e5-335M model [62] as the LM backbone.
 For the graph encoder, our setup includes a single layer with a hidden dimension of $1024$.
@@
 Finally, GNNs select the suitable tasks by leveraging both the generated text embeddings and the task embeddings.
-For a fair performance comparison, we fine-tune the LM as e5-355M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$.
+For a fair performance comparison, we fine-tune the LM as e5-335M and configure the GNN as a $2$-layer GraphSAGE with a hidden dimension of $1024$.
 The training data for both the LM and GNNs are consistent, utilizing the same split of the training dataset.
-GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-355M.
+GraphToken: For adapting GraphToken, we first encode the task graph $G(V,E,X)$ by computing each node’s representation $\bm{x}_{v}$ by feeding its descriptive text into the pre-trained LM, e5-335M.
 Then, a GNN transforms the initial node embedding matrix into $H$.
@@
 The parameter scales for GNN variants range from 1M to 4M, and the total training time for each dataset requires only 2-4 minutes, comparable to the time taken by GraphSearch to fulfill a single request.
-For LM+GNN co-trained mode, where e5-355M [61] serves as the LM backbone, training times increase to approximately 6-12 minutes.
+For LM+GNN co-trained mode, where e5-335M [62] serves as the LM backbone, training times increase to approximately 6-12 minutes.
 In summary, both modes demonstrate high efficiency, with total training times spanning minutes, showcasing their ability to rapidly adapt to new task planning scenarios.
@@
 (2) With accurate planning provided by GNNs, LLMs can leverage their inherent reasoning abilities to analyze the context and correctly fill in the parameters.
-Across four LLMs, improvements of $2\%$ to $20\%$ can be observed, highlighting the advantages of GNN-enhanced planning.
+Across four LLMs, improvements of $3\%$ to $23\%$ can be observed, highlighting the advantages of GNN-enhanced planning.
 (3) Our method is well-suited for training-free scenarios, as the tasks retrieved by the training-free SGC already enable LLMs to infer the parameters effectively, with only a minor performance gap compared to the training-required GraphSAGE.